### v3

# Adaptive Rank, Reduced Forgetting: Knowledge Retention in Continual Learning Vision-Language Models with Dynamic Rank-Selective LoRA

Abstract
We investigate whether the pre-trained knowledge of vision-language models (VLMs), such as CLIP, can be retained or even enhanced during continual learning (CL) while absorbing knowledge from a data stream. Existing methods often rely on additional reference data, isolated components for distribution or domain predictions, leading to high training costs, increased inference complexity, and limited improvement potential for pre-trained models. To address these challenges, we first comprehensively analyze the effects of parameter update locations and ranks on downstream adaptation and knowledge retention. Based on these insights, we propose Dynamic Rank-Selective Low Rank Adaptation (LoRA), a universal and efficient CL approach that adaptively assigns ranks to LoRA modules based on their relevance to the current data. Unlike prior methods, our approach continually enhances the pre-trained VLM by retaining both the pre-trained knowledge and the knowledge acquired during CL. Our approach eliminates the need for explicit domain or distribution prediction and additional reference data, enabling seamless integration of new tasks while preserving pre-trained capabilities. It also maintains the original architecture and deployment pipeline of the pre-trained model without incurring any additional inference overhead. Extensive experiments and analyses demonstrate that our method outperforms state-of-the-art approaches in continually absorbing knowledge of downstream tasks while retaining pre-trained knowledge.
1 Introduction
Figure 1: Illustrations of approaches for CL of VLMs. (a) Continual full fine-tuning is applied toward a universal CL updating. Additional reference data is used to alleviate the forgetting of pre-trained knowledge (zero-shot prediction on unseen data), e.g., [62]. It is computationally expensive and suffers from forgetting. (b) New task-specific components isolated to the pre-trained model are added in CL, e.g., [57, 54]. They require task/domain prediction/gating in inference, increasing inference complexity and restricting the pre-trained model’s improvement potential with limited usage scenarios. (c) We aim to continually learn the model by memory-efficient updating of all parameters in a universal manner, similar to (a), but without requiring reference data. It allows for enhancing the pre-trained model while learning continuously from the data stream, surpassing the limitations of (b). By utilizing dynamic rank-selective LoRA, we reduce forgetting of both the knowledge in the pre-trained model and the knowledge acquired during continual learning.
Continual Learning (CL) [16, 9, 47, 25, 13] focuses on incrementally learning from new data streams without catastrophic forgetting [36, 32] or the need to retrain from scratch. With the advances of large-scale pre-trained models (PTMs), recent CL advances are increasingly leveraging these models. While many PTM-based CL approaches address continual adaptation for downstream tasks (e.g., class-incremental learning [51, 46, 20]), with an emphasis on downstream tasks’ performance, there is growing interest in preserving the inherent capabilities of PTMs, such as multi-domain continual learning for Vision-Language Models (VLMs) [62, 57, 54], or in continual pre-training to enhance PTMs over time [41]. In this paper, we focus on multi-domain continual learning for VLMs like CLIP [39].
The ZSCL [62] framework aims to continually improve pre-trained model by incrementally feeding it with new knowledge. However, this often results in a degradation of both pre-trained and previously acquired knowledge during continual learning steps. To address this challenge, prior works have explored strategies like using reference data [62] or memory replay [25, 51, 40] to distill knowledge during full fine-tuning. However, these methods incur high training costs due to the need for additional data and extensive retraining (Fig. 1(a)), and yet they still suffered from the forgetting problem. Other approaches [57, 54] introduce task-specific components isolated from the pre-trained model during the CL stream, which require task/domain prediction or gating at inference. This increases inference complexity and limits the model’s adaptability to unseen data, restricting its generalization potential (Fig. 1(b)).
In this work, we explore a more general and practical form of CL for VLMs (similar to [62]), which maintains strong performance on seen data with continually updated pre-trained model, and retains or even boosting the performance on unseen data. To ensure memory-efficient learning, we utilize low-rank parameter updates (LoRA) [19]. Rather than inserting LoRA modules at fixed locations [26, 15], we extend them to all pre-trained weight matrices to enhance adaptation, and crucially, to analyze the impact of each weight matrix on downstream learning and knowledge retention.
Our analysis (Sec. 3.2) reveals that not all parameter updates contribute equally to downstream adaptation, with some potentially harming pre-trained capabilities. Manually selecting update locations and ranks can significantly improve the balance between adapting to new tasks and retaining pre-trained knowledge. This insight motivates our proposed method, Dynamic Rank-Selective LoRA, which adaptively determines both the locations within the pre-trained weight matrices and the ranks of parameter updates.
We leverage a reformulation of parameter updates using Singular Value Decomposition (SVD) [60, 12, 28, 35, 59]. Our method introduces learnable importance weights that mimic the singular values in SVD, capturing the significance of each rank in the parameter update. These learnable weights enable the model to dynamically adjust the contribution of each rank during training. To ensure only the most important updates are integrated into the pre-trained weights, we optimize the importance weights using a sparse prompting approach with a soft-thresholding operation. Our contributions are fourfold:
  1. 1.
    In this paper, we explore the universal CL of VLMs in an memory-efficient manner, maintaining strong performance on both seen and unseen data with reduced forgetting. We perform dynamic rank selective parameter updates adaptively to all pre-trained weight matrices.
  2. 2.
    Through comprehensive analyses, we show that the location and rank of parameter updates are critical for achieving an good balance between downstream task adaptation and pre-trained knowledge retention.
  3. 3.
    We propose a general low-rank parameter update method that adaptively perform dynamic rank selective updates, effectively integrating new knowledge while maximizing the retention of pre-trained knowledge.
  4. 4.
    Extensive experiments and in-depth analyses validate our approach across a wide range of benchmarks for downstream adaptation and zero-shot transfer. Additionally, we evaluate on hold-out data to further evaluate knowledge retention.
2 Related Works
Continual Learning. Continual Learning enables models to sequentially learn new tasks while retaining previously acquired knowledge. Experience replay (ER) methods [30, 3, 6, 27, 7, 55, 56] store subsets of past data to refresh the model on prior tasks during new training. Parameter regularization approaches [22, 1, 58, 2, 20] constrain updates to important weights to preserve past knowledge. Meanwhile, dynamic networks [44, 45, 64, 46, 34, 26, 50, 49, 42, 48] adjust their architecture on the fly and balance learning new information while retaining old knowledge.
Multi-Domain Continual Learning of VLMs. Continual learning of VLMs [21, 61] aims to enable VLMs to sequentially learn across diverse domains while retaining their pre-trained generalization capabilities for previously seen tasks. The approach introduced in ZSCL [62] addresses the challenge of preserving zero-shot capabilities while adapting to new tasks, mitigating the risk of catastrophic forgetting. Subsequent works [57, 43, 54] have focused primarily on continual adaptation to downstream tasks while leveraging pre-trained predictions for unseen data. Recently, [54] proposed a more challenging evaluation setting, where the labels from different domains are mixed during testing.
Low-Rank Adaptation. Low-Rank Adaptation (LoRA) [19] has gained popularity as a parameter-efficient method for fine-tuning large pre-trained models. Recent works enhance LoRA by reformulating parameter updates using SVD, focusing on initialization from the SVD of pre-trained weights [59, 35], treating each rank independently [12, 60, 28], or employing a mixture of ranks [52]. These advancements highlight LoRA’s effectiveness in efficiently fine-tuning large models while preserving generalization.
3 Methodology
3.1 Preliminaries
Continual Learning of VLMs. We focus on continual learning for VLMs within two challenging settings: multi-domain task-incremental learning (MTIL) [62, 57] and cross-domain task-agnostic incremental learning (X-TAIL) [54]. In the MTIL setting, the model sequentially learns from a set of $T$ tasks. For each task $t$ , the dataset is represented as $\mathcal{D}^{t}=\{({\mathbf{x}}^{t}_{i},y^{t}_{i})\}_{i=1}^{N^{t}}$, where ${\mathbf{x}}^{t}_{i}\in\mathbb{R}^{H\times W\times C}$ is the input image, $y^{t}_{i}\in\mathcal{C}^{t}$ is its corresponding class label, and $N^{t}$ is the number of samples in the task. The category set $\mathcal{C}^{t}=\{y^{t}_{j}\}_{j=1}^{M^{t}}$ consists of the class names within the task $t$, containing a total of $M^{t}$ classes.
In the X-TAIL setting, however, the category set during testing encompasses both previously seen and unseen domains, denoted as $\mathcal{C}=\mathcal{C}_{\text{seen}}\cup\mathcal{C}_{\text{unseen}}$. Here, $\mathcal{C}_{\text{seen}}=\bigcup_{i=1}^{n}\mathcal{C}_{i}$ includes all classes learned during prior tasks, while $\mathcal{C}_{\text{unseen}}$ represents novel classes that were not encountered during training. This introduces a more challenging scenario, as the model must classify test images across a diverse range of domains, making it more complex than the MTIL setting.
Transformers. A typical transformer model is composed of stacked blocks, where each block consists of two main submodules: an attention (Attn) module utilizing multi-head attention (MHA) and a multi-layer perceptron (MLP). In each transformer layer, the attention module at the $l$-th transformer layer can be defined as: $\text{Attn}(x)=\text{MHA}(x\mathbf{W}^{\text{Q}}_{l},x\mathbf{W}^{\text{K}}_{l},x\mathbf{W}^{\text{V}}_{l})\mathbf{W}^{\text{O}}_{l},$ where $\text{MHA}(\cdot)$ denotes the multi-head attention operation. The input sequence is denoted as $x\in\mathbb{R}^{n\times d}$, where $n$ is the sequence length and $d$ is the hidden dimension. The matrices $\mathbf{W}^{\text{Q}}_{l},\mathbf{W}^{\text{K}}_{l},\mathbf{W}^{\text{V}}_{l},\mathbf{W}^{\text{O}}_{l}\in\mathbb{R}^{d\times d}$ correspond to the Query, Key, Value, and Output projection matrices, respectively.
The MLP module within each transformer block consists of two linear transformations separated by an activation function $\text{Act}(\cdot)$, such as ReLU or GELU [18]. The MLP module at the $l$-th transformer layer is defined as: $\text{MLP}(x)=\text{Act}(x\mathbf{W}^{\text{FC}}_{l}+b^{\text{FC}}_{l})\mathbf{W}^{\text{Proj}}_{l}+b^{\text{Proj}}_{l},$ where $\mathbf{W}^{\text{FC}}_{l}\in\mathbb{R}^{d\times d_{m}}$ and $\mathbf{W}^{\text{Proj}}_{l}\in\mathbb{R}^{d_{m}\times d}$ are the weight matrices, with $d_{m}$ being the hidden dimension of the MLP.
Vision-Language Models (VLMs). In vision-language models like CLIP [39], the image encoder $f_{\theta}$ and text encoder $g_{\psi}$ are implemented as transformers in our case, respectively. Given an input image ${\mathbf{x}}\in\mathbb{R}^{H\times W\times C}$, it is segmented into patches and processed by the vision encoder to produce latent visual features ${\mathbf{z}}^{\text{V}}=f_{\theta}({\mathbf{x}})\in\mathbb{R}^{D}$. For text, the class label $y$ is embedded into prompts like “a photo of a [CLS]”, where [CLS] is the class name, resulting in text input ${\mathbf{t}}$. This prompt is encoded by the text encoder, yielding ${\mathbf{z}}^{\text{T}}=g_{\psi}({\mathbf{t}})\in\mathbb{R}^{D}$. During inference, the probability of classifying image ${\mathbf{x}}$ into class $y_{i}\in\{1,\dots,C\}$ is computed as: $p(y_{i}|{\mathbf{x}})=\frac{\exp(\mathtt{sim}({\mathbf{z}}^{\text{V}},{\mathbf{z}}^{\text{T}}{y_{i}})/\tau)}{\sum_{c=1}^{C}\exp(\mathtt{sim}({\mathbf{z}}^{\text{V}},{\mathbf{z}}^{\text{T}}_{y_{c}})/\tau)},$ where $\mathtt{sim}(\cdot)$ is cosine similarity, and $\tau$ is the temperature.
Figure 2: Evaluations of adaptation to downstream tasks and retention of zero-shot capabilities after training CLIP with various LoRA insertion locations and ranks.
Figure 3: Analysis of the effects of removing trained LoRA modules on downstream task adaptation (Aircraft) and retention of pre-trained capabilities (ImageNet-1k).
Figure 4: Overview of our proposed method: we propose the dynamic rank-selection LoRA that allows the each pre-trained weight matrices to adaptively insert the necessary ranks for downstream adaptation while retaining pre-trained capabilities. After training on each task, the dynamic ranks of the parameter updates are merged back into the pre-trained weights, without incurring additional inference overhead.
3.2 Analyses on LoRA Location and Rank
Analyses of locations and ranks of low-rank updates. We begin by investigating the influence of performing parameter updates at different locations within the pre-trained model, varying the update strength to assess its effects on both downstream adaptation and retention of pre-trained capabilities. The results of this analysis are presented in Fig. 3. We leverage Low-Rank Adaptation (LoRA) [19], where low-rank parameter updates are directly applied to the pre-trained weight matrices.
Our findings reveal the following: (1) Updating only the vision encoder leads to catastrophic forgetting of the pre-trained capabilities, significantly diminishing zero-shot performance. (2) Updating only the attention modules generally reduces forgetting but restricts the model’s ability to adapt to new tasks. (3) Using a higher rank enhances adaptation to downstream tasks without necessarily increasing the forgetting of pre-trained knowledge.
These observations highlight the need for a strategic selection of insertion locations and ranks to achieve an better balance between adaptation and zero-shot retention.
Analyses of pruning subsets of low-rank updates. To further investigate, we analyzed the effects of selectively perturbing individual modules of the pre-trained weights, as shown in Fig. 3. Initially, we inserted LoRA into all pre-trained matrices with a rank of 16, corresponding to the “Vision + Text, Attn + MLP, rank 16” configuration shown in Fig. 3. We then iteratively remove the LoRA updates learned for each module, examining the impact on both downstream adaptation and the knowledge retention.
Our findings reveal that not all parameter updates are equally critical for downstream adaptation, and some may even degrade the model’s pre-trained capabilities. For example, removing updates in the Value projection of the Attention module within the vision encoder had minimal impact on downstream adaptation but significantly restored the model’s pre-trained zero-shot performance. Similarly, removing updates in the MLP layer of the vision encoder slightly reduced downstream task performance but, interestingly, enhanced the model’s pre-trained capabilities, even surpassing its original zero-shot performance.
3.3 CL VLMs with Dynamic Rank-Selective LoRA
Overview. As shown in Fig. 4, we introduce dynamic rank-selective parameter updates for all pre-trained weight matrices, adaptively inserting updates based on the importance of each rank for current training data of task $t$. Specifically, we integrate dynamic rank-selective LoRA into all pre-trained weights within the Attention and MLP modules of both the vision and text encoders. For clarity, for learning on task $t$, we denote the pre-trained weights or those updated from the previous task as $\{\mathbf{W}^{t,m}_{0}\}_{m=1}^{M}$, where $M$ represents the total number of pre-trained weight matrices added with LoRA. After training on each task $t$, these dynamic rank updates are merged back into the original pre-trained weights.
    Method Aircraft [31] Caltech101 [14] DTD [8] EuroSAT [17] Flowers [37] Food [5] MNIST [11] OxfordPet [38] Cars [23] SUN397 [53] Average
CLIP
   Zero-shot 23.5 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 58.5
Transfer
   Zero-shot [39] 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
   LwF [25] 66.6 26.9 19.5 51.0 78.4 26.6 68.9 35.5 56.1 47.7
   WiSE-FT [51] 70.1 31.9 25.3 56.3 79.8 29.9 74.9 45.6 56.8 52.3
   iCaRL [40] 71.7 35.0 43.0 63.4 86.9 43.9 87.8 63.7 60.0 61.7
   ZSCL [62] 73.3 32.6 36.8 62.1 83.8 42.1 83.6 56.5 60.2 59.0
[…]
   Ours 74.3 36.8 44.2 69.9 83.5 42.8 88.9 64.6 63.4 63.2
[…]
Average
   LwF [25] 24.7 79.7 38.3 36.9 63.9 81.0 36.5 71.9 42.7 56.7 53.2
   WiSE-FT [51] 27.1 76.5 40.9 31.3 68.7 81.6 31.4 74.7 51.7 58.4 54.2
   iCaRL [40] 25.4 72.1 37.5 51.6 65.1 87.1 59.1 88.0 63.7 60.1 61.0
   ZSCL [62] 36.0 75.0 40.7 40.5 71.0 85.3 46.3 83.3 60.7 61.5 60.0
[…]
   Ours 41.4 81.0 58.6 77.8 83.4 84.5 64.5 90.3 67.2 64.4 71.3
[…]
Last
   LwF [25] 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
   WiSE-FT [51] 21.8 76.8 42.9 20.8 77.5 84.9 30.7 76.6 75.8 72.5 58.0
   iCaRL [40] 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
   ZSCL [62] 33.1 75.3 43.5 35.2 74.6 87.4 50.4 84.2 77.3 73.4 63.4
[…]
   Ours 37.7 81.5 65.1 89.9 91.4 85.5 96.8 93.3 77.3 73.5 79.2
[…]
Table 1: Comparison of different CL methods on X-TAIL for each domain in terms of “Transfer”, “Average”, and “Last” scores (%). The best and the second best results are highlighted in red and blue, respectively. Methods marked with $\dagger$ indicate the use of domain prediction or distribution detection techniques, as in [57, 54].
Rank-selective LoRA with importance weights. LoRA [19] is a parameter-efficient fine-tuning technique designed to adapt pre-trained models using a minimal number of additional parameters. Instead of directly updating a pre-trained weight matrix $\mathbf{W}_{0}^{t,m}\in\mathbb{R}^{d\times k}$, LoRA introduces two low-rank matrices, $\mathbf{B}^{t,m}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}^{t,m}\in\mathbb{R}^{r\times k}$, such that $\Delta\mathbf{W}^{t,m}=\mathbf{B}^{t,m}\mathbf{A}^{t,m}$, where $r\ll\min(d,k)$. The updated weight matrix is defined as:
$\mathbf{W}^{t,m}=\mathbf{W}_{0}^{t,m}+\Delta\mathbf{W}^{t,m}=\mathbf{W}_{0}^{t,m}+\mathbf{B}^{t,m}\mathbf{A}^{t,m}.$ (1)
In this formulation, the original weights $\mathbf{W}_{0}^{t,m}$ remain fixed, and only the matrices $\mathbf{B}^{t,m}$ and $\mathbf{A}^{t,m}$ are trained, significantly reducing the number of trainable parameters to $r(d+k)$. Notably, during inference, LoRA incurs no additional computational overhead compared to the original pre-trained model, as the low-rank updates can be merged into the pre-trained weights, as shown in Eq. (1).
Previous works reformulate the low-rank parameter update in the form of Singular Value Decomposition (SVD) [60, 12, 28, 35, 59]. We introduce dynamic importance weights into the formulation:
$\Delta\mathbf{W}^{t,m}=\sum_{i=1}^{r}\mathbf{w}_{i}^{t,m}\mathbf{B}_{i}^{t,m}\mathbf{A}_{i}^{t,m},$ (2)
where $\mathbf{w}^{t,m}\in\mathbb{R}^{r}$ is a learnable vector that represents the importance of each rank during training. The importance weights imitate the diagonal elements in the singular value matrix of SVD and we allow these weights to be learned directly from the data.
The introduction of these importance weights incurs only a minimal increase in trainable parameters, adding just $r$ additional parameters per LoRA module. For initialization, we adhere to the convention that randomly initialize the matrix $\mathbf{A}$, set the matrix $\mathbf{B}$ to zero, and we initialize the importance weights $\mathbf{w}^{t,m}$ randomly.
Dynamic rank selection with sparse importance. The formulation in Eq. (2) enables the model to dynamically adjust the importance of each rank through gradient descent. [^1] To align the training process with the pre-trained model, we use CLIP’s contrastive loss during pre-training as the supervised training loss. However, not all ranks contribute equally to downstream task performance (as shown in Fig. 3). To ensure that only the most impactful LoRA ranks are retained in the pre-trained weights, we propose to optimize the learning of importance weights in a sparse-promoting way by minimizing the $\ell_{1}$ norm of the importance weights. The optimization objective for paramters $\{\mathbf{w}^{t,m},\mathbf{B}^{t,m},\mathbf{A}^{t,m}\}_{m=1}^{M}$ learned on task $t$ are as follows:
$\mathcal{L}_{\text{train}}^{t}:=\mathcal{L}_{\text{sup}}^{t}+\lambda\sum_{m=1}^{M}\|\mathbf{w}^{t,m}\|_{1},$ (3)
where $\mathcal{L}_{\text{sup}}^{t}$ is the supervised training loss, $M$ denotes the total number of pre-trained weight matrices inserted with LoRA of rank $r$, and $\mathbf{w}^{t,m}$ represents the importance weights of the each rank of the LoRA module added to the pre-trained weight matrix $\mathbf{W}^{t,m}_{0}$, and $\lambda$ controls the strength of the $\ell_{1}$ regularization.
We adopt the proximal gradient [4] for handling the non-differentiable $\ell_{1}$ regularization applied to the importance weights. Specifically, the dynamic update of the importance weight for the $i$-th rank of LoRA can be written as a soft-thresholding operation:
$\mathbf{w}_{i}^{t,m}:=\mathbbm{1}(\lvert\hat{\mathbf{w}}^{t,m}_{i}\rvert>\kappa)\cdot(\hat{\mathbf{w}}^{t,m}_{i}+\mathrm{sign}(\hat{\mathbf{w}}^{t,m}_{i})\cdot\kappa),$ (4)
where $\hat{\mathbf{w}}^{t,m}_{i}$ denotes the value of ${\mathbf{w}^{t,m}_{i}}$ after gradient update using only the supervised loss $\mathcal{L}_{\text{sup}}^{t}$. The threshold parameter $\kappa$ is initially set to zero and gradually increases until it reaches a maximum value $\kappa_{\text{max}}$. The indicator function $\mathbbm{1}(\cdot)$ returns 1 if condition is met, 0 otherwise, and $\mathrm{sign}(\cdot)$ represent the +/+/-+ / - sign of input.
This encourages that only the ranks with significant importance are preserved, while those with low importance are pruned. By adjusting the weightings iteratively, the model focuses dynamic parameter updates on the most relevant components. In the early stages of training, dense updates are performed without Eq. (4) to allow all ranks to capture task-relevant information.
Continually enhance the pre-trained model. After training is complete, ranks with importance weight of zero are removed, while non-zero ranks are merged into the pre-trained weights by substituting Eq. (2) into the original LoRA formulation Eq. (1). When a new task is introduced, a new set of LoRA modules is initialized, and the computations involve only the newly initialized LoRA modules in conjunction with the continually updated pre-trained model. This approach allows the model to leverage accumulated knowledge efficiently while keeping the computational footprint minimal without additional overhead.
    Method Aircraft [31] Caltech101 [14] CIFAR100 [24] DTD [8] EuroSAT [17] Flowers [37] Food [5] MNIST [11] OxfordPet [38] Cars [23] SUN397 [53] Average
CLIP
   Zero-shot [39] 24.3 88.4 68.2 44.6 54.9 71.0 88.5 59.4 89.0 64.7 65.2 65.3
Transfer
   Zero-shot [39] 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
   LwF [25] 72.1 49.2 35.9 44.5 41.1 66.6 50.5 69.0 19.0 51.7 50.0
   LwF-VR [13] 82.2 62.5 40.1 40.1 56.3 80.0 60.9 77.6 40.5 60.8 60.1
   WiSE-FT [51] 77.6 60.0 41.3 39.4 53.0 76.6 58.1 75.5 37.3 58.2 57.7
   ZSCL [62] 84.0 68.1 44.8 46.8 63.6 84.9 61.4 81.4 55.5 62.2 65.3
[…]
   Ours 92.4 68.4 45.8 54.5 69.6 87.4 65.2 88.5 64.2 64.5 69.9
[…]
Average
   LwF [25] 23.5 77.4 43.5 41.7 43.5 52.2 54.6 63.4 68.0 21.3 52.6 49.2
   LwF-VR [13] 24.9 89.1 64.2 53.4 54.3 70.8 79.2 66.5 79.2 44.1 61.6 62.5
   WiSE-FT [51] 32.0 87.7 61.0 55.8 68.1 69.3 76.8 71.5 77.6 42.0 59.3 63.7
   ZSCL [62] 28.2 88.6 66.5 53.5 56.3 73.4 83.1 56.4 82.4 57.5 62.9 64.4
[…]
   Ours 34.6 95.8 73.9 60.0 77.1 81.3 86.6 75.9 89.9 66.1 65.3 73.3
[…]
Last
   LwF [25] 22.1 58.2 17.9 32.1 28.1 66.7 46.0 84.3 64.1 31.5 60.1 46.5
   LwF-VR [13] 22.9 89.8 59.3 57.1 57.6 79.2 78.3 77.7 83.6 60.1 69.8 66.9
   WiSE-FT [51] 30.8 88.9 59.6 60.3 80.9 81.7 77.1 94.9 83.2 62.8 70.0 71.9
   ZSCL [62] 26.8 88.5 63.7 55.7 60.2 82.1 82.6 58.6 85.9 66.7 70.4 67.4
[…]
   Ours 31.6 95.5 72.8 63.5 85.0 89.7 85.0 94.7 93.2 73.6 73.0 78.0
[…]
Table 2: Comparison with state-of-the-art methods on 5-shot MTIL setting in terms of “Transfer”, “Average”, and “Last” scores (%). The best and the second best results are highlighted in red and blue, respectively. Methods marked with $\dagger$ indicate the use of domain prediction or distribution detection techniques, as in [57, 54].
4 Experiments
4.1 Experimental Setting
Datasets. We evaluate our method on two widely used continual learning settings for vision-language models: MTIL [62, 57] and X-TAIL [54]. For the MTIL setting, a total of 11 datasets are used, where each dataset is treated as an individual task to be incrementally learned. For the X-TAIL setting, CIFAR100 is excluded to prevent domain overlap, following the protocol in [54]. In line with [54], we use a 5-shot split for MTIL and a 16-shot split for X-TAIL.
Evaluation Metrics. We adopt the same evaluation metrics used in prior works, which include “Transfer”, “Average”, and “Last”. The “Transfer” metric evaluates the model’s zero-shot transfer capability on previously unseen data. The “Last” metric assesses the model’s ability to retain knowledge from earlier tasks. The “Average” metric is a composite score that reflects the mean performance across both the “Transfer” and “Last” evaluations.
Implementation Details. We follow the experimental setups in [62, 57, 54] and utilize the CLIP model with a ViT-B/16 backbone [39] for all experiments. By default, our proposed importance-aware sparse LoRA is applied to all pre-trained weight matrices across both the vision encoder and text encoder, with an initial rank set to 16. Each task in both experimental settings is trained over 500 iterations using the AdamW optimizer [29]. The dense training iteration ratio is set to 0.5 for X-TAIL and 0.7 for MTIL, relative to the total training iterations. The maximum threshold value for adaptive pruning is fixed at 0.005 for both settings.
Method Average Transfer Accuracy on Unseen Data
CIFAR100 Places365 ImageNet-1k Average
Zero-Shot [39] 68.24 33.77 66.72 56.24
MoE-Adapter [57] 68.24 33.77 66.72 56.24
RAIL [54] 68.24 33.77 66.72 56.24
Ours 68.95 36.52 68.52 58.00
Table 3: Evaluation of average transfer accuracy on datasets that were never encountered during the continual learning stages.
Methods
Training
Params.
Additional Infer.
Params. / Mem.
LWF [25] 129.6M None
ZSCL [62] 129.6M None
MoE-Adapters [57] 59.8M 13.35M
RAIL [54] N/A 24.18M / 9.01M
Ours 4.4M None
Table 4: Comparison on computation cost with previous methods.
4.2 Experimental Results
Cross-domain task-agnostic incremental learning. We report the evaluation results on the X-TAIL setting in Table 1. Unlike previous works that rely on model parameter updates [25, 51, 40, 13] along with additional memory replay or reference data, our approach performs parameter-efficient updates without requiring any memory replay or reference data. This results in significant improvements across all evaluated metrics.
In contrast to prior methods [57, 54] [^2] Dual version of [54] stores all the training samples’ feature embeddings in the memory, for a fair comparison, we excluded the results here. We leave more detailed comparisons in Appendix., which use zero-shot predictions to preserve pre-trained capabilities and prevent forgetting, our method overcomes this limitation by continually refining the model. Notably, while these approaches are constrained by the upper limit of zero-shot performance, our approach surpasses this limit, achieving a higher average transfer accuracy. Although some predictions on unseen data might be slightly impacted due to the nature of continual parameter updates, our method ultimately achieves a higher overall transfer accuracy by effectively adapting the pre-trained model to new tasks while also enhancing its generalization capabilities.
For both the Average and Last metrics, our method demonstrates superior or comparable performance without relying on domain ID prediction [57], high-dimensional projections, or storing a memory bank of previously seen sample features [54]. Instead, our method focuses purely on updating the pre-trained model, yet still delivers competitive results.
Furthermore, since our approach is entirely orthogonal to the additional techniques used in previous methods [57, 54], these techniques can be seamlessly integrated into our framework to further enhance performance, benefited from our ultimate goal of continually enhancing the pre-trained model’s capabilities.
Multi-domain task-incremental learning. We further evaluate our proposed method under the few-shot MTIL setting, as shown in Table 2, following [57, 54]. Consistent with the results observed in the X-TAIL setting, our method demonstrates clear superiority in this setting. Without employing any additional techniques, our approach outperforms previous methods across all evaluation metrics, achieving an average Transfer accuracy that surpasses the upper bound set by zero-shot predictions in prior works. Moreover, by supplementing our method with techniques used in prior approaches [57, 54], we can further improve the performance.
4.3 Discussions
Figure 5: Visualization and statistical analysis of rank activation on the Aircraft dataset using our proposed method.
Figure 6: Visualization and statistical analysis of rank activation on the Oxford Pets dataset using our proposed method.
Pre-trained knowledge retention and potentially improved pre-trained model. To understand how our method improves transfer performance beyond the upperbound set by zero-shot predictions in previous works, we evaluate on datasets that were never encountered during the continual learning steps, specifically CIFAR100 [24], Places365 [63], and ImageNet-1k [10], as shown in Table 3.
Unlike prior approaches [57, 54] that rely exclusively on zero-shot predictions for unseen data, thereby limiting their performance to that of the original CLIP model, our method goes beyond this limitation. By effectively integrating new knowledge into the pre-trained parameter space without sacrificing existing capabilities, our approach enhances the model’s overall performance. Consequently, our method achieves superior transfer accuracy, surpassing the theoretical upper bound defined by zero-shot performance in previous methods. This not only demonstrates improved adaptation but also enhances the model’s generalization capabilities on completely unseen datasets during CL.
Rank allocation results on different datasets. We visualize and analyze the distribution of activated ranks merged into the pre-trained weights across different datasets in Fig. 6 and Fig. 6. Overall, it is evident that the Aircraft dataset generally requires more ranks to be updated in the pre-trained model compared to the Pets dataset.
We observe that for the Aircraft dataset, a greater number of ranks are allocated to the Output projection of the Attention module within the text encoder (Fig. 6(b)). Additionally, a higher concentration of ranks is assigned to the deeper layers, particularly layers 6 to 10 (Fig. 6(c)).
In contrast, the analysis of the Pets dataset (Fig. 6(b) and Fig. 6(c)) reveals that more ranks are allocated to the Attentions’ Key and Value projection, and MLP’s FC layer of vision encoder. Furthermore, the majority of rank allocations are in the first and last layers, suggesting a different pattern for promoting knowledge retention and adaptation across datasets. These findings highlight the adaptive nature of our approach, where rank allocation varies based on the characteristics of the dataset, optimizing the model’s ability to capture domain-specific information effectively.
Computation Cost. We compare the computational costs in terms of trainable parameters and additional inference requirements, as summarized in Fig. 4. For MoE-Adapter, the difference between training and testing parameters stems from its mixture-of-experts design: during training, all experts are active, while at inference, only the top-2 experts are selected, reducing the active parameter count. RAIL-Primal incurs additional computational overhead during inference due to its use of high-dimensional feature expansion through projection and regression layers. RAIL-Dual maintains a memory bank of image features for all training samples across 1,100 classes, which significantly increases memory usage.
Notably, these methods experience linearly growing computational costs as the number of tasks increases. In contrast, our method minimizes training overhead by selectively retaining only ranks with high importance. Once training is complete, these updates are merged into the pre-trained weights, resulting in zero additional memory overhead during inference.
Figure 7: Ablation studies on different insertion locations and the impact of varying the initial rank.
Figure 8: Analyses of the effects of (a) varying the maximum pruning threshold and (b) different ratios of dense training iterations.
4.4 Ablation Studies
Effects of insertion locations and number of initial ranks. We examine the impact of different insertion locations and initial ranks for our method, with results presented in Fig. 7. These findings are consistent with our analytical experiments shown in Fig. 3. Our results indicate that when training is restricted to only the vision encoder, transfer accuracy is significantly reduced, suggesting that the pre-trained capabilities are negatively impacted. Conversely, increasing the maximum rank improves adaptation ability without necessarily reducing transfer accuracy. Furthermore, when limiting insertion to only the Attention modules, we observe that transfer accuracy can surpass our previously reported state-of-the-art results in Table 1 and Table 2. However, this comes at the expense of reduced adaptation capabilities. To achieve an better balance between adaptation and retention, we enable insertion into all modules by default. This configuration ensures a better trade-off between adapting to new tasks while preserving the model’s zero-shot transfer capabilities.
Effect of maximum threshold value and ratio for dense iterations. We analyze the effects of varying the maximum threshold value $\kappa_{\text{max}}$ and the ratio for dense training iterations in Fig. 8. The results indicate that as the threshold increases, the Transfer accuracy shows an increasing trend. This suggests that pruning more ranks can enhance the model’s performance on unseen data (i.e., improve transfer capability), but it also limits the model’s ability to adapt to downstream tasks, as evidenced by changes in the Last accuracy metric. Our experiments show that a threshold value around 0.005 strikes a good balance between retaining pre-trained knowledge and adapting to new tasks.
The ratio of dense training iterations determines when the model begins the adaptive pruning process. Setting this ratio to 0 (i.e., starting pruning from the beginning of training) severely hinders adaptation, as LoRA lacks sufficient time to learn essential task-related knowledge before pruning begins. On the other hand, delaying pruning for too long can result in the model forgetting pre-trained capabilities, thereby reducing transfer accuracy. Our findings indicate that setting the ratio around 0.5 to 0.7 works best across all experimental settings, providing a good balance between adaptation and knowledge retention.
5 Conclusion
In this paper, we introduced a novel Dynamic Rank-Selective LoRA approach to address the challenges of continual learning for vision-language models. By dynamically adjusting the ranks based on each ranks’ importance, our method optimizes knowledge retention while enabling efficient adaptation to new tasks without adding inference overhead. Extensive evaluations on multi-domain and cross-domain settings demonstrated that our method achieves state-of-the-art performance, effectively balancing downstream adaptation and zero-shot generalization.
Limitations and future works. Our current approach directly merges parameter updates into the pre-trained weights without explicitly considering correlations between tasks. Exploring methods to leverage inter-task relationships and shared information build a promising direction for future research.
Acknowledgment
This work was partially supported by the ARC DECRA Fellowship (DE230101591) awarded to D. Gong. H. Lu is affiliated with CSIRO Data61 through a PhD scholarship and acknowledges the support of the Google PhD Fellowship.
Supplementary Material
Appendix A Additional Details of the Proposed Method
In Sec. 3.3 of the main paper, we introduce our method for learning dynamic rank-selective parameter updates to pre-trained weight matrices. We leverage a learnable vector to represent the importance of each rank, which is updated through a soft-thresholding operation. In this section, we provide more details to enhance clarity and understanding.
Reproducibility Statement. The source code will be made publicly available upon acceptance of the paper.
A.1 Full Derivation of Eq. (4)
For simplicity, we omit the notation for the training task $t$ in the following derivations. With a slight abuse of notation, we use subscript $t$ to represent $t$-th training iteration. At the $t$-th training iteration, the training loss with $\ell_{1}$ sparse regularization, as defined in the main paper, is:
$\mathcal{L}_{\text{train}}(\Delta_{t}):=\mathcal{L}_{\text{sup}}(\Delta_{t})+\lambda\sum_{m=1}^{M}\|\mathbf{w}^{m}_{t}\|_{1},$ (3)
where $\mathcal{L}_{\text{sup}}$ is the supervised training loss, $\Delta_{t}$ is the all trainable parameters $\{\mathbf{w}^{m}_{t},\mathbf{B}^{m}_{t},\mathbf{A}^{m}_{t}\}_{m=1}^{M}$ at $t$-th iteration, $M$ is the total number of pre-trained weight matrices with LoRA of rank $r$, $\mathbf{w}^{m}_{t}$ represents the importance weights of the $m$-th LoRA module, and $\lambda>0$ controls the strength of the $\ell_{1}$ regularization.
We apply proximal gradient descent [4] to handle the non-differentiable $\ell_{1}$ regularization of the importance weights [12]. At the $t$-th iteration, the update rule for the importance weights of each LoRA becomes:
$\mathbf{w}_{t+1}^{m}$ $\leftarrow\arg\min_{\mathbf{w}^{m}}\,\eta_{t}\cdot\lambda\|\mathbf{w}^{m}\|_{1}$
$\quad+\frac{1}{2}\|\mathbf{w}^{m}-(\mathbf{w}_{t}^{m}-\eta_{t}\nabla_{\mathbf{w}^{m}}\mathcal{L}_{\text{sup}}(\Delta_{t}))\|_{2}^{2},$ (5)
where $\eta_{t}>0$ is the step-size of $t$-th iteration, and $\nabla_{\mathbf{w}^{m}}\mathcal{L}_{\text{sup}}(\Delta_{t})$ is the gradient of the supervised loss with respect to $\mathbf{w}^{m}$. This update balances the $\ell_{1}$ sparsity regularization term $\|\mathbf{w}^{m}\|_{1}$, with the quadratic penalty term $\frac{1}{2}\|\mathbf{w}^{m}-\hat{\mathbf{w}}^{m}_{t}\|_{2}^{2}$ , which ensures proximity to the gradient update if importance weights $\hat{\mathbf{w}}^{m}_{t}:=\mathbf{w}_{t}^{m}-\eta_{t}\nabla_{\mathbf{w}^{m}}\mathcal{L}_{\text{sup}}(\Delta_{t})$. This can be achieved through a soft-thresholding operator $\mathcal{T}$:
$\mathbf{w}_{t+1}^{m}\leftarrow\mathcal{T}_{\eta_{t}\cdot\lambda}(\mathbf{w}_{t}^{m}-\eta_{t}\nabla_{\mathbf{w}^{m}}\mathcal{L}_{\text{sup}}(\Delta_{t})),$ (6)
where the soft-thresholding operator $\mathcal{T}$ is defined as:
$\mathcal{T}_{\kappa}(x)=\mathbbm{1}(\lvert x\rvert>\kappa)\cdot(x+\mathrm{sign}(x)\cdot\kappa),$ (7)
such that the input to the operator $\mathcal{T}_{\kappa}$ is the elements in $\hat{\mathbf{w}}^{m}_{t}$, $\hat{\mathbf{w}}^{m}_{t}$ is the value of $\mathbf{w}^{m}_{t}$ after the gradient update using only $\mathcal{L}_{\text{sup}}(\Delta_{t})$, and $\kappa=\eta_{t}\cdot\lambda$ is the threshold. The indicator function $\mathbbm{1}(\cdot)$ returns 1 if the condition is met, and 0 otherwise, while $\mathrm{sign}(\cdot)$ denotes the sign (+/)(+/-)( + / - ) of the input. This concludes our derivation of the dynamic update for the importance weights as presented in the main paper Eq. (4).
Appendix B More Experimental Results and Ablations
B.1 Effects of Additional Adaptive Prediction Head on Visual Representations
Related continual learning methods focus on updating representations through the backbone model (e.g., [50, 62, 57], and ours) or by adding a representation alignment or adaptive prediction head [33, 54]. These two approaches can be considered orthogonal. We investigate how the representation updates introduced by our proposed method can complement the additional adaptive prediction head, utilizing the aggression-based adapters described in [54].
We provide a more detailed analysis of the effects of the additional prediction head on the visual representations in X-TAIL setting in Table 5. Note that the Dual version of RAIL [54] stores all visual representations of all samples for all seen classes and domains in memory, it is not included in Tables 1 and 2 for a fair comparison.
Since RAIL heavily relies on the frozen pre-trained model, the Transfer accuracy is limited by the zero-shot performance of the pre-trained model. In contrast, our approach gradually accumulates knowledge into the pre-trained model, enhancing its capability on unseen data and achieving improved Transfer accuracy. Moreover, by leveraging our continually enhanced pre-trained model, coupling our method with RAIL significantly outperforms the original RAIL, which uses a frozen pre-trained model, on the Average and Last metrics.
Transfer Average Last
RAIL-Primal 62.4 70.7 79.1
Ours + RAIL-Primal 63.1 74.1 83.4
RAIL-Dual 62.4 71.9 82.4
Ours + RAIL-Dual 63.1 74.2 83.6
Table 5: Experimental results for integrating aggression-based adapters in the X-TAIL setting. Metrics for “Transfer”, “Average”, and “Last” are reported as averages across all datasets.
B.2 Analysis of Changes/Interference w.r.t. Pre-Trained Weights
In Sec. 4.3 of the main paper, Fig. 6 and Fig. 6 visualize and analyze the dynamic ranks assigned to each module. Here, we extend the analysis by comparing parameter changes to pre-trained weights using the amplification factor as in [19]. The amplification factor is given by:
$\text{Amp}=\frac{\|\Delta\mathbf{W}\|_{F}}{\|\mathbf{U}^{T}\mathbf{W}\mathbf{V}^{T}\|_{F}},$ (8)
where $\mathbf{W}$ is the pre-trained weight matrix, and $\Delta\mathbf{W}$ is the parameter change introduced by LoRA. For vanilla LoRA and our method, $\Delta\mathbf{W}$ is computed using Eq. (1) and Eq. (2), respectively. $\mathbf{U}$ and $\mathbf{V}$ represent the top $r$ singular vectors of $\text{SVD}(\Delta\mathbf{W})$, where $r$ is the number of interested ranks in analysis.
A higher amplification factor indicates that low-rank parameter updates amplify directions less emphasized in the pre-trained weights [19]. We visualize the amplification factors for both our method and vanilla LoRA trained on the Aircraft, EuroSAT and Oxford Pets datasets in Fig. 9.
The results show that, compared to vanilla LoRA, our method tends to concentrate parameter updates on a more specific subset of transformer modules while reducing updates to pre-trained weights less relevant to the current data. This demonstrates that our method more effectively identifies critical pre-trained weight matrices relevant to the current data and applies targeted dynamic parameter updates. Additionally, by reducing the strength of parameter updates for pre-trained weights less related to the current data, our method minimizes the impact on pre-trained knowledge and knowledge acquired from previous tasks compared to vanilla LoRA.
Note that this analysis differs from the visualizations of number of activated ranks in Fig. 6 and Fig. 6. The number of activated ranks reflects the quantity of significant ranks contributing to parameter updates for leaning each task. In this analysis, the amplification factor directly measures the parameter change to each pre-trained weight matrix.
(a) Aircraft - LoRA
(b) EuroSAT - LoRA
(c) Pets - LoRA
(d) Aircraft - Ours
(e) EuroSAT - Ours
(f) Pets - Ours
Figure 9: Visualization of amplification factors when trained on (a, d) Aircraft, (b, e) EuroSAT, and (c, f) Pets, comparing (d, e, f) our method of dynamic sparse rank selective LoRA and (a, b, c) vanilla LoRA. It shows that the proposed method with adaptive rank selection can achieve more focused/concentrated updating and less interference.
B.3 More Visualizations and Analyses of Rank Allocation Across Datasets
In Sec. 4.3, Fig. 6 and Fig. 6, we visualize and analyze the dynamic ranks assigned to each module for the Aircraft and Oxford Pets datasets under the X-TAIL experimental settings. Here, we extend these visualizations and analyses to additional datasets, as shown from Fig. 11 to Fig. 17.
The visualizations, along with statistical analyses of transformer modules and layers, reveal distinct rank allocation patterns across datasets. These findings suggest that the rank of parameter updates needed for achieving a good downstream improvements and knowledge retention is distinct across each kind of data.
Figure 10: Visualization and statistical analysis of rank activation on the Caltech101 dataset using our proposed method.
Figure 11: Visualization and statistical analysis of rank activation on the DTD dataset using our proposed method.
Figure 12: Visualization and statistical analysis of rank activation on the EuroSAT dataset using our proposed method.
Figure 13: Visualization and statistical analysis of rank activation on the Flowers dataset using our proposed method.
Figure 14: Visualization and statistical analysis of rank activation on the Food101 dataset using our proposed method.
Figure 15: Visualization and statistical analysis of rank activation on the MNIST dataset using our proposed method.
Figure 16: Visualization and statistical analysis of rank activation on the Stanford Cars dataset using our proposed method.
Figure 17: Visualization and statistical analysis of rank activation on the SUN397 dataset using our proposed method.

### v4

@@
 Abstract
-We investigate whether the pre-trained knowledge of vision-language models (VLMs), such as CLIP, can be retained or even enhanced during continual learning (CL) while absorbing knowledge from a data stream.
-Existing methods often rely on additional reference data, isolated components for distribution or domain predictions, leading to high training costs, increased inference complexity, and limited improvement potential for pre-trained models.
-To address these challenges, we first comprehensively analyze the effects of parameter update locations and ranks on downstream adaptation and knowledge retention.
-Based on these insights, we propose Dynamic Rank-Selective Low Rank Adaptation (LoRA), a universal and efficient CL approach that adaptively assigns ranks to LoRA modules based on their relevance to the current data.
-Unlike prior methods, our approach continually enhances the pre-trained VLM by retaining both the pre-trained knowledge and the knowledge acquired during CL.
-Our approach eliminates the need for explicit domain or distribution prediction and additional reference data, enabling seamless integration of new tasks while preserving pre-trained capabilities.
-It also maintains the original architecture and deployment pipeline of the pre-trained model without incurring any additional inference overhead.
-Extensive experiments and analyses demonstrate that our method outperforms state-of-the-art approaches in continually absorbing knowledge of downstream tasks while retaining pre-trained knowledge.
+We investigate whether the pre-trained knowledge in vision-language models (VLMs), such as CLIP, can be retained—or even enhanced—in continual learning (CL) while incorporating new knowledge from the data stream.
+Existing CL methods primarily focus on continual downstream adaptation using components isolated from pre-trained model (PTM), increasing inference complexity and limiting improvements to the PTM itself; some also retain knowledge relying on additional reference data, leading to high training costs.
+To address these limitations, we propose a universal and efficient Continual Learning approach for VLM based on Dynamic Rank-Selective LoRA (CoDyRA), which directly improves the PTMs while preserving the existing knowledge from both pre-training and CL.
+Through analyses on how LoRA rank and placement impact/regularize the learning and forgetting in CL, we propose CoDyRA that adaptively performs rank-minimized parameter updates in different modules, based on their importance to the current data.
+This ensures a balance between knowledge acquisition (plasticity) and forgetting mitigation (stability).
+Our method operates without explicit domain or distribution prediction and does not rely on reference data, enabling seamless task integration while maintaining pre-trained capabilities.
+Moreover, CoDyRA preserves the original model architecture and deployment pipeline, introducing no additional inference overhead.
+Extensive experiments demonstrate that our approach enhances representations based on new downstream data while retaining pre-trained knowledge, achieving state-of-the-art results.
 1 Introduction
-Figure 1: Illustrations of approaches for CL of VLMs.
-(a) Continual full fine-tuning is applied toward a universal CL updating.
-Additional reference data is used to alleviate the forgetting of pre-trained knowledge (zero-shot prediction on unseen data), e.g., [62].
+Figure 1: Illustrations of different approaches for CL in VLMs.
+(a) Continual full fine-tuning is applied to achieve universal CL updates, often incorporating additional reference data to mitigate the forgetting of pre-trained knowledge (e.g., zero-shot prediction on unseen data, as in [63]).
 It is computationally expensive and suffers from forgetting.
-(b) New task-specific components isolated to the pre-trained model are added in CL, e.g., [57, 54].
-They require task/domain prediction/gating in inference, increasing inference complexity and restricting the pre-trained model’s improvement potential with limited usage scenarios.
-(c) We aim to continually learn the model by memory-efficient updating of all parameters in a universal manner, similar to (a), but without requiring reference data.
-It allows for enhancing the pre-trained model while learning continuously from the data stream, surpassing the limitations of (b).
-By utilizing dynamic rank-selective LoRA, we reduce forgetting of both the knowledge in the pre-trained model and the knowledge acquired during continual learning.
-Continual Learning (CL) [16, 9, 47, 25, 13] focuses on incrementally learning from new data streams without catastrophic forgetting [36, 32] or the need to retrain from scratch.
-With the advances of large-scale pre-trained models (PTMs), recent CL advances are increasingly leveraging these models.
-While many PTM-based CL approaches address continual adaptation for downstream tasks (e.g., class-incremental learning [51, 46, 20]), with an emphasis on downstream tasks’ performance, there is growing interest in preserving the inherent capabilities of PTMs, such as multi-domain continual learning for Vision-Language Models (VLMs) [62, 57, 54], or in continual pre-training to enhance PTMs over time [41].
-In this paper, we focus on multi-domain continual learning for VLMs like CLIP [39].
-The ZSCL [62] framework aims to continually improve pre-trained model by incrementally feeding it with new knowledge.
-However, this often results in a degradation of both pre-trained and previously acquired knowledge during continual learning steps.
-To address this challenge, prior works have explored strategies like using reference data [62] or memory replay [25, 51, 40] to distill knowledge during full fine-tuning.
-However, these methods incur high training costs due to the need for additional data and extensive retraining (Fig.
-1(a)), and yet they still suffered from the forgetting problem.
-Other approaches [57, 54] introduce task-specific components isolated from the pre-trained model during the CL stream, which require task/domain prediction or gating at inference.
+(b) New task-specific components isolated from the PTM are added for CL, e.g., [58, 55].
+They require task/domain prediction/gating in inference, increasing inference complexity and restricting the PTM’s improvement potential with restricted usage scenarios.
+(c) We aim to continually train the model via parameter-and-data-efficient adaptation in a universal manner, without requiring reference data or full fine-tuning as in (a).
+It enhances the PTM while continually learning from the data stream, surpassing the limitations of (b) on unseen data.
+Figure 2: Overview of our proposed method: we propose the dynamic rank-selection LoRA that allows the each pre-trained weight matrices to adaptively insert the necessary ranks for downstream adaptation while retaining pre-trained capabilities.
+After training on each task, the dynamic ranks of the parameter updates are merged back into the pre-trained weights, without incurring additional inference overhead.
+Continual learning (CL) [17, 10, 48, 27, 14] focuses on incrementally learning from new data streams without catastrophic forgetting [37, 34] or the need to retrain with all seen data.
+With the advances of large-scale pre-trained models (PTMs), recent studies increasingly perform CL with these models [66, 35, 47, 22, 55].
+Many PTM-based CL approaches primarily focus on continual adaptation for downstream tasks (e.g., class-incremental learning, CIL [43, 52, 22, 35]), with an emphasis on downstream task performance.
+While task-specific designs [66, 35, 55] offer promising improvements in adaptation performance, preserving the inherent capabilities of PTMs during continual learning is crucial for maintaining and enhancing generalizable representations.
+This has garnered growing interest in research areas such as multi-domain continual learning for VLMs [63], or in continual pre-training to enhance PTMs over time [42].
+In this paper, we focus on CL of VLMs like CLIP [40] with multi-domain data.
+Some approaches, such as ZSCL [63], aim to continually update PTMs by incrementally incorporating new data with full fine-tuning while mitigating forgetting of pre-trained knowledge through memory replay [27, 52, 41] using additional reference data.
+However, these methods incur high training costs due to the reliance on additional data and extensive retraining (Fig.
+1(a)), yet they still struggle with forgetting.
+To retain PTM capabilities during CL on downstream tasks, some approaches [58, 55] introduce task-specific components that isolated from the PTM during the CL stream, which require task/domain prediction or gating at inference.
 This increases inference complexity and limits the model’s adaptability to unseen data, restricting its generalization potential (Fig.
 1(b)).
-In this work, we explore a more general and practical form of CL for VLMs (similar to [62]), which maintains strong performance on seen data with continually updated pre-trained model, and retains or even boosting the performance on unseen data.
-To ensure memory-efficient learning, we utilize low-rank parameter updates (LoRA) [19].
-Rather than inserting LoRA modules at fixed locations [26, 15], we extend them to all pre-trained weight matrices to enhance adaptation, and crucially, to analyze the impact of each weight matrix on downstream learning and knowledge retention.
-Our analysis (Sec.
-3.2) reveals that not all parameter updates contribute equally to downstream adaptation, with some potentially harming pre-trained capabilities.
-Manually selecting update locations and ranks can significantly improve the balance between adapting to new tasks and retaining pre-trained knowledge.
-This insight motivates our proposed method, Dynamic Rank-Selective LoRA, which adaptively determines both the locations within the pre-trained weight matrices and the ranks of parameter updates.
-We leverage a reformulation of parameter updates using Singular Value Decomposition (SVD) [60, 12, 28, 35, 59].
-Our method introduces learnable importance weights that mimic the singular values in SVD, capturing the significance of each rank in the parameter update.
-These learnable weights enable the model to dynamically adjust the contribution of each rank during training.
-To ensure only the most important updates are integrated into the pre-trained weights, we optimize the importance weights using a sparse prompting approach with a soft-thresholding operation.
+In this work, we propose a practical and efficient CL approach for VLMs that continually trains and updates pre-trained models without relying on downstream task-specific components.
+Our method preserves and even enhances the generalizable representations of PTMs while seamlessly integrating new knowledge from data streams.
+To achieve efficient and effective learning, we conduct comprehensive analyses of low-rank adaptation (LoRA) [20] for CL in PTMs.
+Rather than restricting LoRA to fixed locations as in prior works [28, 16], we extend it to all parameter weights for a thorough investigation, aiming to approximate the capabilities of full fine-tuning.
+We analyze how the rank and placement of LoRA updates impact or regularize the learning dynamics, influencing both downstream adaptation and the retention of existing knowledge (Sec.
+3.2).
+Our analyses reveal that not all LoRA ranks and placements contribute equally to downstream learning and forgetting.
+Generally, for arbitrary weights, relatively higher-rank LoRA facilitates learning new tasks but tends to increase forgetting, while lower-rank LoRA mitigates forgetting but limits adaptation.
+[^1] Similar findings are also observed in the language domain [5].
+However, we identify that there usually exists a balance between plasticity (learning new knowledge) and stability (regularizing forgetting) at a moderately small (but not too small) rank, potentially maximizing CL benefits.
+Moreover, we observe that this balance varies significantly across different parameter placements.
+The analysis results motivate us to propose Continual Dynamic Rank-Selective LoRA (CoDyRA), which performs parameter updates via dynamically optimizing for a minimized LoRA rank suitable for the current data.
+The rank selection for each LoRA is achieved through a Singular Value Decomposition (SVD)-based reformulation [61, 13, 30, 36, 60], with learnable importance weights corresponding to the singular values.
+With sparsity-promoting regularization, these learnable weights allow the model to dynamically adjust the contribution of each rank during training.
+This ensures that only the most essential updates (i.e., LoRA with a minimized rank) are integrated into the pre-trained weights, enabling learning from new data while mitigating forgetting.
 Our contributions are fourfold:
   1. 1.
-    In this paper, we explore the universal CL of VLMs in an memory-efficient manner, maintaining strong performance on both seen and unseen data with reduced forgetting.
-We perform dynamic rank selective parameter updates adaptively to all pre-trained weight matrices.
+    We investigate a universal and practical CL paradigm for efficiently updating pre-trained VLMs on new data streams while mitigating forgetting, without relying on reference data or additional components.
   2. 2.
-    Through comprehensive analyses, we show that the location and rank of parameter updates are critical for achieving an good balance between downstream task adaptation and pre-trained knowledge retention.
+    We conduct in-depth analyses on the impact of LoRA ranks and placements, revealing their crucial role in balancing downstream learning and knowledge retention.
   3. 3.
-    We propose a general low-rank parameter update method that adaptively perform dynamic rank selective updates, effectively integrating new knowledge while maximizing the retention of pre-trained knowledge.
+    We propose CoDyRA, which dynamically optimizes a minimized rank for parameter updates based on current data, with learnable importance weights for each rank.
   4. 4.
-    Extensive experiments and in-depth analyses validate our approach across a wide range of benchmarks for downstream adaptation and zero-shot transfer.
-Additionally, we evaluate on hold-out data to further evaluate knowledge retention.
-2 Related Works
+    Extensive experiments across multiple CL benchmarks validate our method’s superior performance in both downstream learning and mitigating forgetting, and additionally prove superiority on pre-trained knowledge retention using hold-out data.
+2 Related Work
 Continual Learning.
@@
 Continual Learning of VLMs.
-We focus on continual learning for VLMs within two challenging settings: multi-domain task-incremental learning (MTIL) [62, 57] and cross-domain task-agnostic incremental learning (X-TAIL) [54].
+We focus on CL of VLMs with multiple domain data, e.g., multi-domain task-incremental learning (MTIL) [63, 58] and cross-domain task-agnostic incremental learning (X-TAIL) [55].
 In the MTIL setting, the model sequentially learns from a set of $T$ tasks.
 For each task $t$ , the dataset is represented as $\mathcal{D}^{t}=\{({\mathbf{x}}^{t}_{i},y^{t}_{i})\}_{i=1}^{N^{t}}$, where ${\mathbf{x}}^{t}_{i}\in\mathbb{R}^{H\times W\times C}$ is the input image, $y^{t}_{i}\in\mathcal{C}^{t}$ is its corresponding class label, and $N^{t}$ is the number of samples in the task.
-The category set $\mathcal{C}^{t}=\{y^{t}_{j}\}_{j=1}^{M^{t}}$ consists of the class names within the task $t$, containing a total of $M^{t}$ classes.
-In the X-TAIL setting, however, the category set during testing encompasses both previously seen and unseen domains, denoted as $\mathcal{C}=\mathcal{C}_{\text{seen}}\cup\mathcal{C}_{\text{unseen}}$.
+The category set $\mathcal{C}^{t}=\{y^{t}_{j}\}_{j=1}^{M^{t}}$ represents the class labels for task $t$, consisting of $M^{t}$ classes in total.
+In the X-TAIL setting, the category set during testing includes classes from both previously seen and unseen domains, denoted as $\mathcal{C}=\mathcal{C}_{\text{seen}}\cup\mathcal{C}_{\text{unseen}}$.
 Here, $\mathcal{C}_{\text{seen}}=\bigcup_{i=1}^{n}\mathcal{C}_{i}$ includes all classes learned during prior tasks, while $\mathcal{C}_{\text{unseen}}$ represents novel classes that were not encountered during training.
-This introduces a more challenging scenario, as the model must classify test images across a diverse range of domains, making it more complex than the MTIL setting.
+This poses a more challenging scenario, as the model must classify test images across a broader range of domains.
 Transformers.
-A typical transformer model is composed of stacked blocks, where each block consists of two main submodules: an attention (Attn) module utilizing multi-head attention (MHA) and a multi-layer perceptron (MLP).
-In each transformer layer, the attention module at the $l$-th transformer layer can be defined as: $\text{Attn}(x)=\text{MHA}(x\mathbf{W}^{\text{Q}}_{l},x\mathbf{W}^{\text{K}}_{l},x\mathbf{W}^{\text{V}}_{l})\mathbf{W}^{\text{O}}_{l},$ where $\text{MHA}(\cdot)$ denotes the multi-head attention operation.
+A typical transformer model is composed of stacked blocks, where each block consists of two main submodules: an attention (Attn) module utilizing multi-head attention (MHA) and a multilayer perceptron (MLP).
+In each transformer layer, the attention module at the $l$-th layer is defined as: $\text{Attn}(x)=\text{MHA}(x\mathbf{W}^{\text{Q}}_{l},x\mathbf{W}^{\text{K}}_{l},x\mathbf{W}^{\text{V}}_{l})\mathbf{W}^{\text{O}}_{l},$ where $\text{MHA}(\cdot)$ denotes the multi-head attention operation.
 The input sequence is denoted as $x\in\mathbb{R}^{n\times d}$, where $n$ is the sequence length and $d$ is the hidden dimension.
@@
 Vision-Language Models (VLMs).
-In vision-language models like CLIP [39], the image encoder $f_{\theta}$ and text encoder $g_{\psi}$ are implemented as transformers in our case, respectively.
+In vision-language models like CLIP [40], both the image encoder $f_{\theta}$ and text encoder $g_{\psi}$ are implemented as transformers in our case.
 Given an input image ${\mathbf{x}}\in\mathbb{R}^{H\times W\times C}$, it is segmented into patches and processed by the vision encoder to produce latent visual features ${\mathbf{z}}^{\text{V}}=f_{\theta}({\mathbf{x}})\in\mathbb{R}^{D}$.
@@
 This prompt is encoded by the text encoder, yielding ${\mathbf{z}}^{\text{T}}=g_{\psi}({\mathbf{t}})\in\mathbb{R}^{D}$.
-During inference, the probability of classifying image ${\mathbf{x}}$ into class $y_{i}\in\{1,\dots,C\}$ is computed as: $p(y_{i}|{\mathbf{x}})=\frac{\exp(\mathtt{sim}({\mathbf{z}}^{\text{V}},{\mathbf{z}}^{\text{T}}{y_{i}})/\tau)}{\sum_{c=1}^{C}\exp(\mathtt{sim}({\mathbf{z}}^{\text{V}},{\mathbf{z}}^{\text{T}}_{y_{c}})/\tau)},$ where $\mathtt{sim}(\cdot)$ is cosine similarity, and $\tau$ is the temperature.
-Figure 2: Evaluations of adaptation to downstream tasks and retention of zero-shot capabilities after training CLIP with various LoRA insertion locations and ranks.
-Figure 3: Analysis of the effects of removing trained LoRA modules on downstream task adaptation (Aircraft) and retention of pre-trained capabilities (ImageNet-1k).
-Figure 4: Overview of our proposed method: we propose the dynamic rank-selection LoRA that allows the each pre-trained weight matrices to adaptively insert the necessary ranks for downstream adaptation while retaining pre-trained capabilities.
-After training on each task, the dynamic ranks of the parameter updates are merged back into the pre-trained weights, without incurring additional inference overhead.
+During inference, the probability of classifying image ${\mathbf{x}}$ into class $y_{i}\in\{1,\dots,C\}$ is computed using the softmax function as: $p(y_{i}|{\mathbf{x}})=\frac{\exp(\mathtt{sim}({\mathbf{z}}^{\text{V}},{\mathbf{z}}^{\text{T}}{y_{i}})/\tau)}{\sum_{c=1}^{C}\exp(\mathtt{sim}({\mathbf{z}}^{\text{V}},{\mathbf{z}}^{\text{T}}_{y_{c}})/\tau)},$ where $\mathtt{sim}(\cdot)$ is cosine similarity, and $\tau$ is the temperature.
 3.2 Analyses on LoRA Location and Rank
-Analyses of locations and ranks of low-rank updates.
-We begin by investigating the influence of performing parameter updates at different locations within the pre-trained model, varying the update strength to assess its effects on both downstream adaptation and retention of pre-trained capabilities.
-The results of this analysis are presented in Fig.
-3.
-We leverage Low-Rank Adaptation (LoRA) [19], where low-rank parameter updates are directly applied to the pre-trained weight matrices.
-Our findings reveal the following: (1) Updating only the vision encoder leads to catastrophic forgetting of the pre-trained capabilities, significantly diminishing zero-shot performance.
-(2) Updating only the attention modules generally reduces forgetting but restricts the model’s ability to adapt to new tasks.
-(3) Using a higher rank enhances adaptation to downstream tasks without necessarily increasing the forgetting of pre-trained knowledge.
-These observations highlight the need for a strategic selection of insertion locations and ranks to achieve an better balance between adaptation and zero-shot retention.
-Analyses of pruning subsets of low-rank updates.
-To further investigate, we analyzed the effects of selectively perturbing individual modules of the pre-trained weights, as shown in Fig.
-3.
-Initially, we inserted LoRA into all pre-trained matrices with a rank of 16, corresponding to the “Vision + Text, Attn + MLP, rank 16” configuration shown in Fig.
-3.
-We then iteratively remove the LoRA updates learned for each module, examining the impact on both downstream adaptation and the knowledge retention.
-Our findings reveal that not all parameter updates are equally critical for downstream adaptation, and some may even degrade the model’s pre-trained capabilities.
-For example, removing updates in the Value projection of the Attention module within the vision encoder had minimal impact on downstream adaptation but significantly restored the model’s pre-trained zero-shot performance.
-Similarly, removing updates in the MLP layer of the vision encoder slightly reduced downstream task performance but, interestingly, enhanced the model’s pre-trained capabilities, even surpassing its original zero-shot performance.
-3.3 CL VLMs with Dynamic Rank-Selective LoRA
+To achieve efficient learning, we explore LoRA and aim to optimize its usage strategy.
+To analyze the impact of LoRA’s rank and placement on downstream learning and forgetting, we conduct a series of experiments: (1) We apply LoRA components with varying ranks at different locations in the pre-trained CLIP model (Fig.
+4); (2) We apply LoRA across all PTM weights and systematically prune them for evaluation (Fig.
+4).
+We train the model on downstream datasets and assess its performance on both (1) the new task and (2) an additional reference dataset to evaluate the preservation of pre-trained capabilities.
+Applying LoRA on all weights.
+As shown in Fig.
+4 and Fig.
+4, our analysis of LoRA placement shows that applying LoRA to specific modules alone may limit learning capacity in different ways: (1) The different LoRA placement locations significantly affect downstream performance and the forgetting/retention of pre-trained capabilities; (2) Updating only the vision encoder leads to more forgetting of pre-trained capabilities; (3) Updating only the attention modules helps mitigate forgetting but limits the learning ability on new tasks.
+These observations align with the distinct roles of different PTM components.
+Rather than manually selecting LoRA placement, we apply it across all weights and optimize for an adaptive configuration.
+The effects of different LoRA ranks on balancing learning and forgetting.
+From Fig.
+4, we can observe that, despite learning low-rank space, a relatively higher-rank LoRA facilitates learning new tasks but tends to increase forgetting, while lower-rank LoRA mitigates forgetting but limits adaptation.
+This suggests that a lower-rank LoRA produces stronger regularization in maintaining the learned knowledge by sacrificing the learning of new knowledge.
+Importantly, we identify a balance at a moderately small rank, which reduces forgetting while maintaining adaptability.
+This implies that minimizing the rank while fitting on new data may find an optimized trade-off for CL.
+The balance varies by location and task.
+The ranks for a good balance vary by placement (Fig.
+4), and pruning LoRA at different locations has varying effects (Fig.
+4), indicating that not all LoRA components contributed equally for a task.
+For example, removing updates in the Value projection of the Attention module in the vision encoder minimally affected adaptation but significantly restored pre-trained performance.
+Interestingly, pruning the MLP layer of the vision encoder slightly reduced task performance but enhanced pre-trained capabilities.
+These findings motivate our adaptive approach to dynamically determine ranks based on location and task.
+Figure 3: Evaluations of downstream task adaptation and zero-shot capability retention after training CLIP with different LoRA insertion locations and ranks.
+Different shapes represent encoders, colors indicate transformer modules, and sizes denote ranks.
+Figure 4: Analysis of the impact of removing trained LoRA modules on downstream task adaptation (Aircraft) and retention of pre-trained capabilities (ImageNet-1k).
+3.3 Proposed CoDyRA
 Overview.
+Our analyses in Sec.
+3.2 suggest applying LoRA across all CLIP modules while adaptively selecting an appropriate rank for each task and location to balance learning and forgetting.
+While finding an ideally global optimal rank is challenging, we approach it by jointly optimizing downstream task performance and minimizing the number of ranks using sparsity-promoting regularization.
 As shown in Fig.
-4, we introduce dynamic rank-selective parameter updates for all pre-trained weight matrices, adaptively inserting updates based on the importance of each rank for current training data of task $t$.
-Specifically, we integrate dynamic rank-selective LoRA into all pre-trained weights within the Attention and MLP modules of both the vision and text encoders.
-For clarity, for learning on task $t$, we denote the pre-trained weights or those updated from the previous task as $\{\mathbf{W}^{t,m}_{0}\}_{m=1}^{M}$, where $M$ represents the total number of pre-trained weight matrices added with LoRA.
+2, we introduce dynamic rank-selective parameter updates for all pre-trained weight matrices within the Attention and MLP modules of both the vision and text encoders.
+These updates are adaptively inserted based on the importance of each rank for the current training data.
+For clarity, during training on task $t$, we denote the pre-trained weights—or those updated from the previous tasks—as $\{\mathbf{W}^{t,m}_{0}\}_{m=1}^{M}$, where $M$ represents the total number of pre-trained weight matrices augmented with LoRA.
 After training on each task $t$, these dynamic rank updates are merged back into the original pre-trained weights.
+Low-rank updates with rank importance.
+For efficient adaptation, instead of directly updating a pre-trained weight matrix $\mathbf{W}_{0}^{t,m}\in\mathbb{R}^{d\times k}$, LoRA [20] introduces two low-rank matrices, $\mathbf{B}^{t,m}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}^{t,m}\in\mathbb{R}^{r\times k}$, such that $\Delta\mathbf{W}=\mathbf{B}^{t,m}\mathbf{A}^{t,m}$, where $r\ll\min(d,k)$.
+The updated weight matrix is then defined as:
+$\mathbf{W}^{t,m}=\mathbf{W}_{0}^{t,m}+\Delta\mathbf{W}^{t,m}=\mathbf{W}_{0}^{t,m}+\mathbf{B}^{t,m}\mathbf{A}^{t,m}.$ (1)
+In this formulation, the original weights $\mathbf{W}_{0}^{t,m}$ remain fixed, and only $\mathbf{B}^{t,m}$ and $\mathbf{A}^{t,m}$ are trained, reducing the number of trainable parameters to $r(d+k)$.
+Notably, LoRA incurs no additional computational overhead during inference, as the low-rank updates can be efficiently merged into the pre-trained weights, as shown in Eq.
+(1).
+To automatically and adaptively decide the rank in the LoRA updates, we introduce a learnable importance weight vector $\mathbf{w}^{t,m}\in\mathbb{R}^{r}$ and incorporate them into the formulation of LoRA as follows:
+$\Delta\mathbf{W}^{t,m}=\sum\nolimits_{i=1}^{r}\mathbf{w}_{i}^{t,m}\mathbf{B}_{i}^{t,m}\mathbf{A}_{i}^{t,m},$ (2)
+where $\mathbf{w}^{t,m}$ is learnable and represents the importance of each rank.
+The formulation is based on a reformulation of LoRA updates in Eq.
+(1) based on Singular Value Decomposition (SVD) [61, 13, 30, 36, 60].
+The importance weights are corresponding to the diagonal elements in the singular value matrix of SVD.
+Introducing importance weights incurs minimal overhead, adding only $r$ additional parameters per module.
+For initialization, we follow the standard convention: matrix $\mathbf{A}$ is randomly initialized, matrix $\mathbf{B}$ is set to zero, and the importance weights $\mathbf{w}$ are randomly initialized.
     Method Aircraft [33] Caltech101 [15] DTD [9] EuroSAT [18] Flowers [38] Food [6] MNIST [12] OxfordPet [39] Cars [24] SUN397 [54] Average
@@
 […]
-   Ours 74.3 36.8 44.2 69.9 83.5 42.8 88.9 64.6 63.4 63.2
+   CoDyRA 74.3 36.8 44.2 69.9 83.5 42.8 88.9 64.6 63.4 63.2
 […]
@@
 […]
-   Ours 41.4 81.0 58.6 77.8 83.4 84.5 64.5 90.3 67.2 64.4 71.3
+   CoDyRA 41.4 81.0 58.7 77.8 83.4 84.6 64.5 90.4 67.2 64.4 71.3
 […]
@@
 […]
-   Ours 37.7 81.5 65.1 89.9 91.4 85.5 96.8 93.3 77.3 73.5 79.2
+   CoDyRA 37.7 81.5 65.1 89.9 91.4 85.5 96.8 93.3 77.3 73.5 79.2
 […]
@@
 Methods marked with $\dagger$ indicate the use of domain prediction or distribution detection techniques, as in [58, 55].
-Rank-selective LoRA with importance weights.
-LoRA [19] is a parameter-efficient fine-tuning technique designed to adapt pre-trained models using a minimal number of additional parameters.
-Instead of directly updating a pre-trained weight matrix $\mathbf{W}_{0}^{t,m}\in\mathbb{R}^{d\times k}$, LoRA introduces two low-rank matrices, $\mathbf{B}^{t,m}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}^{t,m}\in\mathbb{R}^{r\times k}$, such that $\Delta\mathbf{W}^{t,m}=\mathbf{B}^{t,m}\mathbf{A}^{t,m}$, where $r\ll\min(d,k)$.
-The updated weight matrix is defined as:
-$\mathbf{W}^{t,m}=\mathbf{W}_{0}^{t,m}+\Delta\mathbf{W}^{t,m}=\mathbf{W}_{0}^{t,m}+\mathbf{B}^{t,m}\mathbf{A}^{t,m}.$ (1)
-In this formulation, the original weights $\mathbf{W}_{0}^{t,m}$ remain fixed, and only the matrices $\mathbf{B}^{t,m}$ and $\mathbf{A}^{t,m}$ are trained, significantly reducing the number of trainable parameters to $r(d+k)$.
-Notably, during inference, LoRA incurs no additional computational overhead compared to the original pre-trained model, as the low-rank updates can be merged into the pre-trained weights, as shown in Eq.
-(1).
-Previous works reformulate the low-rank parameter update in the form of Singular Value Decomposition (SVD) [60, 12, 28, 35, 59]. We introduce dynamic importance weights into the formulation:
-$\Delta\mathbf{W}^{t,m}=\sum_{i=1}^{r}\mathbf{w}_{i}^{t,m}\mathbf{B}_{i}^{t,m}\mathbf{A}_{i}^{t,m},$ (2)
-where $\mathbf{w}^{t,m}\in\mathbb{R}^{r}$ is a learnable vector that represents the importance of each rank during training.
-The importance weights imitate the diagonal elements in the singular value matrix of SVD and we allow these weights to be learned directly from the data.
-The introduction of these importance weights incurs only a minimal increase in trainable parameters, adding just $r$ additional parameters per LoRA module.
-For initialization, we adhere to the convention that randomly initialize the matrix $\mathbf{A}$, set the matrix $\mathbf{B}$ to zero, and we initialize the importance weights $\mathbf{w}^{t,m}$ randomly.
-Dynamic rank selection with sparse importance.
+Dynamic rank-selective updates with sparse rank importance.
 The formulation in Eq.
 (2) enables the model to dynamically adjust the importance of each rank through gradient descent.
-[^1] To align the training process with the pre-trained model, we use CLIP’s contrastive loss during pre-training as the supervised training loss.
+[^2] To align the training process with the pre-trained model, we use CLIP’s contrastive loss as the optimization objective during pre-training.
 However, not all ranks contribute equally to downstream task performance (as shown in Fig.
-3).
-To ensure that only the most impactful LoRA ranks are retained in the pre-trained weights, we propose to optimize the learning of importance weights in a sparse-promoting way by minimizing the $\ell_{1}$ norm of the importance weights.
-The optimization objective for paramters $\{\mathbf{w}^{t,m},\mathbf{B}^{t,m},\mathbf{A}^{t,m}\}_{m=1}^{M}$ learned on task $t$ are as follows:
-$\mathcal{L}_{\text{train}}^{t}:=\mathcal{L}_{\text{sup}}^{t}+\lambda\sum_{m=1}^{M}\|\mathbf{w}^{t,m}\|_{1},$ (3)
-where $\mathcal{L}_{\text{sup}}^{t}$ is the supervised training loss, $M$ denotes the total number of pre-trained weight matrices inserted with LoRA of rank $r$, and $\mathbf{w}^{t,m}$ represents the importance weights of the each rank of the LoRA module added to the pre-trained weight matrix $\mathbf{W}^{t,m}_{0}$, and $\lambda$ controls the strength of the $\ell_{1}$ regularization.
-We adopt the proximal gradient [4] for handling the non-differentiable $\ell_{1}$ regularization applied to the importance weights.
-Specifically, the dynamic update of the importance weight for the $i$-th rank of LoRA can be written as a soft-thresholding operation:
+4).
+To ensure that only the most impactful ranks are retained for learning on new task while removing unnecessary ranks to mitigate forgetting of already-learned knowledge, we propose optimizing the importance weights with a sparsity-promoting regularization through minimizing the $\ell_{1}$ norm.
+The optimization objective for parameters $\{\mathbf{w}^{t,m},\mathbf{B}^{t,m},\mathbf{A}^{t,m}\}_{m=1}^{M}$ learned on task $t$ is:
+$\mathcal{L}_{\text{train}}^{t}:=\mathcal{L}_{\text{sup}}^{t}+\lambda\sum\nolimits_{m=1}^{M}\|\mathbf{w}^{t,m}\|_{1},$ (3)
+where $\mathcal{L}_{\text{sup}}^{t}$ is the supervised training loss, $M$ is the total number of pre-trained weight matrices inserted $r$-rank update, and $\mathbf{w}^{t,m}$ represents the importance weights of each rank added to the weight matrix $\mathbf{W}^{t,m}_{0}$.
+The $\ell_{1}$ regularization strength is controlled by $\lambda$.
+To handle the non-differentiable $\ell_{1}$ regularization applied to the importance weights, we adopt the proximal gradient method [4].
+The $i$-th element of $\hat{\mathbf{w}}^{t,m}$ for rank $i$ is updated via soft-thresholding:
 $\mathbf{w}_{i}^{t,m}:=\mathbbm{1}(\lvert\hat{\mathbf{w}}^{t,m}_{i}\rvert>\kappa)\cdot(\hat{\mathbf{w}}^{t,m}_{i}+\mathrm{sign}(\hat{\mathbf{w}}^{t,m}_{i})\cdot\kappa),$ (4)
-where $\hat{\mathbf{w}}^{t,m}_{i}$ denotes the value of ${\mathbf{w}^{t,m}_{i}}$ after gradient update using only the supervised loss $\mathcal{L}_{\text{sup}}^{t}$.
-The threshold parameter $\kappa$ is initially set to zero and gradually increases until it reaches a maximum value $\kappa_{\text{max}}$.
-The indicator function $\mathbbm{1}(\cdot)$ returns 1 if condition is met, 0 otherwise, and $\mathrm{sign}(\cdot)$ represent the +/+/-+ / - sign of input.
-This encourages that only the ranks with significant importance are preserved, while those with low importance are pruned.
-By adjusting the weightings iteratively, the model focuses dynamic parameter updates on the most relevant components.
-In the early stages of training, dense updates are performed without Eq.
-(4) to allow all ranks to capture task-relevant information.
-Continually enhance the pre-trained model.
-After training is complete, ranks with importance weight of zero are removed, while non-zero ranks are merged into the pre-trained weights by substituting Eq.
+where $\hat{\mathbf{w}}^{t,m}_{i}$ denotes the value of ${\mathbf{w}^{t,m}_{i}}$ after applying the gradient update from the supervised loss $\mathcal{L}_{\text{sup}}^{t}$.
+The threshold parameter $\kappa$ is initially set to zero and gradually increases to a maximum value $\kappa_{\text{max}}$, analogous to linearly scheduling the regularization strength $\lambda$.
+The indicator function $\mathbbm{1}(\cdot)$ returns 1 if the condition is met and 0 otherwise, while $\mathrm{sign}(\cdot)$ represents the +/+/-+ / - sign of the input.
+This approach preserves only the ranks with significant importance, pruning those with low relevance.
+By iteratively adjusting weightings, the model focuses dynamic parameter updates on the most relevant components.
+During the early stages of training, dense updates are performed without Eq.
+(4) to allow all ranks to capture task-relevant information before applying sparsity constraints.
+Continually enhancing the pre-trained model.
+After training, ranks with zero importance weights are pruned, while non-zero ranks are merged into the pre-trained weights by substituting Eq.
 (2) into the original LoRA formulation Eq.
 (1).
-When a new task is introduced, a new set of LoRA modules is initialized, and the computations involve only the newly initialized LoRA modules in conjunction with the continually updated pre-trained model.
-This approach allows the model to leverage accumulated knowledge efficiently while keeping the computational footprint minimal without additional overhead.
+When a new task is introduced, a fresh set of CoDyRA modules is initialized, and computations involve only these newly initialized CoDyRA modules alongside the continually updated pre-trained model.
+This approach enables the model to efficiently leverage accumulated knowledge while maintaining a minimal computational footprint, ensuring no additional overhead.
     Method Aircraft [33] Caltech101 [15] CIFAR100 [25] DTD [9] EuroSAT [18] Flowers [38] Food [6] MNIST [12] OxfordPet [39] Cars [24] SUN397 [54] Average
@@
 […]
-   Ours 92.4 68.4 45.8 54.5 69.6 87.4 65.2 88.5 64.2 64.5 69.9
+   CoDyRA 92.4 68.4 45.8 54.5 69.6 87.4 65.2 88.5 64.2 64.5 69.9
 […]
@@
 […]
-   Ours 34.6 95.8 73.9 60.0 77.1 81.3 86.6 75.9 89.9 66.1 65.3 73.3
+   CoDyRA 34.6 95.8 73.9 60.0 77.1 81.3 86.6 75.9 89.9 66.1 65.3 73.3
 […]
@@
 […]
-   Ours 31.6 95.5 72.8 63.5 85.0 89.7 85.0 94.7 93.2 73.6 73.0 78.0
+   CoDyRA 31.6 95.5 72.8 63.5 85.0 89.7 85.0 94.7 93.2 73.6 73.0 78.0
 […]
@@
 Evaluation Metrics.
-We adopt the same evaluation metrics used in prior works, which include “Transfer”, “Average”, and “Last”.
-The “Transfer” metric evaluates the model’s zero-shot transfer capability on previously unseen data.
-The “Last” metric assesses the model’s ability to retain knowledge from earlier tasks.
-The “Average” metric is a composite score that reflects the mean performance across both the “Transfer” and “Last” evaluations.
+We adopt the same evaluation metrics as prior works, including “Transfer”, “Average” , and “Last”.
+“Transfer”evaluates the model’s zero-shot transfer capability on unseen data.
+“Last” assesses the model’s ability to retain knowledge from earlier tasks, and“Average” provides a composite score, representing the mean performance across both the “Transfer” and “Last” evaluations.
 Implementation Details.
-We follow the experimental setups in [62, 57, 54] and utilize the CLIP model with a ViT-B/16 backbone [39] for all experiments.
-By default, our proposed importance-aware sparse LoRA is applied to all pre-trained weight matrices across both the vision encoder and text encoder, with an initial rank set to 16.
+We follow the experimental setups in [63, 58, 55] and use the CLIP model with a ViT-B/16 backbone [40] for all experiments.
+By default, CoDyRA is applied to all pre-trained weight matrices in both the vision and text encoders, with an initial rank of 16.
 Each task in both experimental settings is trained over 500 iterations using the AdamW optimizer [31].
 The dense training iteration ratio is set to 0.5 for X-TAIL and 0.7 for MTIL, relative to the total training iterations.
-The maximum threshold value for adaptive pruning is fixed at 0.005 for both settings.
-Method Average Transfer Accuracy on Unseen Data
-CIFAR100 Places365 ImageNet-1k Average
-Zero-Shot [39] 68.24 33.77 66.72 56.24
-MoE-Adapter [57] 68.24 33.77 66.72 56.24
-RAIL [54] 68.24 33.77 66.72 56.24
-Ours 68.95 36.52 68.52 58.00
-Table 3: Evaluation of average transfer accuracy on datasets that were never encountered during the continual learning stages.
+The maximum threshold adaptive pruning is fixed at 0.005 for both settings.
 Methods
@@
 RAIL [55] N/A 24.18M / 9.01M
-Ours 4.4M None
-Table 4: Comparison on computation cost with previous methods.
+CoDyRA 4.4M None
+Table 3: Comparison of computational cost with prior methods.
 4.2 Experimental Results
 Cross-domain task-agnostic incremental learning.
-We report the evaluation results on the X-TAIL setting in Table 1.
-Unlike previous works that rely on model parameter updates [25, 51, 40, 13] along with additional memory replay or reference data, our approach performs parameter-efficient updates without requiring any memory replay or reference data.
-This results in significant improvements across all evaluated metrics.
-In contrast to prior methods [57, 54] [^2] Dual version of [54] stores all the training samples’ feature embeddings in the memory, for a fair comparison, we excluded the results here.
-We leave more detailed comparisons in Appendix., which use zero-shot predictions to preserve pre-trained capabilities and prevent forgetting, our method overcomes this limitation by continually refining the model.
-Notably, while these approaches are constrained by the upper limit of zero-shot performance, our approach surpasses this limit, achieving a higher average transfer accuracy.
-Although some predictions on unseen data might be slightly impacted due to the nature of continual parameter updates, our method ultimately achieves a higher overall transfer accuracy by effectively adapting the pre-trained model to new tasks while also enhancing its generalization capabilities.
-For both the Average and Last metrics, our method demonstrates superior or comparable performance without relying on domain ID prediction [57], high-dimensional projections, or storing a memory bank of previously seen sample features [54].
-Instead, our method focuses purely on updating the pre-trained model, yet still delivers competitive results.
-Furthermore, since our approach is entirely orthogonal to the additional techniques used in previous methods [57, 54], these techniques can be seamlessly integrated into our framework to further enhance performance, benefited from our ultimate goal of continually enhancing the pre-trained model’s capabilities.
+We report the X-TAIL setting evaluation results in Table 1.
+In contrast to prior methods [58, 55] [^3] The DUAL version of [55] stores all the training samples’ feature embeddings in memory.
+For a fair comparison, we exclude these results here and provide more detailed comparisons in the Appendix., which rely on zero-shot predictions to preserve pre-trained capabilities and mitigate forgetting, our method overcomes this limitation by continually refining the model.
+Notably, while these approaches are bounded by the upper limit of zero-shot performance, our approach surpasses this limit, achieving a higher average transfer accuracy.
+Although some predictions on unseen data might be slightly affected due to the continual parameter updates, our method ultimately improves overall transfer accuracy by adapting the pre-trained model to new tasks while enhancing its generalization capabilities.
+[…]
 Multi-domain task-incremental learning.
 We further evaluate our proposed method under the few-shot MTIL setting, as shown in Table 2, following [58, 55].
-Consistent with the results observed in the X-TAIL setting, our method demonstrates clear superiority in this setting.
-Without employing any additional techniques, our approach outperforms previous methods across all evaluation metrics, achieving an average Transfer accuracy that surpasses the upper bound set by zero-shot predictions in prior works.
-Moreover, by supplementing our method with techniques used in prior approaches [57, 54], we can further improve the performance.
+Consistent with the results observed in the X-TAIL setting, our method demonstrates clear superiority in this scenario.
+Without incorporating any additional techniques, our approach outperforms previous methods across all evaluation metrics, achieving an average transfer accuracy that surpasses the upper bound set by zero-shot predictions in prior works.
+Furthermore, by integrating techniques from previous approaches [58, 55], our method can be further enhanced, leading to even greater performance gains.
 4.3 Discussions
+Method Average Transfer Accuracy on Unseen Data
+CIFAR100 Places365 ImageNet-1k Average
+Zero-Shot [40] 68.24 33.77 66.72 56.24
+MoE-Adapter [58] 68.24 33.77 66.72 56.24
+RAIL [55] 68.24 33.77 66.72 56.24
+CoDyRA 68.95 36.52 68.52 58.00
+Table 4: Evaluation of average transfer accuracy on unseen datasets not encountered during continual learning.
 Figure 5: Visualization and statistical analysis of rank activation on the Aircraft dataset using our proposed method.
 Figure 6: Visualization and statistical analysis of rank activation on the Oxford Pets dataset using our proposed method.
-Pre-trained knowledge retention and potentially improved pre-trained model.
-To understand how our method improves transfer performance beyond the upperbound set by zero-shot predictions in previous works, we evaluate on datasets that were never encountered during the continual learning steps, specifically CIFAR100 [24], Places365 [63], and ImageNet-1k [10], as shown in Table 3.
-Unlike prior approaches [57, 54] that rely exclusively on zero-shot predictions for unseen data, thereby limiting their performance to that of the original CLIP model, our method goes beyond this limitation.
-By effectively integrating new knowledge into the pre-trained parameter space without sacrificing existing capabilities, our approach enhances the model’s overall performance.
-Consequently, our method achieves superior transfer accuracy, surpassing the theoretical upper bound defined by zero-shot performance in previous methods.
-This not only demonstrates improved adaptation but also enhances the model’s generalization capabilities on completely unseen datasets during CL.
-Rank allocation results on different datasets.
+Pre-trained knowledge retention and potential pre-trained model improvement.
+To understand how our method improves transfer performance beyond the upper bound set by zero-shot predictions in prior works, we evaluated on hold-out datasets never encountered during continual learning, specifically CIFAR100 [25], Places365 [64], and ImageNet-1k [11], as shown in Table 4.
+Unlike prior approaches [58, 55], which rely exclusively on zero-shot predictions for unseen data—thereby constraining their performance to that of the original CLIP model—our method overcomes this limitation.
+By effectively integrating new knowledge into the pre-trained parameter space while preserving existing capabilities, our approach enhances overall model performance.
+As a result, it achieves superior zero-shot transfer accuracy, surpassing the zero-shot performance limits of prior works.
+This demonstrates that CoDyRA effectively incorporates general knowledge into its knowledge base, further improving its overall representation capability.
+Rank allocation across different datasets.
 We visualize and analyze the distribution of activated ranks merged into the pre-trained weights across different datasets in Fig.
@@
 6.
-Overall, it is evident that the Aircraft dataset generally requires more ranks to be updated in the pre-trained model compared to the Pets dataset.
-We observe that for the Aircraft dataset, a greater number of ranks are allocated to the Output projection of the Attention module within the text encoder (Fig.
+Overall, the Aircraft dataset generally requires more rank updates in the PTM than the Pets dataset.
+For the Aircraft dataset, we observe that a greater number of ranks are allocated to the Output projection of the Attention module within the text encoder (Fig.
 6(b)).
-Additionally, a higher concentration of ranks is assigned to the deeper layers, particularly layers 6 to 10 (Fig.
+Additionally, a higher concentration of ranks is assigned to deeper layers, particularly layers 6 to 10 (Fig.
 6(c)).
-In contrast, the analysis of the Pets dataset (Fig.
-6(b) and Fig.
-6(c)) reveals that more ranks are allocated to the Attentions’ Key and Value projection, and MLP’s FC layer of vision encoder.
-Furthermore, the majority of rank allocations are in the first and last layers, suggesting a different pattern for promoting knowledge retention and adaptation across datasets.
-These findings highlight the adaptive nature of our approach, where rank allocation varies based on the characteristics of the dataset, optimizing the model’s ability to capture domain-specific information effectively.
+In contrast, for the Pets dataset, more ranks are allocated to the Attention module’s Value and Output projections and the MLP’s Projection layer within the vision encoder (Fig.
+6(b)).
+Additionally, rank allocation is primarily concentrated in the first and last layers (Fig.
+6(c)), indicating a distinct adaptation pattern that balances knowledge retention and domain adaptation across datasets.
+These observations highlight the adaptive nature of our approach, where rank allocation dynamically adjusts based on dataset characteristics, ensuring effective domain adaptation while preserving pre-trained knowledge.
+We provide more visualizations and analyses in Appendix B.3.
 Computation cost.
 We compare the computational costs in terms of trainable parameters and additional inference requirements, as summarized in Fig.
-4.
-For MoE-Adapter, the difference between training and testing parameters stems from its mixture-of-experts design: during training, all experts are active, while at inference, only the top-2 experts are selected, reducing the active parameter count.
-RAIL-Primal incurs additional computational overhead during inference due to its use of high-dimensional feature expansion through projection and regression layers.
-RAIL-Dual maintains a memory bank of image features for all training samples across 1,100 classes, which significantly increases memory usage.
+3.
+For MoE-Adapter, the difference between training and testing parameters arises from its mixture-of-experts design: during training, all experts are active, whereas during inference, only the top-2 experts are selected, reducing the number of active parameters.
+RAIL-Primal incurs additional inference overhead due to its use of high-dimensional feature expansion through projection and regression layers.
+Meanwhile, RAIL-DUAL significantly increases memory usage by maintaining a memory bank of image features for all training samples across 1,100 classes.
 Notably, these methods experience linearly growing computational costs as the number of tasks increases.
 In contrast, our method minimizes training overhead by selectively retaining only ranks with high importance.
-Once training is complete, these updates are merged into the pre-trained weights, resulting in zero additional memory overhead during inference.
+Once training is complete, these updates are merged into the pre-trained weights, ensuring zero additional memory overhead during inference.
 Figure 7: Ablation studies on different insertion locations and the impact of varying the initial rank.
@@
 Effects of insertion locations and number of initial ranks.
-We examine the impact of different insertion locations and initial ranks for our method, with results presented in Fig.
+We examine the impact of different insertion locations and initial ranks in our method, with results presented in Fig.
 7.
-These findings are consistent with our analytical experiments shown in Fig.
-3.
-Our results indicate that when training is restricted to only the vision encoder, transfer accuracy is significantly reduced, suggesting that the pre-trained capabilities are negatively impacted.
-Conversely, increasing the maximum rank improves adaptation ability without necessarily reducing transfer accuracy.
-Furthermore, when limiting insertion to only the Attention modules, we observe that transfer accuracy can surpass our previously reported state-of-the-art results in Table 1 and Table 2.
+These findings are consistent with our analyses shown in Fig.
+4.
+Our results indicate that when training is restricted to the vision encoder, transfer accuracy drops significantly, suggesting that the pre-trained capabilities are negatively impacted.
+Moreover, at specific locations, increasing the maximum rank improves adaptation ability while not always reducing transfer accuracy.
+Interestingly, restricting updates to only the Attention modules can surpass our reported state-of-the-art Transfer metric in Table 1 and Table 2.
 However, this comes at the expense of reduced adaptation capabilities.
-To achieve an better balance between adaptation and retention, we enable insertion into all modules by default.
-This configuration ensures a better trade-off between adapting to new tasks while preserving the model’s zero-shot transfer capabilities.
-Effect of maximum threshold value and ratio for dense iterations.
-We analyze the effects of varying the maximum threshold value $\kappa_{\text{max}}$ and the ratio for dense training iterations in Fig.
+To strike a better balance between adaptation and retention, we enable updates across all modules by default, ensuring an optimal trade-off between learning new tasks and retaining the model’s pre-trained capabilities.
+Effect of maximum threshold value and dense iteration ratio.
+We analyze the effects of varying the maximum threshold value $\kappa_{\text{max}}$ and the ratio of dense training iterations in Fig.
 8.
-The results indicate that as the threshold increases, the Transfer accuracy shows an increasing trend.
-This suggests that pruning more ranks can enhance the model’s performance on unseen data (i.e., improve transfer capability), but it also limits the model’s ability to adapt to downstream tasks, as evidenced by changes in the Last accuracy metric.
-Our experiments show that a threshold value around 0.005 strikes a good balance between retaining pre-trained knowledge and adapting to new tasks.
-The ratio of dense training iterations determines when the model begins the adaptive pruning process.
-Setting this ratio to 0 (i.e., starting pruning from the beginning of training) severely hinders adaptation, as LoRA lacks sufficient time to learn essential task-related knowledge before pruning begins.
-On the other hand, delaying pruning for too long can result in the model forgetting pre-trained capabilities, thereby reducing transfer accuracy.
-Our findings indicate that setting the ratio around 0.5 to 0.7 works best across all experimental settings, providing a good balance between adaptation and knowledge retention.
+Our results indicate that as the threshold increases, Last accuracy improves, suggesting that pruning more ranks can enhance transfer performance on unseen data.
+However, aggressive pruning limits adaption ability to to downstream tasks, as reflected in changes to the Last accuracy metric.
+We find that setting the threshold value around 0.005 strikes a good balance between retaining pre-trained knowledge and adapting to new tasks.
+The dense training iteration ratio determines when adaptive pruning begins.
+Setting this ratio to 0 (i.e., pruning from the start of training) hinders adaptation, as the model lacks sufficient time to learn task-relevant knowledge before pruning begins.
+Conversely, delaying pruning too long increases the risk of forgetting pre-trained capabilities, ultimately reducing transfer accuracy.
+We find that setting this ratio between 0.5 and 0.7 achieves the best trade-off across all experimental settings, effectively balancing both adaptation and knowledge retention.
 5 Conclusion
-In this paper, we introduced a novel Dynamic Rank-Selective LoRA approach to address the challenges of continual learning for vision-language models.
-By dynamically adjusting the ranks based on each ranks’ importance, our method optimizes knowledge retention while enabling efficient adaptation to new tasks without adding inference overhead.
-Extensive evaluations on multi-domain and cross-domain settings demonstrated that our method achieves state-of-the-art performance, effectively balancing downstream adaptation and zero-shot generalization.
+In this paper, we introduced CoDyRA, a novel approach to addressing continual learning challenges in vision-language models.
+By dynamically adjusting ranks based on their importance, our method optimizes knowledge retention while enabling efficient adaptation to new tasks without adding inference overhead.
+Extensive evaluations across various continual learning settings demonstrate that our method achieves state-of-the-art performance, effectively balancing downstream adaptation and pre-trained retention.
 Limitations and future works.
-Our current approach directly merges parameter updates into the pre-trained weights without explicitly considering correlations between tasks.
-Exploring methods to leverage inter-task relationships and shared information build a promising direction for future research.
+Our approach merges updates into pre-trained weights without modeling inter-task correlations.
+Future works could explore leveraging inter-task relationships and extending our method to other VLMs.
 Acknowledgment
@@
 Lu is affiliated with CSIRO Data61 through a PhD scholarship and acknowledges the support of the Google PhD Fellowship.
-Supplementary Material
 Appendix A Additional Details of the Proposed Method
@@
 B.1 Effects of Additional Adaptive Prediction Head on Visual Representations
-Related continual learning methods focus on updating representations through the backbone model (e.g., [50, 62, 57], and ours) or by adding a representation alignment or adaptive prediction head [33, 54].
-These two approaches can be considered orthogonal.
-We investigate how the representation updates introduced by our proposed method can complement the additional adaptive prediction head, utilizing the aggression-based adapters described in [54].
+The related continual learning methods focus on updating the representation with the backbone model (e.g., [51, 63, 58], and CoDyRA) or newly added representation alignment or prediction head [35, 55].
+These two types of methods can be seen as orthogonal.
+We study how the representation updated by the proposed can work with the additional adaptive prediction head, relying on the aggression-based adapters in [55].
 We provide a more detailed analysis of the effects of the additional prediction head on the visual representations in X-TAIL setting in Table 5.
 Note that the Dual version of RAIL [55] stores all visual representations of all samples for all seen classes and domains in memory, it is not included in Tables 1 and 2 for a fair comparison.
-Since RAIL heavily relies on the frozen pre-trained model, the Transfer accuracy is limited by the zero-shot performance of the pre-trained model.
+RAIL is a regression-based method that applies an additional prediction head on top of the visual representations extracted by the frozen pre-trained vision encoder of the CLIP model.
+This approach is fully orthogonal to our method, as our focus is on continually incorporating new knowledge into the pre-trained model through dynamic low-rank parameter updates, and we can seamlessly integrate their techniques into our method.
+Because RAIL heavily relies on the frozen pre-trained model, the Transfer accuracy is limited by the zero-shot performance of the pre-trained model.
 In contrast, our approach gradually accumulates knowledge into the pre-trained model, enhancing its capability on unseen data and achieving improved Transfer accuracy.
@@
 RAIL-Primal 62.4 70.7 79.1
-Ours + RAIL-Primal 63.1 74.1 83.4
+CoDyRA + RAIL-Primal 63.1 74.1 83.4
 RAIL-Dual 62.4 71.9 82.4
-Ours + RAIL-Dual 63.1 74.2 83.6
+CoDyRA + RAIL-Dual 63.1 74.2 83.6
 Table 5: Experimental results for integrating aggression-based adapters in the X-TAIL setting. Metrics for “Transfer”, “Average”, and “Last” are reported as averages across all datasets.
@@
 The amplification factor is given by:
-$\text{Amp}=\frac{\|\Delta\mathbf{W}\|_{F}}{\|\mathbf{U}^{T}\mathbf{W}\mathbf{V}^{T}\|_{F}},$ (8)
+$\text{Amp}=\frac{\|\Delta\mathbf{W}\|_{F}}{\|\mathbf{U}^{\text{T}}\mathbf{W}\mathbf{V}^{\text{T}}\|_{F}},$ (8)
 where $\mathbf{W}$ is the pre-trained weight matrix, and $\Delta\mathbf{W}$ is the parameter change introduced by LoRA.
-For vanilla LoRA and our method, $\Delta\mathbf{W}$ is computed using Eq.
+For vanilla LoRA and CoDyRA, $\Delta\mathbf{W}$ is computed using Eq.
 (1) and Eq.
@@
 (c) Pets - LoRA
-(d) Aircraft - Ours
-(e) EuroSAT - Ours
-(f) Pets - Ours
+(d) Aircraft - CoDyRA
+(e) EuroSAT - CoDyRA
+(f) Pets - CoDyRA
 Figure 9: Visualization of amplification factors when trained on (a, d) Aircraft, (b, e) EuroSAT, and (c, f) Pets, comparing (d, e, f) our method of dynamic sparse rank selective LoRA and (a, b, c) vanilla LoRA.
 It shows that the proposed method with adaptive rank selection can achieve more focused/concentrated updating and less interference.
-B.3 More Visualizations and Analyses of Rank Allocation Across Datasets
+B.3 More Visualizations and Analyses on Rank allocation results of different datasets
 In Sec.
@@
 Figure 17: Visualization and statistical analysis of rank activation on the SUN397 dataset using our proposed method.
+B.4 Evaluation with Other VLMs
+To evaluate the broader applicability of our method, we conducted preliminary experiments with BLIP [26] on X-TAIL (Table 6).
+Our method consistently outperformed prior works and exceeded the pre-trained model’s upper-bound performance, demonstrating its general effectiveness across different VLMs.
+Transfer Average Last
+Zero-shot 51.03 46.45
+[…]
+[…]
+Table 6: Averaged performance using BLIP.
+B.5 Evaluation on Standard Parameter-Efficient Fine-Tuning (PEFT) Task
+Leveraging our dynamic rank-adaptive parameter updates, our method naturally extends to Parameter-Efficient Fine-Tuning (PEFT) tasks.
+Table 7 presents 16-shot PEFT results, using the same hyperparameters as the main paper, showing that CoDyRA outperforms vanilla LoRA while reducing final trainable parameters by pruning ranks with zero importance.
+Notably, across multiple downstream tasks (Caltech101, Cars, and OxfordPet), CoDyRA achieves higher accuracy while using fewer parameters compared to LoRA with a fixed rank of 16.
+These results highlight a more adaptive and efficient alternative to conventional low-rank tuning by automatically selecting the most relevant ranks based on training data.
+Crucially, by amplifying important ranks through learned higher importance weights, our approach significantly enhances downstream adaptation.
+Acc. $\uparrow$ / Final Used Params. $\downarrow$
+Caltech101 Cars OxfordPet
+Zero-shot 88.40 64.70 89.00
+LoRA ($r=16$) 96.43 / 4.4M 80.66 / 4.4M 92.94 / 4.4M
+CoDyRA  ($r_{init}=16$) 97.20 / 3.95M 81.61 / 3.75M 94.44 / 3.85M
+Table 7: Performance of 16-shot PEFT on downstream tasks.

### v5

@@
 Abstract
-We investigate whether the pre-trained knowledge in vision-language models (VLMs), such as CLIP, can be retained—or even enhanced—in continual learning (CL) while incorporating new knowledge from the data stream.
-Existing CL methods primarily focus on continual downstream adaptation using components isolated from pre-trained model (PTM), increasing inference complexity and limiting improvements to the PTM itself; some also retain knowledge relying on additional reference data, leading to high training costs.
-To address these limitations, we propose a universal and efficient Continual Learning approach for VLM based on Dynamic Rank-Selective LoRA (CoDyRA), which directly improves the PTMs while preserving the existing knowledge from both pre-training and CL.
-Through analyses on how LoRA rank and placement impact/regularize the learning and forgetting in CL, we propose CoDyRA that adaptively performs rank-minimized parameter updates in different modules, based on their importance to the current data.
+Continual learning (CL) aims to accumulate knowledge from sequential data and task streams.
+Leveraging their strong generalization and flexibility, pre-trained vision-language embedding models such as CLIP (Contrastive Language-Image Pre-training) have been widely adopted and validated in CL.
+In addition to learning new knowledge, we investigate whether the pre-trained knowledge in CLIP, can be retained or even enhanced, in CL, while incorporating new knowledge from a data stream.
+Existing CL methods primarily focus on continual downstream adaptation using components isolated from the pre-trained model (PTM), increasing inference complexity and limiting improvements to the PTM itself; some also retain knowledge by relying on additional reference data, resulting in high training costs.
+To address these limitations, we propose a universal and efficient CL approach for CLIP based on Dynamic Rank-Selective LoRA (CoDyRA), which directly improves the PTMs while preserving the existing knowledge from both pre-training and CL.
+By analyzing how LoRA rank and placement affect learning and forgetting in CL, we design CoDyRA that adaptively performs rank-minimized parameter updates in different modules, based on their importance to the current data.
 This ensures a balance between knowledge acquisition (plasticity) and forgetting mitigation (stability).
@@
 Moreover, CoDyRA preserves the original model architecture and deployment pipeline, introducing no additional inference overhead.
-Extensive experiments demonstrate that our approach enhances representations based on new downstream data while retaining pre-trained knowledge, achieving state-of-the-art results.
+Extensive experiments show that our approach enhances representations for new downstream data while retaining pre-trained knowledge, achieving state-of-the-art results.
 1 Introduction
-Figure 1: Illustrations of different approaches for CL in VLMs.
-(a) Continual full fine-tuning is applied to achieve universal CL updates, often incorporating additional reference data to mitigate the forgetting of pre-trained knowledge (e.g., zero-shot prediction on unseen data, as in [63]).
-It is computationally expensive and suffers from forgetting.
-(b) New task-specific components isolated from the PTM are added for CL, e.g., [58, 55].
-They require task/domain prediction/gating in inference, increasing inference complexity and restricting the PTM’s improvement potential with restricted usage scenarios.
-(c) We aim to continually train the model via parameter-and-data-efficient adaptation in a universal manner, without requiring reference data or full fine-tuning as in (a).
-It enhances the PTM while continually learning from the data stream, surpassing the limitations of (b) on unseen data.
-Figure 2: Overview of our proposed method: we propose the dynamic rank-selection LoRA that allows the each pre-trained weight matrices to adaptively insert the necessary ranks for downstream adaptation while retaining pre-trained capabilities.
-After training on each task, the dynamic ranks of the parameter updates are merged back into the pre-trained weights, without incurring additional inference overhead.
-Continual learning (CL) [17, 10, 48, 27, 14] focuses on incrementally learning from new data streams without catastrophic forgetting [37, 34] or the need to retrain with all seen data.
-With the advances of large-scale pre-trained models (PTMs), recent studies increasingly perform CL with these models [66, 35, 47, 22, 55].
+Continual learning (CL) [18, 9, 49, 29, 13] focuses on incrementally learning from new data streams without catastrophic forgetting [39, 36, 17] or the need to retrain on all previously seen data.
+With the rise of large-scale pre-trained models (PTMs), recent studies increasingly perform CL in this context [70, 37, 48, 24, 58].
 Many PTM-based CL approaches primarily focus on continual adaptation for downstream tasks (e.g., class-incremental learning, CIL [44, 54, 24, 37]), with an emphasis on downstream task performance.
-While task-specific designs [66, 35, 55] offer promising improvements in adaptation performance, preserving the inherent capabilities of PTMs during continual learning is crucial for maintaining and enhancing generalizable representations.
-This has garnered growing interest in research areas such as multi-domain continual learning for VLMs [63], or in continual pre-training to enhance PTMs over time [42].
-In this paper, we focus on CL of VLMs like CLIP [40] with multi-domain data.
-Some approaches, such as ZSCL [63], aim to continually update PTMs by incrementally incorporating new data with full fine-tuning while mitigating forgetting of pre-trained knowledge through memory replay [27, 52, 41] using additional reference data.
-However, these methods incur high training costs due to the reliance on additional data and extensive retraining (Fig.
-1(a)), yet they still struggle with forgetting.
-To retain PTM capabilities during CL on downstream tasks, some approaches [58, 55] introduce task-specific components that isolated from the PTM during the CL stream, which require task/domain prediction or gating at inference.
-This increases inference complexity and limits the model’s adaptability to unseen data, restricting its generalization potential (Fig.
+While these task-specific designs [70, 37, 58] yield promising results, preserving the inherent capabilities of PTMs during CL is critical to maintaining and improving generalizable representations.
+Benefiting from the strong generalization and flexibility of jointly learned vision-text embeddings, vision-language models such as CLIP have been widely adopted and studied in CL [24, 66, 50, 69], including continual updating [17] or multi-domain CL [66] of CLIP.
+Some approaches, such as ZSCL [66], aim to update PTM, i.e., CLIP, by incrementally incorporating new data through full fine-tuning while mitigating forgetting via memory replay [29, 54, 43] using additional reference data.
+However, this leads to high training costs and still struggles with forgetting (Fig.
+1(a)).
+Other methods [61, 58] retain PTM capabilities during downstream CL by introducing task-specific components that are isolated from the PTM, requiring task/domain prediction or gating at inference.
+This increases inference complexity and limits adaptability to unseen data, restricting the model’s generalization potential (Fig.
 1(b)).
-In this work, we propose a practical and efficient CL approach for VLMs that continually trains and updates pre-trained models without relying on downstream task-specific components.
+Figure 1: CL methods for CLIP: (a) Continual full fine-tuning uses reference data to apply universal updates and mitigate forgetting of pre-trained knowledge (e.g., zero-shot performance on unseen data [66]), but is computationally expensive and still susceptible to forgetting; (b) Task-specific modular methods (e.g., [61, 58]) introduce isolated components that require domain prediction or gating at inference, increasing complexity and limiting adaptability and generalization; (c) Our approach enables universal, parameter- and data-efficient adaptation without reference data or task-specific modules.
+It continually improves the pre-trained model and achieves superior performance on unseen data.
+In this work, we propose a practical and efficient CL approach for CLIP that continually trains and updates pre-trained models without relying on downstream task-specific components.
 Our method preserves and even enhances the generalizable representations of PTMs while seamlessly integrating new knowledge from data streams.
@@
 Generally, for arbitrary weights, relatively higher-rank LoRA facilitates learning new tasks but tends to increase forgetting, while lower-rank LoRA mitigates forgetting but limits adaptation.
-[^1] Similar findings are also observed in the language domain [5].
 However, we identify that there usually exists a balance between plasticity (learning new knowledge) and stability (regularizing forgetting) at a moderately small (but not too small) rank, potentially maximizing CL benefits.
 Moreover, we observe that this balance varies significantly across different parameter placements.
-The analysis results motivate us to propose Continual Dynamic Rank-Selective LoRA (CoDyRA), which performs parameter updates via dynamically optimizing for a minimized LoRA rank suitable for the current data.
+The analysis results motivate us to propose Continual Dynamic Rank-Selective LoRA (CoDyRA), which performs parameter updates by dynamically optimizing for a minimized LoRA rank suitable for the current data.
 The rank selection for each LoRA is achieved through a Singular Value Decomposition (SVD)-based reformulation [64, 12, 32, 38, 63], with learnable importance weights corresponding to the singular values.
@@
   1. 1.
-    We investigate a universal and practical CL paradigm for efficiently updating pre-trained VLMs on new data streams while mitigating forgetting, without relying on reference data or additional components.
+    We investigate a universal and practical CL paradigm for efficiently updating pre-trained CLIP on new data streams while mitigating forgetting, without reference data or additional components.
   2. 2.
@@
 Parameter regularization approaches [25, 1, 62, 2, 23] constrain updates to important weights to preserve past knowledge.
-Meanwhile, dynamic networks [45, 46, 65, 47, 35, 28, 51, 50, 43, 49] adjust their architecture on the fly and balance learning new information while retaining old knowledge.
-Multi-Domain Continual Learning of VLMs.
-Continual learning of VLMs [22, 62] aims to enable VLMs to sequentially learn across diverse domains while retaining their pre-trained generalization capabilities for previously seen tasks.
+Meanwhile, dynamic networks [46, 47, 68, 48, 37, 30, 53, 52, 44, 51] adjust their architecture on the fly, balancing the acquisition of new information with the retention of old knowledge.
+Continual Learning of CLIP.
+Continual learning of CLIP [24, 65] aims to enable PTMs to sequentially learn across diverse domains while retaining their pre-trained generalization capabilities for previously seen tasks.
+[17] continually adapt CLIP to temporal distribution shifts via timestamped data updates.
 The approach introduced in ZSCL [66] addresses the challenge of preserving zero-shot capabilities while adapting to new tasks, mitigating the risk of catastrophic forgetting.
@@
 Low-Rank Adaptation (LoRA) [21] has gained popularity as a parameter-efficient method for fine-tuning large pre-trained models.
-Recent works enhance LoRA by reformulating parameter updates using SVD, focusing on initialization from the SVD of pre-trained weights [60, 36], treating each rank independently [13, 61, 30], or employing a mixture of ranks [53].
+Similar to LoRA, [56] learns a small set of task-specific representation vectors at each layer.
+Building on LoRA’s framework, recent works enhance LoRA by reformulating parameter updates using SVD, focusing on initialization from the SVD of pre-trained weights [63, 38], treating each rank independently [12, 64, 32], or employing a mixture of ranks [55].
 These advancements highlight LoRA’s effectiveness in efficiently fine-tuning large models while preserving generalization.
+Figure 2: Overview of CoDyRA: we propose the dynamic rank-selection LoRA that allows the each pre-trained weight matrices to adaptively insert the necessary ranks for downstream adaptation while retaining pre-trained capabilities.
+After training on each task, the dynamic ranks of the parameter updates are merged back into the pre-trained weights, without incurring additional inference overhead.
 3 Methodology
 3.1 Preliminaries
-Continual Learning of VLMs.
-We focus on CL of VLMs with multiple domain data, e.g., multi-domain task-incremental learning (MTIL) [63, 58] and cross-domain task-agnostic incremental learning (X-TAIL) [55].
-In the MTIL setting, the model sequentially learns from a set of $T$ tasks.
-For each task $t$ , the dataset is represented as $\mathcal{D}^{t}=\{({\mathbf{x}}^{t}_{i},y^{t}_{i})\}_{i=1}^{N^{t}}$, where ${\mathbf{x}}^{t}_{i}\in\mathbb{R}^{H\times W\times C}$ is the input image, $y^{t}_{i}\in\mathcal{C}^{t}$ is its corresponding class label, and $N^{t}$ is the number of samples in the task.
-The category set $\mathcal{C}^{t}=\{y^{t}_{j}\}_{j=1}^{M^{t}}$ represents the class labels for task $t$, consisting of $M^{t}$ classes in total.
-In the X-TAIL setting, the category set during testing includes classes from both previously seen and unseen domains, denoted as $\mathcal{C}=\mathcal{C}_{\text{seen}}\cup\mathcal{C}_{\text{unseen}}$.
-Here, $\mathcal{C}_{\text{seen}}=\bigcup_{i=1}^{n}\mathcal{C}_{i}$ includes all classes learned during prior tasks, while $\mathcal{C}_{\text{unseen}}$ represents novel classes that were not encountered during training.
-This poses a more challenging scenario, as the model must classify test images across a broader range of domains.
 Transformers.
@@
 The MLP module at the $l$-th transformer layer is defined as: $\text{MLP}(x)=\text{Act}(x\mathbf{W}^{\text{FC}}_{l}+b^{\text{FC}}_{l})\mathbf{W}^{\text{Proj}}_{l}+b^{\text{Proj}}_{l},$ where $\mathbf{W}^{\text{FC}}_{l}\in\mathbb{R}^{d\times d_{m}}$ and $\mathbf{W}^{\text{Proj}}_{l}\in\mathbb{R}^{d_{m}\times d}$ are the weight matrices, with $d_{m}$ being the hidden dimension of the MLP.
-Vision-Language Models (VLMs).
-In vision-language models like CLIP [40], both the image encoder $f_{\theta}$ and text encoder $g_{\psi}$ are implemented as transformers in our case.
+Vision-language CLIP model.
+In CLIP model with joint vision-language embedding [42, 22, 14, 28], both the image encoder $f_{\theta}$ and text encoder $g_{\psi}$ are implemented as transformers in our case.
 Given an input image ${\mathbf{x}}\in\mathbb{R}^{H\times W\times C}$, it is segmented into patches and processed by the vision encoder to produce latent visual features ${\mathbf{z}}^{\text{V}}=f_{\theta}({\mathbf{x}})\in\mathbb{R}^{D}$.
@@
 During inference, the probability of classifying image ${\mathbf{x}}$ into class $y_{i}\in\{1,\dots,C\}$ is computed using the softmax function as: $p(y_{i}|{\mathbf{x}})=\frac{\exp(\mathtt{sim}({\mathbf{z}}^{\text{V}},{\mathbf{z}}^{\text{T}}{y_{i}})/\tau)}{\sum_{c=1}^{C}\exp(\mathtt{sim}({\mathbf{z}}^{\text{V}},{\mathbf{z}}^{\text{T}}_{y_{c}})/\tau)},$ where $\mathtt{sim}(\cdot)$ is cosine similarity, and $\tau$ is the temperature.
+CL of CLIP.
+We focus on CL of CLIP with multiple domain data, e.g., multi-domain task-incremental learning (MTIL) [66, 61] and cross-domain task-agnostic incremental learning (X-TAIL) [58].
+In the MTIL setting, the model sequentially learns from a set of $T$ tasks.
+For each task $t$ , the dataset is represented as $\mathcal{D}^{t}=\{({\mathbf{x}}^{t}_{i},y^{t}_{i})\}_{i=1}^{N^{t}}$, where ${\mathbf{x}}^{t}_{i}\in\mathbb{R}^{H\times W\times C}$ is the input image, $y^{t}_{i}\in\mathcal{C}^{t}$ is its corresponding class label, and $N^{t}$ is the number of samples in the task.
+The category set $\mathcal{C}^{t}=\{y^{t}_{j}\}_{j=1}^{M^{t}}$ represents the class labels for task $t$, consisting of $M^{t}$ classes in total.
+In the X-TAIL setting, the category set during testing includes classes from both previously seen and unseen domains, denoted as $\mathcal{C}=\mathcal{C}_{\text{seen}}\cup\mathcal{C}_{\text{unseen}}$.
+Here, $\mathcal{C}_{\text{seen}}=\bigcup_{i=1}^{n}\mathcal{C}_{i}$ includes all classes learned during prior tasks, while $\mathcal{C}_{\text{unseen}}$ represents novel classes that were not encountered during training.
+This poses a more challenging scenario, as the model must classify test images across a broader range of domains.
 3.2 Analyses on LoRA Location and Rank
+Figure 3: Evaluations of downstream task adaptation and zero-shot capability retention when training CLIP with varying LoRA insertion points and ranks.
+Shapes indicate encoders, colors denote transformer modules, and sizes reflect rank values.
+Figure 4: Analysis of the impact of removing trained LoRA modules on downstream task adaptation (Aircraft) and retention of pre-trained capabilities (ImageNet-1k).
 To achieve efficient learning, we explore LoRA and aim to optimize its usage strategy.
@@
 4 and Fig.
-4, our analysis of LoRA placement shows that applying LoRA to specific modules alone may limit learning capacity in different ways: (1) The different LoRA placement locations significantly affect downstream performance and the forgetting/retention of pre-trained capabilities; (2) Updating only the vision encoder leads to more forgetting of pre-trained capabilities; (3) Updating only the attention modules helps mitigate forgetting but limits the learning ability on new tasks.
+4, applying LoRA to specific modules alone may limit learning capacity in different ways: (1) The different LoRA placement locations significantly affect downstream performance and the forgetting/retention of pre-trained capabilities; (2) Updating only the vision encoder leads to more forgetting of pre-trained capabilities; (3) Updating only the attention modules helps mitigate forgetting but limits the learning ability on new tasks.
 These observations align with the distinct roles of different PTM components.
@@
 These findings motivate our adaptive approach to dynamically determine ranks based on location and task.
-Figure 3: Evaluations of downstream task adaptation and zero-shot capability retention after training CLIP with different LoRA insertion locations and ranks.
-Different shapes represent encoders, colors indicate transformer modules, and sizes denote ranks.
-Figure 4: Analysis of the impact of removing trained LoRA modules on downstream task adaptation (Aircraft) and retention of pre-trained capabilities (ImageNet-1k).
 3.3 Proposed CoDyRA
@@
 While finding an ideally global optimal rank is challenging, we approach it by jointly optimizing downstream task performance and minimizing the number of ranks using sparsity-promoting regularization.
-As shown in Fig.
-2, we introduce dynamic rank-selective parameter updates for all pre-trained weight matrices within the Attention and MLP modules of both the vision and text encoders.
+We introduce dynamic rank-selective parameter updates for all pre-trained weight matrices within the Attention and MLP modules of both the vision and text encoders, as shown in Fig.
+2.
 These updates are adaptively inserted based on the importance of each rank for the current training data.
@@
 (1).
-To automatically and adaptively decide the rank in the LoRA updates, we introduce a learnable importance weight vector $\mathbf{w}^{t,m}\in\mathbb{R}^{r}$ and incorporate them into the formulation of LoRA as follows:
+To automatically and adaptively decide the rank in the LoRA updates, we introduce a learnable importance weight vector $\mathbf{w}^{t,m}\in\mathbb{R}^{r}$ and incorporate it into the formulation of LoRA as follows:
 $\Delta\mathbf{W}^{t,m}=\sum\nolimits_{i=1}^{r}\mathbf{w}_{i}^{t,m}\mathbf{B}_{i}^{t,m}\mathbf{A}_{i}^{t,m},$ (2)
@@
 For initialization, we follow the standard convention: matrix $\mathbf{A}$ is randomly initialized, matrix $\mathbf{B}$ is set to zero, and the importance weights $\mathbf{w}$ are randomly initialized.
+Table 1: Comparisons on X-TAIL for each domain in terms of “Transfer”, “Average”, and “Last” scores (%).
+The best and the second best results are highlighted in red and blue, respectively.
+Methods marked with $\dagger$ indicate the use of domain prediction or distribution detection [61, 58].
     Method Aircraft [35] Caltech101 [15] DTD [8] EuroSAT [19] Flowers [40] Food [5] MNIST [11] OxfordPet [41] Cars [26] SUN397 [57] Average
@@
 […]
-Table 1: Comparison of different CL methods on X-TAIL for each domain in terms of “Transfer”, “Average”, and “Last” scores (%).
-The best and the second best results are highlighted in red and blue, respectively.
-Methods marked with $\dagger$ indicate the use of domain prediction or distribution detection techniques, as in [58, 55].
 Dynamic rank-selective updates with sparse rank importance.
@@
 After training, ranks with zero importance weights are pruned, while non-zero ranks are merged into the pre-trained weights by substituting Eq.
-(2) into the original LoRA formulation Eq.
+(2) into the formulation Eq.
 (1).
@@
 This approach enables the model to efficiently leverage accumulated knowledge while maintaining a minimal computational footprint, ensuring no additional overhead.
+Table 2: Comparisons on 5-shot MTIL setting, evaluated in a similar manner as in Table 1.
     Method Aircraft [35] Caltech101 [15] CIFAR100 [27] DTD [8] EuroSAT [19] Flowers [40] Food [5] MNIST [11] OxfordPet [41] Cars [26] SUN397 [57] Average
@@
 […]
-Table 2: Comparison with state-of-the-art methods on 5-shot MTIL setting in terms of “Transfer”, “Average”, and “Last” scores (%).
-The best and the second best results are highlighted in red and blue, respectively.
-Methods marked with $\dagger$ indicate the use of domain prediction or distribution detection techniques, as in [58, 55].
 4 Experiments
@@
 “Transfer”evaluates the model’s zero-shot transfer capability on unseen data.
-“Last” assesses the model’s ability to retain knowledge from earlier tasks, and“Average” provides a composite score, representing the mean performance across both the “Transfer” and “Last” evaluations.
+“Last” assesses the model’s ability to retain knowledge from earlier tasks, and“Average” provides a composite score, representing the mean performance across both the “Transfer” and “Last”.
 Implementation Details.
@@
 The dense training iteration ratio is set to 0.5 for X-TAIL and 0.7 for MTIL, relative to the total training iterations.
-The maximum threshold adaptive pruning is fixed at 0.005 for both settings.
+The maximum threshold for adaptive pruning $\kappa$ is fixed at 0.005 for both settings.
+4.2 Experimental Results
+Cross-domain task-agnostic incremental learning.
+We report the X-TAIL setting evaluation results in Table 1.
+In contrast to prior methods [61, 58] , which rely on zero-shot predictions to preserve pre-trained capabilities and mitigate forgetting, our method overcomes this limitation by continually refining the model.
+Notably, while these approaches are bounded by the upper limit of zero-shot performance, our approach surpasses this limit, achieving a higher average transfer accuracy.
+Although some predictions on unseen data might be slightly affected due to the continual parameter updates, our method ultimately improves overall transfer accuracy by adapting the pre-trained model to new tasks while enhancing its generalization capabilities.
+[…]
+Multi-domain task-incremental learning.
+We further evaluate our proposed method under the few-shot MTIL setting, as shown in Table 2, following [61, 58].
+Consistent with the results observed in the X-TAIL setting, our method demonstrates clear superiority in this scenario.
+Without incorporating any additional techniques, our approach outperforms previous methods across all evaluation metrics, achieving an average transfer accuracy that surpasses the upper bound set by zero-shot predictions in prior works.
+Furthermore, by integrating techniques from previous approaches [61, 58], our method can be further enhanced, leading to even greater performance gains.
+4.3 Discussions
+Pre-trained knowledge retention and potential pre-trained model improvement.
+To understand how our method improves transfer performance beyond the upper bound set by zero-shot predictions in prior works, we evaluated on hold-out datasets never encountered during continual learning, specifically CIFAR100 [27], Places365 [67], and ImageNet-1k [10], as shown in Table 3.
+Table 3: Average transfer accuracy on unseen datasets not encountered during CL.
+Method Average Transfer Accuracy on Unseen Data
+CIFAR100 Places365 ImageNet-1k Average
+Zero-Shot [42] 68.24 33.77 66.72 56.24
+MoE-Adapter [61] 68.24 33.77 66.72 56.24
+RAIL [58] 68.24 33.77 66.72 56.24
+CoDyRA 68.95 36.52 68.52 58.00
+Unlike prior approaches [61, 58], which rely exclusively on zero-shot predictions for unseen data—thereby constraining their performance to that of the original CLIP model—our method overcomes this limitation.
+By effectively integrating new knowledge into the pre-trained parameter space while preserving existing capabilities, our approach enhances overall model performance.
+As a result, it achieves superior zero-shot transfer accuracy, surpassing the zero-shot performance limits of prior works.
+This demonstrates that CoDyRA effectively incorporates general knowledge into its knowledge base, further improving its overall representation capability.
+Figure 5: Visualization and statistical analysis of rank activation on the Aircraft dataset.
+Figure 6: Visualization and statistical analysis of rank activation on the Oxford Pets dataset.
+Rank allocation across different datasets.
+We visualize and analyze the distribution of activated ranks merged into the pre-trained weights across different datasets in Fig.
+6 and Fig.
+6.
+Overall, the Aircraft dataset generally requires more rank updates in the PTM than the Pets dataset.
+Table 4: Computation cost.
 Methods
@@
 CoDyRA 4.4M None
-Table 3: Comparison of computational cost with prior methods.
-4.2 Experimental Results
-Cross-domain task-agnostic incremental learning.
-We report the X-TAIL setting evaluation results in Table 1.
-In contrast to prior methods [58, 55] [^3] The DUAL version of [55] stores all the training samples’ feature embeddings in memory.
-For a fair comparison, we exclude these results here and provide more detailed comparisons in the Appendix., which rely on zero-shot predictions to preserve pre-trained capabilities and mitigate forgetting, our method overcomes this limitation by continually refining the model.
-Notably, while these approaches are bounded by the upper limit of zero-shot performance, our approach surpasses this limit, achieving a higher average transfer accuracy.
-Although some predictions on unseen data might be slightly affected due to the continual parameter updates, our method ultimately improves overall transfer accuracy by adapting the pre-trained model to new tasks while enhancing its generalization capabilities.
-[…]
-Multi-domain task-incremental learning.
-We further evaluate our proposed method under the few-shot MTIL setting, as shown in Table 2, following [58, 55].
-Consistent with the results observed in the X-TAIL setting, our method demonstrates clear superiority in this scenario.
-Without incorporating any additional techniques, our approach outperforms previous methods across all evaluation metrics, achieving an average transfer accuracy that surpasses the upper bound set by zero-shot predictions in prior works.
-Furthermore, by integrating techniques from previous approaches [58, 55], our method can be further enhanced, leading to even greater performance gains.
-4.3 Discussions
-Method Average Transfer Accuracy on Unseen Data
-CIFAR100 Places365 ImageNet-1k Average
-Zero-Shot [40] 68.24 33.77 66.72 56.24
-MoE-Adapter [58] 68.24 33.77 66.72 56.24
-RAIL [55] 68.24 33.77 66.72 56.24
-CoDyRA 68.95 36.52 68.52 58.00
-Table 4: Evaluation of average transfer accuracy on unseen datasets not encountered during continual learning.
-Figure 5: Visualization and statistical analysis of rank activation on the Aircraft dataset using our proposed method.
-Figure 6: Visualization and statistical analysis of rank activation on the Oxford Pets dataset using our proposed method.
-Pre-trained knowledge retention and potential pre-trained model improvement.
-To understand how our method improves transfer performance beyond the upper bound set by zero-shot predictions in prior works, we evaluated on hold-out datasets never encountered during continual learning, specifically CIFAR100 [25], Places365 [64], and ImageNet-1k [11], as shown in Table 4.
-Unlike prior approaches [58, 55], which rely exclusively on zero-shot predictions for unseen data—thereby constraining their performance to that of the original CLIP model—our method overcomes this limitation.
-By effectively integrating new knowledge into the pre-trained parameter space while preserving existing capabilities, our approach enhances overall model performance.
-As a result, it achieves superior zero-shot transfer accuracy, surpassing the zero-shot performance limits of prior works.
-This demonstrates that CoDyRA effectively incorporates general knowledge into its knowledge base, further improving its overall representation capability.
-Rank allocation across different datasets.
-We visualize and analyze the distribution of activated ranks merged into the pre-trained weights across different datasets in Fig.
-6 and Fig.
-6.
-Overall, the Aircraft dataset generally requires more rank updates in the PTM than the Pets dataset.
 For the Aircraft dataset, we observe that a greater number of ranks are allocated to the Output projection of the Attention module within the text encoder (Fig.
@@
 Additionally, rank allocation is primarily concentrated in the first and last layers (Fig.
-6(c)), indicating a distinct adaptation pattern that balances knowledge retention and domain adaptation across datasets.
+6(c)), indicating a distinct adaptation pattern that balances knowledge retention and adaptation across datasets.
 These observations highlight the adaptive nature of our approach, where rank allocation dynamically adjusts based on dataset characteristics, ensuring effective domain adaptation while preserving pre-trained knowledge.
-We provide more visualizations and analyses in Appendix B.3.
+We provide more visualizations and analyses in Appendix.
 Computation cost.
 We compare the computational costs in terms of trainable parameters and additional inference requirements, as summarized in Fig.
-3.
+4.
 For MoE-Adapter, the difference between training and testing parameters arises from its mixture-of-experts design: during training, all experts are active, whereas during inference, only the top-2 experts are selected, reducing the number of active parameters.
@@
 Once training is complete, these updates are merged into the pre-trained weights, ensuring zero additional memory overhead during inference.
-Figure 7: Ablation studies on different insertion locations and the impact of varying the initial rank.
-Figure 8: Analyses of the effects of (a) varying the maximum pruning threshold and (b) different ratios of dense training iterations.
+Figure 7: Ablations on different insertion locations and the impact of varying the initial rank.
+Figure 8: Analyses (a) maximum pruning threshold and (b) ratios of dense training iterations.
 4.4 Ablation Studies
@@
 To strike a better balance between adaptation and retention, we enable updates across all modules by default, ensuring an optimal trade-off between learning new tasks and retaining the model’s pre-trained capabilities.
-Effect of maximum threshold value and dense iteration ratio.
-We analyze the effects of varying the maximum threshold value $\kappa_{\text{max}}$ and the ratio of dense training iterations in Fig.
-8.
-Our results indicate that as the threshold increases, Last accuracy improves, suggesting that pruning more ranks can enhance transfer performance on unseen data.
-However, aggressive pruning limits adaption ability to to downstream tasks, as reflected in changes to the Last accuracy metric.
-We find that setting the threshold value around 0.005 strikes a good balance between retaining pre-trained knowledge and adapting to new tasks.
-The dense training iteration ratio determines when adaptive pruning begins.
-Setting this ratio to 0 (i.e., pruning from the start of training) hinders adaptation, as the model lacks sufficient time to learn task-relevant knowledge before pruning begins.
-Conversely, delaying pruning too long increases the risk of forgetting pre-trained capabilities, ultimately reducing transfer accuracy.
-We find that setting this ratio between 0.5 and 0.7 achieves the best trade-off across all experimental settings, effectively balancing both adaptation and knowledge retention.
+Analyses of pruning threshold and dense-iteration ratio.
+To better understand the trade-off between downstream learning and forgetting, we control the number of adaptive ranks by varying the maximum threshold $\kappa_{\max}$ and the fraction of dense training iterations (Fig.
+8).
+Gradually raising $\kappa_{\max}$ improves Last accuracy by pruning more useless or noisy ranks, whereas overly aggressive pruning limits task adaptation; a modest setting of 0.005 strikes a good balance.
+The dense-iteration ratio governs pruning onset: pruning from the start (ratio = 0) removes ranks before it can learn from the training data, while delaying pruning too long increases the risk of retaining useless or noisy ranks that do not get pruned eventually.
 5 Conclusion
@@
 Limitations and future works.
-Our approach merges updates into pre-trained weights without modeling inter-task correlations.
-Future works could explore leveraging inter-task relationships and extending our method to other VLMs.
-Acknowledgment
+Our approach merges updates into pre-trained weights without modeling inter-task correlations, mainly evaluated using pre-trained CLIP model.
+Future works could explore leveraging inter-task relationships and extending our method to other PTMs and settings.
+Acknowledgments and Disclosure of Funding
 This work was partially supported by the ARC DECRA Fellowship (DE230101591) awarded to D.
@@
 In this section, we provide more details to enhance clarity and understanding.
-Reproducibility Statement. The source code will be made publicly available upon acceptance of the paper.
 A.1 Full Derivation of Eq. (4)
@@
 Appendix B More Experimental Results and Ablations
+Reproducibility Statement.
+The source code will be made publicly available upon acceptance of the paper.
+We report experiment results averaged over 3 independent runs for reproducibility.
+Experiments were conducted on a single NVIDIA GeForce RTX 3090 GPU.
 B.1 Effects of Additional Adaptive Prediction Head on Visual Representations
@@
 Moreover, by leveraging our continually enhanced pre-trained model, coupling our method with RAIL significantly outperforms the original RAIL, which uses a frozen pre-trained model, on the Average and Last metrics.
+Table 5: Experimental results for integrating aggression-based adapters in the X-TAIL setting. Metrics for “Transfer”, “Average”, and “Last” are reported as averages across all datasets.
 Transfer Average Last
@@
 CoDyRA + RAIL-Dual 63.1 74.2 83.6
-Table 5: Experimental results for integrating aggression-based adapters in the X-TAIL setting. Metrics for “Transfer”, “Average”, and “Last” are reported as averages across all datasets.
 B.2 Analysis of Changes/Interference w.r.t. Pre-Trained Weights
@@
 Our method consistently outperformed prior works and exceeded the pre-trained model’s upper-bound performance, demonstrating its general effectiveness across different VLMs.
+Table 6: Averaged performance using BLIP.
 Transfer Average Last
@@
 […]
-Table 6: Averaged performance using BLIP.
+Table 7: Performance of 16-shot PEFT on downstream tasks.
+Acc. $\uparrow$ / Final Used Params. $\downarrow$
+Caltech101 Cars OxfordPet
+Zero-shot 88.40 64.70 89.00
+LoRA ($r=16$) 96.43 / 4.4M 80.66 / 4.4M 92.94 / 4.4M
+CoDyRA  ($r_{init}=16$) 97.20 / 3.95M 81.61 / 3.75M 94.44 / 3.85M
 B.5 Evaluation on Standard Parameter-Efficient Fine-Tuning (PEFT) Task
@@
 Crucially, by amplifying important ranks through learned higher importance weights, our approach significantly enhances downstream adaptation.
-Acc. $\uparrow$ / Final Used Params. $\downarrow$
-Caltech101 Cars OxfordPet
-Zero-shot 88.40 64.70 89.00
-LoRA ($r=16$) 96.43 / 4.4M 80.66 / 4.4M 92.94 / 4.4M
-CoDyRA  ($r_{init}=16$) 97.20 / 3.95M 81.61 / 3.75M 94.44 / 3.85M
-Table 7: Performance of 16-shot PEFT on downstream tasks.
+Broader Impacts
+CoDyRA’s sparse, adaptive, importance-aware low-rank subspace learning significantly reduces compute and memory demands for continual learning, enabling on-device personalization in domains such as mobile health and education and empowering resource-constrained users and organizations.
+As a fundamental research effort, CodyRA does not address potential biases or hazards in training data, and its incremental updates could introduce privacy risks or stealthy backdoors.
+We therefore recommend transparent update logging, regular audits, and differential-privacy safeguards.
+Although more efficient than full fine-tuning, large-scale deployment still incurs energy costs; developers should monitor resource use, leverage low-power hardware, and explore carbon offsets.

### v6

@@
 Abstract
-Continual learning (CL) aims to accumulate knowledge from sequential data and task streams.
-Leveraging their strong generalization and flexibility, pre-trained vision-language embedding models such as CLIP (Contrastive Language-Image Pre-training) have been widely adopted and validated in CL.
-In addition to learning new knowledge, we investigate whether the pre-trained knowledge in CLIP, can be retained or even enhanced, in CL, while incorporating new knowledge from a data stream.
-Existing CL methods primarily focus on continual downstream adaptation using components isolated from the pre-trained model (PTM), increasing inference complexity and limiting improvements to the PTM itself; some also retain knowledge by relying on additional reference data, resulting in high training costs.
-To address these limitations, we propose a universal and efficient CL approach for CLIP based on Dynamic Rank-Selective LoRA (CoDyRA), which directly improves the PTMs while preserving the existing knowledge from both pre-training and CL.
-By analyzing how LoRA rank and placement affect learning and forgetting in CL, we design CoDyRA that adaptively performs rank-minimized parameter updates in different modules, based on their importance to the current data.
-This ensures a balance between knowledge acquisition (plasticity) and forgetting mitigation (stability).
-Our method operates without explicit domain or distribution prediction and does not rely on reference data, enabling seamless task integration while maintaining pre-trained capabilities.
-Moreover, CoDyRA preserves the original model architecture and deployment pipeline, introducing no additional inference overhead.
-Extensive experiments show that our approach enhances representations for new downstream data while retaining pre-trained knowledge, achieving state-of-the-art results.
+Continual learning (CL) aims to accumulate knowledge from sequential tasks without catastrophic forgetting.
+Vision–language models like CLIP, with strong generalization, are widely used for CL.
+Existing methods often adapt isolated PTM components, adding inference complexity and limiting PTM improvement, or rely on replay, stored information, or assumptions, incurring high costs and limited applicability.
+To advance models as continual learners, we explore CL via natural, efficient PTM updates instead of complex task-specific additions.
+We thus study continual low-rank learning and systematically analyze how LoRA ranks and placements affect learning and forgetting.
+We find that a relatively higher-rank LoRA improves task learning (i.e., plasticity) but increases forgetting, while a relatively lower-rank LoRA reduces forgetting (i.e., stability) but limits adaptation.
+Crucially, we find a plasticity–stability balance tied to rank across parameters and tasks, with moderately small ranks maximizing CL benefits.
+Motivated by this, we propose Continual Dynamic Rank-Selective LoRA (CoDyRA), which continually updates PTMs with LoRA adapters for all parameters of adaptively optimized rank number.
+While the new-task objective drives learning, CoDyRA adaptively minimizes ranks with sparsity-promoting regularization to reduce interference and forgetting, achieving a plasticity–stability balance tailored to different parameters and tasks.
+Despite allowing updates to all parameters, the adaptively selected and minimized LoRA ranks help keep the updated model closer to its previous state, while still enabling comprehensive and effective learning of new tasks.
+CoDyRA enables efficient CL as a sequence of LoRA-based tasks without storing past data, task information, or relying on assumptions.
+It preserves the original model architecture and deployment pipeline, adding no inference overhead.
+Extensive experiments show CoDyRA improves new representations while retaining old knowledge, achieving state-of-the-art results.
+Code is available.
 1 Introduction
-Continual learning (CL) [18, 9, 49, 29, 13] focuses on incrementally learning from new data streams without catastrophic forgetting [39, 36, 17] or the need to retrain on all previously seen data.
-With the rise of large-scale pre-trained models (PTMs), recent studies increasingly perform CL in this context [70, 37, 48, 24, 58].
-Many PTM-based CL approaches primarily focus on continual adaptation for downstream tasks (e.g., class-incremental learning, CIL [44, 54, 24, 37]), with an emphasis on downstream task performance.
-While these task-specific designs [70, 37, 58] yield promising results, preserving the inherent capabilities of PTMs during CL is critical to maintaining and improving generalizable representations.
-Benefiting from the strong generalization and flexibility of jointly learned vision-text embeddings, vision-language models such as CLIP have been widely adopted and studied in CL [24, 66, 50, 69], including continual updating [17] or multi-domain CL [66] of CLIP.
-Some approaches, such as ZSCL [66], aim to update PTM, i.e., CLIP, by incrementally incorporating new data through full fine-tuning while mitigating forgetting via memory replay [29, 54, 43] using additional reference data.
-However, this leads to high training costs and still struggles with forgetting (Fig.
+Continual learning (CL) (Hadsell et al., 2020; De Lange et al., 2021; Wang et al., 2024b) focuses on incrementally learning from new data streams without catastrophic forgetting (Nguyen et al., 2019; McCloskey & Cohen, 1989) or the need to retrain on all previously seen data.
+With the rise of large-scale pre-trained models (PTMs), recent studies increasingly perform CL in this context (Zhou et al., 2024; McDonnell et al., 2024; Wang et al., 2024a; Jha et al., 2024).
+Many PTM-based CL approaches primarily focus on continual adaptation for downstream tasks (e.g., class-incremental learning, CIL (Smith et al., 2023; Wortsman et al., 2022)), with an emphasis on downstream tasks.
+While task-specific and downstream-oriented designs (Zhou et al., 2024; McDonnell et al., 2024; Xu et al., 2024; Yu et al., 2024) show promising results on particular CL tasks, PTMs remain largely isolated from the CL process, mitigating forgetting but undermining interaction with new tasks and thus the potential for continual improvement.
+To advance toward models that function as continual learners, we consider CL with PTMs that enable flexible and efficient updates, preserving or even enhancing generalizable representations while acquiring new knowledge.
+Leveraging the strong generalization and flexibility of jointly learned vision–text embeddings, CLIP as a vision–language model has been widely adopted in CL (Jha et al., 2024; Zheng et al., 2023; Wang et al., 2023; Zhou et al., 2023), including continual updating (Garg et al., 2023) and multi-domain CL (Zheng et al., 2023).
+Beyond downstream performance, the retention of knowledge in PTMs is also evaluated.
+Some approaches, such as ZSCL (Zheng et al., 2023), aim to continually update PTM, i.e., CLIP, through full fine-tuning while mitigating forgetting via memory replay (Li & Hoiem, 2017; Wortsman et al., 2022; Rebuffi et al., 2017b) using additional reference data.
+However, this leads to high training costs and still struggles with catastrophic forgetting (Fig.
 1(a)).
-Other methods [61, 58] retain PTM capabilities during downstream CL by introducing task-specific components that are isolated from the PTM, requiring task/domain prediction or gating at inference.
-This increases inference complexity and limits adaptability to unseen data, restricting the model’s generalization potential (Fig.
+To directly control the forgetting, some methods (Yu et al., 2024; Xu et al., 2024) introduce task-specific components isolated from the frozen PTMs, requiring task/domain prediction or gating at inference.
+These methods can deliver strong results with task-specific gating and branching; however, by isolating tasks and relying heavily on per-task modules and accurate task prediction, they become restricted to certain domains and limited in both general CL applicability and inference efficiency.
+Moreover, with restricted interaction between PTMs and CL process (e.g., isolation) limits PTMs’ generalization improvement potential (Fig.
 1(b)).
-Figure 1: CL methods for CLIP: (a) Continual full fine-tuning uses reference data to apply universal updates and mitigate forgetting of pre-trained knowledge (e.g., zero-shot performance on unseen data [66]), but is computationally expensive and still susceptible to forgetting; (b) Task-specific modular methods (e.g., [61, 58]) introduce isolated components that require domain prediction or gating at inference, increasing complexity and limiting adaptability and generalization; (c) Our approach enables universal, parameter- and data-efficient adaptation without reference data or task-specific modules.
-It continually improves the pre-trained model and achieves superior performance on unseen data.
-In this work, we propose a practical and efficient CL approach for CLIP that continually trains and updates pre-trained models without relying on downstream task-specific components.
-Our method preserves and even enhances the generalizable representations of PTMs while seamlessly integrating new knowledge from data streams.
-To achieve efficient and effective learning, we conduct comprehensive analyses of low-rank adaptation (LoRA) [21] for CL in PTMs.
-Rather than restricting LoRA to fixed locations as in prior works [30, 16], we extend it to all parameter weights for a thorough investigation, aiming to approximate the capabilities of full fine-tuning.
-We analyze how the rank and placement of LoRA updates impact or regularize the learning dynamics, influencing both downstream adaptation and the retention of existing knowledge (Sec.
+Figure 1: (a) Continual full fine-tuning leverages reference data to reduce forgetting of pre-trained knowledge (e.g., zero-shot performance (Zheng et al., 2023)) but remains costly and prone to forgetting.
+(b) Task-specific modular methods (e.g., (Yu et al., 2024; Xu et al., 2024)) add isolated components requiring domain prediction or gating, increasing complexity and limiting adaptability and generalization.
+(c) Our approach enables efficient, universal continual PTM updating without reference data or task-specific designs, retaining PTM knowledge and improving performance on unseen data.
+Straightforward adaptation of PTMs can be achieved through direct model updating through full fine-tuning (FT) or parameter-efficient fine-tuning (PEFT) methods (Zhou et al., 2022c; b), such as Low-Rank Adaptation (LoRA) (Hu et al., 2021).
+This raises a fundamental question: instead of relying on complex additions or task-specific components, can CL be performed by leveraging these natural PTM update mechanisms?
+Our work investigates how to mitigate catastrophic forgetting within these standard frameworks without imposing extensive assumptions.
+We propose an efficient CL approach for CLIP that enables the model to be updated naturally across a stream of tasks, while retaining foundational pre-trained knowledge.
+We build upon LoRA for efficiency.
+While there have been some LoRA-based CL methods (Liang & Li, 2024; Wu et al., 2025; Zhu et al., 2025), they rely on specific regularization (e.g., orthogonal property) for specific settings.
+We first systematically analyze and ablate how LoRA adapters at different locations in the network with different ranks influence learning and forgetting (Sec.
 3.2).
 Our analyses reveal that not all LoRA ranks and placements contribute equally to downstream learning and forgetting.
-Generally, for arbitrary weights, relatively higher-rank LoRA facilitates learning new tasks but tends to increase forgetting, while lower-rank LoRA mitigates forgetting but limits adaptation.
-However, we identify that there usually exists a balance between plasticity (learning new knowledge) and stability (regularizing forgetting) at a moderately small (but not too small) rank, potentially maximizing CL benefits.
+Our analyses empirically demonstrate that, for arbitrary weights, relatively higher-rank LoRA facilitates learning new tasks (i.e., plasticity) but tends to increase forgetting, while lower-rank LoRA mitigates forgetting (i.e., stability) but limits adaptation.
+Importantly, we identify that there exists a balance between plasticity (learning new knowledge) and stability (regularizing forgetting) at a moderately small (but not too small) rank, potentially maximizing CL benefits.
 Moreover, we observe that this balance varies significantly across different parameter placements.
-The analysis results motivate us to propose Continual Dynamic Rank-Selective LoRA (CoDyRA), which performs parameter updates by dynamically optimizing for a minimized LoRA rank suitable for the current data.
-The rank selection for each LoRA is achieved through a Singular Value Decomposition (SVD)-based reformulation [64, 12, 32, 38, 63], with learnable importance weights corresponding to the singular values.
-With sparsity-promoting regularization, these learnable weights allow the model to dynamically adjust the contribution of each rank during training.
-This ensures that only the most essential updates (i.e., LoRA with a minimized rank) are integrated into the pre-trained weights, enabling learning from new data while mitigating forgetting.
-Our contributions are fourfold:
-  1. 1.
-    We investigate a universal and practical CL paradigm for efficiently updating pre-trained CLIP on new data streams while mitigating forgetting, without reference data or additional components.
-  2. 2.
-    We conduct in-depth analyses on the impact of LoRA ranks and placements, revealing their crucial role in balancing downstream learning and knowledge retention.
-  3. 3.
-    We propose CoDyRA, which dynamically optimizes a minimized rank for parameter updates based on current data, with learnable importance weights for each rank.
-  4. 4.
-    Extensive experiments across multiple CL benchmarks validate our method’s superior performance in both downstream learning and mitigating forgetting, and additionally prove superiority on pre-trained knowledge retention using hold-out data.
+The analysis results motivate us to propose Continual Dynamic Rank-Selective LoRA (CoDyRA), which continually updates the PTM using LoRA adapters with dynamically and adaptively optimized ranks.
+CoDyRA performs new task learning while jointly minimizing the rank of each LoRA adaptively.
+Rank selection is guided by sparsity-promoting regularization and learnable importance weights, which dynamically adjust the contribution of each rank during training.
+By updating with LoRA adapters of minimized ranks, CoDyRA inherently biases models toward remaining closer to their previous state, thereby alleviating forgetting in principle without task-specific assumptions.
+Our contributions are fourfold: (1) We study PTM-based CL with natural and efficient continual update of the model, not relying on complex task-specific additions, without storing past data, task information, or relying on assumptions.
+(2) We comprehensively analyze learning–forgetting trade-offs across LoRA ranks and placements, identifying a plasticity–stability balance at moderately small ranks and deriving design strategies.
+This study guides our CL approach with regularized forgetting and provides evidence for future research.
+(3) We propose CoDyRA, which adaptively minimizes LoRA ranks via sparsity-promoting regularization.
+It handles different parameters and tasks, yielding regularized updates that reduce interference and keep the model closer to its previous state with less forgetting.
+(4) Extensive experiments across multiple CL benchmarks validate our method’s superior performance in both downstream learning and mitigating forgetting, achieving a balanced trade-off between plasticity and stability.
 2 Related Work
@@
 Continual learning enables models to sequentially learn new tasks while retaining previously acquired knowledge.
-Experience replay (ER) methods [34, 3, 6, 31, 7, 59, 60] store subsets of past data to refresh the model on prior tasks during new training.
-Parameter regularization approaches [25, 1, 62, 2, 23] constrain updates to important weights to preserve past knowledge.
-Meanwhile, dynamic networks [46, 47, 68, 48, 37, 30, 53, 52, 44, 51] adjust their architecture on the fly, balancing the acquisition of new information with the retention of old knowledge.
+Experience replay (ER) methods (Luo et al., 2023; Aljundi et al., 2019b; Chaudhry et al., 2018a; Liu et al., 2020; Chaudhry et al., 2018b; Yan et al., 2022; 2021) store subsets of past data to refresh the model on prior tasks during new training.
+Parameter regularization approaches (Kirkpatrick et al., 2017; Aljundi et al., 2018; Zenke et al., 2017; Aljundi et al., 2019a; Jha et al., 2023) constrain updates to important weights to preserve past knowledge.
+Meanwhile, dynamic networks (Wang et al., 2022a; b; Zhou et al., 2022a; Wang et al., 2024a; McDonnell et al., 2024; Liang & Li, 2024; Wang et al., 2022e; d; Smith et al., 2023; Wang et al., 2022c) adjust their architecture on the fly, balancing the acquisition of new information with the retention of old knowledge.
 Continual Learning of CLIP.
-Continual learning of CLIP [24, 65] aims to enable PTMs to sequentially learn across diverse domains while retaining their pre-trained generalization capabilities for previously seen tasks.
-[17] continually adapt CLIP to temporal distribution shifts via timestamped data updates.
-The approach introduced in ZSCL [66] addresses the challenge of preserving zero-shot capabilities while adapting to new tasks, mitigating the risk of catastrophic forgetting.
-Subsequent works [61, 45, 58] have focused primarily on continual adaptation to downstream tasks while leveraging pre-trained predictions for unseen data.
-Recently, [58] proposed a more challenging evaluation setting, where the labels from different domains are mixed during testing.
+Continual learning of CLIP (Jha et al., 2024; Zhang et al., 2024b) aims to enable PTMs to sequentially learn across diverse domains while retaining their pre-trained generalization capabilities for previously seen tasks.
+(Garg et al., 2023) continually adapt CLIP to temporal distribution shifts via timestamped data updates.
+The approach introduced in ZSCL (Zheng et al., 2023) addresses the challenge of preserving zero-shot capabilities while adapting to new tasks, mitigating the risk of catastrophic forgetting.
+Subsequent works (Yu et al., 2024; Tang et al., 2025; Xu et al., 2024; Huang et al., 2025) have focused primarily on continual adaptation to downstream tasks while leveraging pre-trained predictions for unseen data.
+Recently, (Xu et al., 2024) proposed a more challenging setting, where the labels from different domains are mixed during testing.
 Low-Rank Adaptation.
-Low-Rank Adaptation (LoRA) [21] has gained popularity as a parameter-efficient method for fine-tuning large pre-trained models.
-Similar to LoRA, [56] learns a small set of task-specific representation vectors at each layer.
-Building on LoRA’s framework, recent works enhance LoRA by reformulating parameter updates using SVD, focusing on initialization from the SVD of pre-trained weights [63, 38], treating each rank independently [12, 64, 32], or employing a mixture of ranks [55].
+Low-Rank Adaptation (LoRA) (Hu et al., 2021) has gained popularity as a parameter-efficient method for fine-tuning large pre-trained models.
+Similar to LoRA, (Wu et al., 2024b) learns a small set of task‐specific representation vectors at each layer.
+Building on LoRA’s framework, recent works enhance LoRA by reformulating parameter updates, focusing on initialization based on the pre-trained weights (Zhang et al., 2024a; Meng et al., 2024), treating each rank independently (Ding et al., 2023; Zhang et al., 2023; Liu et al., 2024), or employing a mixture of subspaces (Wu et al., 2024a).
 These advancements highlight LoRA’s effectiveness in efficiently fine-tuning large models while preserving generalization.
-Figure 2: Overview of CoDyRA: we propose the dynamic rank-selection LoRA that allows the each pre-trained weight matrices to adaptively insert the necessary ranks for downstream adaptation while retaining pre-trained capabilities.
-After training on each task, the dynamic ranks of the parameter updates are merged back into the pre-trained weights, without incurring additional inference overhead.
 3 Methodology
@@
 Transformers.
-A typical transformer model is composed of stacked blocks, where each block consists of two main submodules: an attention (Attn) module utilizing multi-head attention (MHA) and a multilayer perceptron (MLP).
-In each transformer layer, the attention module at the $l$-th layer is defined as: $\text{Attn}(x)=\text{MHA}(x\mathbf{W}^{\text{Q}}_{l},x\mathbf{W}^{\text{K}}_{l},x\mathbf{W}^{\text{V}}_{l})\mathbf{W}^{\text{O}}_{l},$ where $\text{MHA}(\cdot)$ denotes the multi-head attention operation.
-The input sequence is denoted as $x\in\mathbb{R}^{n\times d}$, where $n$ is the sequence length and $d$ is the hidden dimension.
+A typical transformer model is composed of stacked blocks, each containing two main submodules: an multi-head attention (MHA) module (Attn) and a multilayer perceptron (MLP).
+The attention module at the $l$-th layer is defined as: $\text{Attn}(x)=\text{MHA}(x\mathbf{W}^{\text{Q}}_{l},x\mathbf{W}^{\text{K}}_{l},x\mathbf{W}^{\text{V}}_{l})\mathbf{W}^{\text{O}}_{l}.$ The input sequence is denoted as $x\in\mathbb{R}^{n\times d}$, where $n$ is the sequence length and $d$ is the hidden dimension.
 The matrices $\mathbf{W}^{\text{Q}}_{l},\mathbf{W}^{\text{K}}_{l},\mathbf{W}^{\text{V}}_{l},\mathbf{W}^{\text{O}}_{l}\in\mathbb{R}^{d\times d}$ correspond to the Query, Key, Value, and Output projection matrices, respectively.
-The MLP module within each transformer block consists of two linear transformations separated by an activation function $\text{Act}(\cdot)$, such as ReLU or GELU [20].
-The MLP module at the $l$-th transformer layer is defined as: $\text{MLP}(x)=\text{Act}(x\mathbf{W}^{\text{FC}}_{l}+b^{\text{FC}}_{l})\mathbf{W}^{\text{Proj}}_{l}+b^{\text{Proj}}_{l},$ where $\mathbf{W}^{\text{FC}}_{l}\in\mathbb{R}^{d\times d_{m}}$ and $\mathbf{W}^{\text{Proj}}_{l}\in\mathbb{R}^{d_{m}\times d}$ are the weight matrices, with $d_{m}$ being the hidden dimension of the MLP.
-Vision-language CLIP model.
-In CLIP model with joint vision-language embedding [42, 22, 14, 28], both the image encoder $f_{\theta}$ and text encoder $g_{\psi}$ are implemented as transformers in our case.
-Given an input image ${\mathbf{x}}\in\mathbb{R}^{H\times W\times C}$, it is segmented into patches and processed by the vision encoder to produce latent visual features ${\mathbf{z}}^{\text{V}}=f_{\theta}({\mathbf{x}})\in\mathbb{R}^{D}$.
-For text, the class label $y$ is embedded into prompts like “a photo of a [CLS]”, where [CLS] is the class name, resulting in text input ${\mathbf{t}}$.
-This prompt is encoded by the text encoder, yielding ${\mathbf{z}}^{\text{T}}=g_{\psi}({\mathbf{t}})\in\mathbb{R}^{D}$.
-During inference, the probability of classifying image ${\mathbf{x}}$ into class $y_{i}\in\{1,\dots,C\}$ is computed using the softmax function as: $p(y_{i}|{\mathbf{x}})=\frac{\exp(\mathtt{sim}({\mathbf{z}}^{\text{V}},{\mathbf{z}}^{\text{T}}{y_{i}})/\tau)}{\sum_{c=1}^{C}\exp(\mathtt{sim}({\mathbf{z}}^{\text{V}},{\mathbf{z}}^{\text{T}}_{y_{c}})/\tau)},$ where $\mathtt{sim}(\cdot)$ is cosine similarity, and $\tau$ is the temperature.
-CL of CLIP.
-We focus on CL of CLIP with multiple domain data, e.g., multi-domain task-incremental learning (MTIL) [66, 61] and cross-domain task-agnostic incremental learning (X-TAIL) [58].
-In the MTIL setting, the model sequentially learns from a set of $T$ tasks.
-For each task $t$ , the dataset is represented as $\mathcal{D}^{t}=\{({\mathbf{x}}^{t}_{i},y^{t}_{i})\}_{i=1}^{N^{t}}$, where ${\mathbf{x}}^{t}_{i}\in\mathbb{R}^{H\times W\times C}$ is the input image, $y^{t}_{i}\in\mathcal{C}^{t}$ is its corresponding class label, and $N^{t}$ is the number of samples in the task.
-The category set $\mathcal{C}^{t}=\{y^{t}_{j}\}_{j=1}^{M^{t}}$ represents the class labels for task $t$, consisting of $M^{t}$ classes in total.
-In the X-TAIL setting, the category set during testing includes classes from both previously seen and unseen domains, denoted as $\mathcal{C}=\mathcal{C}_{\text{seen}}\cup\mathcal{C}_{\text{unseen}}$.
-Here, $\mathcal{C}_{\text{seen}}=\bigcup_{i=1}^{n}\mathcal{C}_{i}$ includes all classes learned during prior tasks, while $\mathcal{C}_{\text{unseen}}$ represents novel classes that were not encountered during training.
-This poses a more challenging scenario, as the model must classify test images across a broader range of domains.
+The MLP in each transformer block consists of two linear layers separated by an activation function $\text{Act}(\cdot)$, such as ReLU or GELU (Hendrycks & Gimpel, 2016).
+The MLP module at the $l$-th transformer layer is defined as: $\text{MLP}(x)=\text{Act}(x\mathbf{W}^{\text{FC}}_{l}+b^{\text{FC}}_{l})\mathbf{W}^{\text{Proj}}_{l}+b^{\text{Proj}}_{l},$ where $\mathbf{W}^{\text{FC}}_{l}\in\mathbb{R}^{d\times d_{m}}$ and $\mathbf{W}^{\text{Proj}}_{l}\in\mathbb{R}^{d_{m}\times d}$ are the weight matrices, with $d_{m}$ being the hidden dimension.
+Figure 2: Overview of CoDyRA: we propose dynamic rank-selection LoRA, enabling each pre-trained weight matrix to adaptively add necessary ranks for downstream adaptation while retaining pre-trained capabilities.
+After each task, dynamic rank updates are merged into the pre-trained weights with no inference overhead.
+CLIP model.
+In CLIP (Radford et al., 2021; Ilharco et al., 2021; Fang et al., 2024; Li et al., 2022), both the image encoder $f_{\theta}$ and text encoder $g_{\psi}$ are transformers that learn a joint vision–language embedding.
+Given an image ${\mathbf{x}}\in\mathbb{R}^{H\times W\times C}$, the vision encoder produces features ${\mathbf{z}}_{\text{V}}=f_{\theta}({\mathbf{x}})\in\mathbb{R}^{D}$.
+A class label $y$ is embedded into a prompt such as “a photo of a [CLS]” to form text input $\mathbf{t}$}, which the text encoder maps to ${\mathbf{z}}_{\text{T}}=g_{\psi}({\mathbf{t}})\in\mathbb{R}^{D}$.
+At inference, the probability of classifying $\mathbf{x}$} into class $y_{i}\in\{1,\dots,C\}$ is: $p(y_{i}|{\mathbf{x}})=\frac{\exp(\mathtt{sim}({\mathbf{z}}_{\text{V}},{\mathbf{z}}_{\text{T}}^{y_{i}})/\tau)}{\sum_{c=1}^{C}\exp(\mathtt{sim}({\mathbf{z}}_{\text{V}},{\mathbf{z}}_{\text{T}}^{y_{c}})/\tau)},$ where $\mathtt{sim}(\cdot)$ is cosine similarity and $\tau$ is a temperature parameter.
+Figure 3: Task adaptation and zero-shot capability retention after training CLIP with different LoRA insertion points and ranks.
+Shapes indicate encoders, colors denote transformer modules, and sizes reflect rank values.
+Figure 4: Impact of removing trained LoRA modules on downstream task adaptation (Aircraft, left y-axis) and retention of existing capabilities (ImageNet-1k, right y-axis).
+“T” and “V” indicate text-encoder and vision-encoder modules, respectively.
+CL with CLIP.
+We study continual learning of CLIP with multi-domain data, e.g., multi-domain task-incremental learning (MTIL) (Zheng et al., 2023; Yu et al., 2024) and cross-domain task-agnostic incremental learning (X-TAIL) (Xu et al., 2024).
+In MTIL, the model sequentially learns from $T$ tasks, where each task $t$ has dataset $\mathcal{D}^{t}=\{({\mathbf{x}}^{t}_{i},y^{t}_{i})\}_{i=1}^{N^{t}}$, with input image ${\mathbf{x}}^{t}_{i}\in\mathbb{R}^{H\times W\times C}$, label $y^{t}_{i}\in\mathcal{C}^{t}$, and $M^{t}$ classes in $\mathcal{C}^{t}=\{y^{t}_{j}\}_{j=1}^{M^{t}}$.
+In X-TAIL, the test-time category set includes both seen and unseen domains, $\mathcal{C}=\mathcal{C}_{\text{seen}}\cup\mathcal{C}_{\text{unseen}}$, where $\mathcal{C}_{\text{seen}}=\bigcup_{i=1}^{n}\mathcal{C}_{i}$ contains all classes from prior tasks and $\mathcal{C}_{\text{unseen}}$ denotes novel classes never encountered.
 3.2 Analyses on LoRA Location and Rank
-Figure 3: Evaluations of downstream task adaptation and zero-shot capability retention when training CLIP with varying LoRA insertion points and ranks.
-Shapes indicate encoders, colors denote transformer modules, and sizes reflect rank values.
-Figure 4: Analysis of the impact of removing trained LoRA modules on downstream task adaptation (Aircraft) and retention of pre-trained capabilities (ImageNet-1k).
-To achieve efficient learning, we explore LoRA and aim to optimize its usage strategy.
-To analyze the impact of LoRA’s rank and placement on downstream learning and forgetting, we conduct a series of experiments: (1) We apply LoRA components with varying ranks at different locations in the pre-trained CLIP model (Fig.
-4); (2) We apply LoRA across all PTM weights and systematically prune them for evaluation (Fig.
+To enable efficient and natural continual updating of PTMs, we employ LoRA and analyze its learning and forgetting capability for optimizing its usage.
+We examine how LoRA’s rank and placement affect downstream learning and forgetting through two types of analyses (Pearl, 2022; Meng et al., 2022), providing a comprehensive view of their overall impact.
+(1) We first apply LoRA components with varying ranks at different locations in the pre-trained CLIP, for measurement of importance or contribution (Fig.
+4); (2) We apply LoRA across all PTM weights and systematically prune them for evaluation, for testing the standalone capability (Fig.
 4).
-We train the model on downstream datasets and assess its performance on both (1) the new task and (2) an additional reference dataset to evaluate the preservation of pre-trained capabilities.
+We train the model on downstream datasets and assess its performance on both (1) the new task, i.e., learning new knowledge, and (2) an additional reference dataset to evaluate the retention and forgetting of existing capabilities.
+We achieve three main results to guide methodology design.
+Without loss of generality, the learning–forgetting analyses can be extended to each learning step in straightforward, assumption-free CL.
 Applying LoRA on all weights.
+Unlike prior works (Hu et al., 2021; Zhang et al., 2024a; Meng et al., 2024) that only apply LoRA to limited components by default, we first examine the impact of different LoRA placements.
 As shown in Fig.
 4 and Fig.
-4, applying LoRA to specific modules alone may limit learning capacity in different ways: (1) The different LoRA placement locations significantly affect downstream performance and the forgetting/retention of pre-trained capabilities; (2) Updating only the vision encoder leads to more forgetting of pre-trained capabilities; (3) Updating only the attention modules helps mitigate forgetting but limits the learning ability on new tasks.
-These observations align with the distinct roles of different PTM components.
-Rather than manually selecting LoRA placement, we apply it across all weights and optimize for an adaptive configuration.
-The effects of different LoRA ranks on balancing learning and forgetting.
-From Fig.
-4, we can observe that, despite learning low-rank space, a relatively higher-rank LoRA facilitates learning new tasks but tends to increase forgetting, while lower-rank LoRA mitigates forgetting but limits adaptation.
-This suggests that a lower-rank LoRA produces stronger regularization in maintaining the learned knowledge by sacrificing the learning of new knowledge.
-Importantly, we identify a balance at a moderately small rank, which reduces forgetting while maintaining adaptability.
-This implies that minimizing the rank while fitting on new data may find an optimized trade-off for CL.
+4, applying LoRA to only specific modules alone may limit learning capacity in different ways: (1) LoRA placement strongly impacts downstream performance and retention of pre-trained capabilities; (2) Updating only the vision encoder causes more forgetting; (3) Updating only attention modules mitigates forgetting but limits learning on new tasks.
+These observations reflect the distinct roles of PTM components.
+Takeaway 1: Instead of manual or fixed LoRA placement at specific parameters of PTM like previous methods (Hu et al., 2021; Liang & Li, 2024; Yu et al., 2024; Zhang et al., 2024a; Meng et al., 2024), we apply LoRA to all weights and optimize for an adaptive configuration.
+The effects of LoRA ranks on balancing learning and forgetting.
+We identified that the rank number of LoRA plays a critical role in balancing learning and forgetting.
+As shown in Fig.
+4, we can observe that a relatively higher-rank LoRA facilitates learning new tasks but tends to increase forgetting, while a lower-rank LoRA mitigates forgetting but limits adaptation.
+This suggests that lower-rank LoRA interferes less with existing knowledge (i.e., reduces forgetting) while learning less new knowledge.
+Fortunately, we identify a balance (top-right in Fig.
+4), where a moderately small rank reduces forgetting while enabling enough adaptability.
+During training, low-rank parameter updates tend to enable the updated model to be closer to the original model, with less interference and forgetting.
+In the extreme case, a rank-zero LoRA learns nothing and introduces no forgetting.
+Thus, minimizing LoRA rank for each new task in CL can reduce forgetting without extra assumptions.
+On the other hand, the task-learning objective emphasizes new knowledge acquisition, enabling a balance.
+Takeaway 2: A plasticity–stability balance exists and associates with LoRA rank (high vs.
+low), which can be adaptively achieved by jointly optimizing the task objective and minimizing LoRA ranks.
 The balance varies by location and task.
-The ranks for a good balance vary by placement (Fig.
-4), and pruning LoRA at different locations has varying effects (Fig.
-4), indicating that not all LoRA components contributed equally for a task.
-For example, removing updates in the Value projection of the Attention module in the vision encoder minimally affected adaptation but significantly restored pre-trained performance.
-Interestingly, pruning the MLP layer of the vision encoder slightly reduced task performance but enhanced pre-trained capabilities.
-These findings motivate our adaptive approach to dynamically determine ranks based on location and task.
-3.3 Proposed CoDyRA
+Fig.
+4 shows that the optimal rank for balance depends on both parameter placement and the learning task, consistent with the pruning-based sufficiency analysis in Fig.
+4.
+For example, removing updates to the Value projection in the Attention module of the vision encoder had little impact on adaptation but substantially restored pre-trained performance.
+Interestingly, pruning the MLP layer slightly reduced task performance while enhancing pre-trained capabilities.
+Takeaway 3: The learning–forgetting balance point tied to rank varies across modules and tasks, necessitating adaptive optimization.
+3.3 Proposed CoDyRA with Sparsity-Induced Rank Selection and Minimization
 Overview.
 Our analyses in Sec.
-3.2 suggest applying LoRA across all CLIP modules while adaptively selecting an appropriate rank for each task and location to balance learning and forgetting.
-While finding an ideally global optimal rank is challenging, we approach it by jointly optimizing downstream task performance and minimizing the number of ranks using sparsity-promoting regularization.
-We introduce dynamic rank-selective parameter updates for all pre-trained weight matrices within the Attention and MLP modules of both the vision and text encoders, as shown in Fig.
-2.
-These updates are adaptively inserted based on the importance of each rank for the current training data.
-For clarity, during training on task $t$, we denote the pre-trained weights—or those updated from the previous tasks—as $\{\mathbf{W}^{t,m}_{0}\}_{m=1}^{M}$, where $M$ represents the total number of pre-trained weight matrices augmented with LoRA.
-After training on each task $t$, these dynamic rank updates are merged back into the original pre-trained weights.
-Low-rank updates with rank importance.
-For efficient adaptation, instead of directly updating a pre-trained weight matrix $\mathbf{W}_{0}^{t,m}\in\mathbb{R}^{d\times k}$, LoRA [21] introduces two low-rank matrices, $\mathbf{B}^{t,m}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}^{t,m}\in\mathbb{R}^{r\times k}$, such that $\Delta\mathbf{W}=\mathbf{B}^{t,m}\mathbf{A}^{t,m}$, where $r\ll\min(d,k)$.
+3.2 suggest applying LoRA across all CLIP modules while adaptively optimizing for an appropriate rank for each specific task and location to balance learning and forgetting.
+Together with the task learning objective, minimizing active LoRA ranks acts as a regularizer, keeping the updated model closer to its previous state and alleviating forgetting.
+However, the balanced rank between learning and forgetting is not consistent across parameters or tasks (Takeaway 3).
+We address this with a sparsity-promoting regularization for the LoRA adapters.
+We introduce adaptive rank-selective updates for all pre-trained weight matrices in the Attention and MLP modules of both vision and text encoders (Fig.
+2).
+The sparsity-promoting regularization enables adaptive rank selection and minimization for each LoRA adapter on each task.
+For each task in CL, beyond adaptive rank minimization, CoDyRA requires no additional priors or assumptions and does not store or maintain information (e.g., data, parameters, or statistics) from previous steps, unlike prior methods (Zheng et al., 2023; Yu et al., 2024; Xu et al., 2024).
+During training on task $t$, we denote the pre-trained weights, or those updated from previous tasks, as $\{\mathbf{W}^{t,m}_{0}\}_{m=1}^{M}$, where $M$ is the total number of weight matrices updated by LoRA.
+After each task, the dynamic rank updates are merged back into the original weights.
+Regularized Low-rank updates with rank importance.
+To update weight matrix $\mathbf{W}_{0}^{t,m}\in\mathbb{R}^{d\times k}$, LoRA (Hu et al., 2021) introduces two low-rank matrices, $\mathbf{B}^{t,m}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}^{t,m}\in\mathbb{R}^{r\times k}$, such that $\Delta\mathbf{W}^{t,m}=\mathbf{B}^{t,m}\mathbf{A}^{t,m}$, where $r\ll\min(d,k)$.
 The updated weight matrix is then defined as:
 $\mathbf{W}^{t,m}=\mathbf{W}_{0}^{t,m}+\Delta\mathbf{W}^{t,m}=\mathbf{W}_{0}^{t,m}+\mathbf{B}^{t,m}\mathbf{A}^{t,m}.$ (1)
-In this formulation, the original weights $\mathbf{W}_{0}^{t,m}$ remain fixed, and only $\mathbf{B}^{t,m}$ and $\mathbf{A}^{t,m}$ are trained, reducing the number of trainable parameters to $r(d+k)$.
-Notably, LoRA incurs no additional computational overhead during inference, as the low-rank updates can be efficiently merged into the pre-trained weights, as shown in Eq.
-(1).
-To automatically and adaptively decide the rank in the LoRA updates, we introduce a learnable importance weight vector $\mathbf{w}^{t,m}\in\mathbb{R}^{r}$ and incorporate it into the formulation of LoRA as follows:
-$\Delta\mathbf{W}^{t,m}=\sum\nolimits_{i=1}^{r}\mathbf{w}_{i}^{t,m}\mathbf{B}_{i}^{t,m}\mathbf{A}_{i}^{t,m},$ (2)
-where $\mathbf{w}^{t,m}$ is learnable and represents the importance of each rank.
-The formulation is based on a reformulation of LoRA updates in Eq.
-(1) based on Singular Value Decomposition (SVD) [64, 12, 32, 38, 63].
-The importance weights are corresponding to the diagonal elements in the singular value matrix of SVD.
-Introducing importance weights incurs minimal overhead, adding only $r$ additional parameters per module.
-For initialization, we follow the standard convention: matrix $\mathbf{A}$ is randomly initialized, matrix $\mathbf{B}$ is set to zero, and the importance weights $\mathbf{w}$ are randomly initialized.
-Table 1: Comparisons on X-TAIL for each domain in terms of “Transfer”, “Average”, and “Last” scores (%).
+Only $\mathbf{B}^{t,m}$ and $\mathbf{A}^{t,m}$ are trained for each task and then merged into the weights. LoRA incurs no additional computational overhead during inference.
+As shown in Sec.
+3.2, minimizing the active rank for each LoRA adapter alleviates forgetting, which can be formulated as minimizing a regularizer $R(\Delta\mathbf{W}^{t,m})\doteq R(\mathbf{B}^{t,m}\mathbf{A}^{t,m})$.
+The minimum of $R(\Delta\mathbf{W}^{t,m})$ is zero, i.e., $\min R(\Delta\mathbf{W}^{t,m})=0$, corresponding to a rank-zero matrix—no update to $\mathbf{W}^{t,m}$ and hence no forgetting.
+By jointly optimizing the new-task learning objective and the minimized $R(\Delta\mathbf{W}^{t,m})$, we achieve a balance between learning and forgetting.
+To automatically and adaptively minimize the active ranks in LoRA, we formulate $R(\Delta\mathbf{W}^{t,m})$ through a rank-selection LoRA.
+We introduce a learnable importance weight vector $\mathbf{w}^{t,m}\in\mathbb{R}^{r}$ to indicate the active ranks for a LoRA adapter:
+$\Delta\mathbf{W}^{t,m}=\sum\nolimits_{i=1}^{r}\mathbf{w}_{i}^{t,m}\mathbf{B}_{:,i}^{t,m}\mathbf{A}_{i,:}^{t,m}$ (2)
+where $\mathbf{w}_{i}^{t,m}$ is the learnable importance weight associated with rank $i$, $\mathbf{B}_{:,i}^{t,m}$ denotes the $i$-th column of $\mathbf{B}^{t,m}$, and $\mathbf{A}_{i,:}^{t,m}$ denotes the $i$-th row of $\mathbf{A}^{t,m}$.
+Minimizing $R(\Delta\mathbf{W}^{t,m})$ can be achieved through optimizing $\mathbf{w}^{t,m}$.
+Adaptive rank-selective updates with sparse-prompting regularization.
+The formulation in Eq.
+(2) enables the model to dynamically optimize the importance of each rank and minimize active ranks through gradient descent.
+[^1] To align the training process with the pre-trained model, we use CLIP’s contrastive loss as the optimization objective during pre-training.
+To only retain the essential ranks for learning new tasks (thus reducing forgetting), we can achieve optimizing $R(\Delta\mathbf{W}^{t,m})$ through an $\ell_{1}$ norm-based sparsity-promoting regularization on $\mathbf{w}^{t,m}$, i.e., $\|\mathbf{w}^{t,m}\|_{1}$.
+The optimization objective for parameters $\{\mathbf{w}^{t,m},\mathbf{B}^{t,m},\mathbf{A}^{t,m}\}_{m=1}^{M}$ learned on task $t$ is:
+$\mathcal{L}_{\text{train}}^{t}:={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{L}_{\text{sup}}^{t}}}+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\lambda\sum\nolimits_{m=1}^{M}\|\mathbf{w}^{t,m}\|_{1}}},$ (3)
+where $\mathcal{L}_{\text{sup}}^{t}$ is the supervised training loss for current task, $M$ is the number of weight matrices inserted $r$-rank update, and $\mathbf{w}^{t,m}$ represents the importance weights of each rank added to the weight matrix $\mathbf{W}^{t,m}_{0}$.
+The $\ell_{1}$ regularization strength is controlled by $\lambda$.
+To handle the non-differentiable $\ell_{1}$ regularization applied to the importance weights, we adopt the proximal gradient method (Beck & Teboulle, 2009).
+The $i$-th element of $\hat{\mathbf{w}}^{t,m}$ for rank $i$ is updated via soft-thresholding:
+$\mathbf{w}_{i}^{t,m}:=\mathbbm{1}(\lvert\hat{\mathbf{w}}^{t,m}_{i}\rvert>\kappa)\cdot(\hat{\mathbf{w}}^{t,m}_{i}+\mathrm{sign}(\hat{\mathbf{w}}^{t,m}_{i})\cdot\kappa),$ (4)
+where $\hat{\mathbf{w}}^{t,m}_{i}$ denotes the value of $\mathbf{w}^{t,m}_{i}$} after applying the gradient update from the supervised loss $\mathcal{L}_{\text{sup}}^{t}$.
+The threshold $\kappa$ increases from zero to $\kappa_{\text{max}}$, analogous to linearly scheduling $\lambda$, and adaptively induces sparsity in $\mathbf{w}$ based on the task learning objective.
+The indicator $\mathbbm{1}(\cdot)$ returns 1 if the condition holds and 0 otherwise, while $\mathrm{sign}(\cdot)$ gives the input’s sign ($\pm$).
+More details are left in Appendix A.
+This approach preserves only the ranks with significant importance for learning the current task and prunes those with low relevance, adapting across weights and tasks.
+The importance weights on preserved ranks softly highlight their relative significance.
+Early in training, dense updates are applied without Eq.
+(4), allowing all ranks to capture task-relevant information before sparsity is enforced.
+Continually learn on new tasks.
+CoDyRA views CL as a sequence of tasks with LoRA, without storing past data Zheng et al.
+(2023), task information Yu et al.
+(2024); Xu et al.
+(2024), or relying on assumptions (Liang & Li, 2024).
+For each new task, new CoDyRA modules are initialized and optimized; low-importance ranks are pruned, and non-zero ranks merged into the weights.
+Sparsity-promoting regularization controls task capacity, keeping the model close to its previous state and reducing forgetting.
+This enables efficient continual updating with only sparsity-based regularization.
+Table 1: Comparisons on X-TAIL for each domain.
 The best and the second best results are highlighted in red and blue, respectively.
-Methods marked with $\dagger$ indicate the use of domain prediction or distribution detection [61, 58].
-    Method Aircraft [35] Caltech101 [15] DTD [8] EuroSAT [19] Flowers [40] Food [5] MNIST [11] OxfordPet [41] Cars [26] SUN397 [57] Average
+Methods marked with $\dagger$ indicate the use of domain prediction or distribution detection (Yu et al., 2024; Xu et al., 2024).
+Method
+Aircraft Caltech101 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
 CLIP
@@
 Transfer
-   Zero-shot [42] 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
-   LwF [29] 66.6 26.9 19.5 51.0 78.4 26.6 68.9 35.5 56.1 47.7
-   WiSE-FT [54] 70.1 31.9 25.3 56.3 79.8 29.9 74.9 45.6 56.8 52.3
-   iCaRL [43] 71.7 35.0 43.0 63.4 86.9 43.9 87.8 63.7 60.0 61.7
-   ZSCL [66] 73.3 32.6 36.8 62.1 83.8 42.1 83.6 56.5 60.2 59.0
+   Zero-shot (Radford et al., 2021) 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
+   LwF (Li & Hoiem, 2017) 66.6 26.9 19.5 51.0 78.4 26.6 68.9 35.5 56.1 47.7
+   WiSE-FT (Wortsman et al., 2022) 70.1 31.9 25.3 56.3 79.8 29.9 74.9 45.6 56.8 52.3
+   iCaRL (Rebuffi et al., 2017b) 71.7 35.0 43.0 63.4 86.9 43.9 87.8 63.7 60.0 61.7
+   ZSCL (Zheng et al., 2023) 73.3 32.6 36.8 62.1 83.8 42.1 83.6 56.5 60.2 59.0
 […]
@@
 Average
-   LwF [29] 24.7 79.7 38.3 36.9 63.9 81.0 36.5 71.9 42.7 56.7 53.2
-   WiSE-FT [54] 27.1 76.5 40.9 31.3 68.7 81.6 31.4 74.7 51.7 58.4 54.2
-   iCaRL [43] 25.4 72.1 37.5 51.6 65.1 87.1 59.1 88.0 63.7 60.1 61.0
-   ZSCL [66] 36.0 75.0 40.7 40.5 71.0 85.3 46.3 83.3 60.7 61.5 60.0
+   LwF (Li & Hoiem, 2017) 24.7 79.7 38.3 36.9 63.9 81.0 36.5 71.9 42.7 56.7 53.2
+   WiSE-FT (Wortsman et al., 2022) 27.1 76.5 40.9 31.3 68.7 81.6 31.4 74.7 51.7 58.4 54.2
+   iCaRL (Rebuffi et al., 2017b) 25.4 72.1 37.5 51.6 65.1 87.1 59.1 88.0 63.7 60.1 61.0
+   ZSCL (Zheng et al., 2023) 36.0 75.0 40.7 40.5 71.0 85.3 46.3 83.3 60.7 61.5 60.0
 […]
@@
 Last
-   LwF [29] 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
-   WiSE-FT [54] 21.8 76.8 42.9 20.8 77.5 84.9 30.7 76.6 75.8 72.5 58.0
-   iCaRL [43] 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
-   ZSCL [66] 33.1 75.3 43.5 35.2 74.6 87.4 50.4 84.2 77.3 73.4 63.4
+   LwF (Li & Hoiem, 2017) 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
+   WiSE-FT (Wortsman et al., 2022) 21.8 76.8 42.9 20.8 77.5 84.9 30.7 76.6 75.8 72.5 58.0
+   iCaRL (Rebuffi et al., 2017b) 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
+   ZSCL (Zheng et al., 2023) 33.1 75.3 43.5 35.2 74.6 87.4 50.4 84.2 77.3 73.4 63.4
 […]
@@
 […]
-Dynamic rank-selective updates with sparse rank importance.
-The formulation in Eq.
-(2) enables the model to dynamically adjust the importance of each rank through gradient descent.
-[^1] To align the training process with the pre-trained model, we use CLIP’s contrastive loss as the optimization objective during pre-training.
-However, not all ranks contribute equally to downstream task performance (as shown in Fig.
-4).
-To ensure that only the most impactful ranks are retained for learning on new task while removing unnecessary ranks to mitigate forgetting of already-learned knowledge, we propose optimizing the importance weights with a sparsity-promoting regularization through minimizing the $\ell_{1}$ norm.
-The optimization objective for parameters $\{\mathbf{w}^{t,m},\mathbf{B}^{t,m},\mathbf{A}^{t,m}\}_{m=1}^{M}$ learned on task $t$ is:
-$\mathcal{L}_{\text{train}}^{t}:=\mathcal{L}_{\text{sup}}^{t}+\lambda\sum\nolimits_{m=1}^{M}\|\mathbf{w}^{t,m}\|_{1},$ (3)
-where $\mathcal{L}_{\text{sup}}^{t}$ is the supervised training loss, $M$ is the total number of pre-trained weight matrices inserted $r$-rank update, and $\mathbf{w}^{t,m}$ represents the importance weights of each rank added to the weight matrix $\mathbf{W}^{t,m}_{0}$.
-The $\ell_{1}$ regularization strength is controlled by $\lambda$.
-To handle the non-differentiable $\ell_{1}$ regularization applied to the importance weights, we adopt the proximal gradient method [4].
-The $i$-th element of $\hat{\mathbf{w}}^{t,m}$ for rank $i$ is updated via soft-thresholding:
-$\mathbf{w}_{i}^{t,m}:=\mathbbm{1}(\lvert\hat{\mathbf{w}}^{t,m}_{i}\rvert>\kappa)\cdot(\hat{\mathbf{w}}^{t,m}_{i}+\mathrm{sign}(\hat{\mathbf{w}}^{t,m}_{i})\cdot\kappa),$ (4)
-where $\hat{\mathbf{w}}^{t,m}_{i}$ denotes the value of ${\mathbf{w}^{t,m}_{i}}$ after applying the gradient update from the supervised loss $\mathcal{L}_{\text{sup}}^{t}$.
-The threshold parameter $\kappa$ is initially set to zero and gradually increases to a maximum value $\kappa_{\text{max}}$, analogous to linearly scheduling the regularization strength $\lambda$.
-The indicator function $\mathbbm{1}(\cdot)$ returns 1 if the condition is met and 0 otherwise, while $\mathrm{sign}(\cdot)$ represents the +/+/-+ / - sign of the input.
-This approach preserves only the ranks with significant importance, pruning those with low relevance.
-By iteratively adjusting weightings, the model focuses dynamic parameter updates on the most relevant components.
-During the early stages of training, dense updates are performed without Eq.
-(4) to allow all ranks to capture task-relevant information before applying sparsity constraints.
-Continually enhancing the pre-trained model.
-After training, ranks with zero importance weights are pruned, while non-zero ranks are merged into the pre-trained weights by substituting Eq.
-(2) into the formulation Eq.
-(1).
-When a new task is introduced, a fresh set of CoDyRA modules is initialized, and computations involve only these newly initialized CoDyRA modules alongside the continually updated pre-trained model.
-This approach enables the model to efficiently leverage accumulated knowledge while maintaining a minimal computational footprint, ensuring no additional overhead.
-Table 2: Comparisons on 5-shot MTIL setting, evaluated in a similar manner as in Table 1.
-    Method Aircraft [35] Caltech101 [15] CIFAR100 [27] DTD [8] EuroSAT [19] Flowers [40] Food [5] MNIST [11] OxfordPet [41] Cars [26] SUN397 [57] Average
+4 Experiments
+4.1 Experimental Setting
+Datasets.
+We evaluate our method on two widely used continual learning settings for vision-language models: MTIL (Zheng et al., 2023; Yu et al., 2024) and X-TAIL (Xu et al., 2024).
+A wide range of fine-grained datasets is used, where each dataset is treated as an individual task to be incrementally learned.
+We leave detailed experimental settings in Appendix B.
+Evaluation Metrics.
+We use the same metrics as prior work (Zheng et al., 2023; Yu et al., 2024; Xu et al., 2024): “Transfer” for zero-shot performance on unseen data, “Last” for retention of earlier tasks, and “Average” for mean accuracy across all datasets of all learning tasks.
+Implementation Details.
+We follow the setups in (Zheng et al., 2023; Yu et al., 2024; Xu et al., 2024) using CLIP with a ViT-B/16 backbone (Radford et al., 2021).
+By default, CoDyRA is applied to all pre-trained weight matrices in the vision and text encoders with an initial rank of 16.
+Each task is trained for 500 iterations with AdamW (Loshchilov & Hutter, 2017), using dense training ratios of 0.5 for X-TAIL and 0.7 for MTIL.
+$\kappa_{\text{max}}$ is set as 0.005.
+4.2 Experimental Results
+Cross-domain task-agnostic incremental learning.
+We report X-TAIL results in Table 1.
+Unlike prior methods (Yu et al., 2024; Xu et al., 2024) that rely on zero-shot predictions to preserve pre-trained capabilities, our approach continually updates the model with minimal forgetting.
+As those methods are bounded by zero-shot performance, ours surpasses this limit, achieving higher transfer accuracy.
+While continual updates may slightly affect some unseen predictions, the overall effect is positive: the model adapts to new tasks while modestly improving generalization.
+[…]
+Table 2: Average transfer accuracy on unseen datasets not encountered during CL.
+Method Average Transfer Accuracy on Unseen Data
+CIFAR100 Places365 ImageNet-1k Average
+Zero-Shot (Radford et al., 2021) 68.24 33.77 66.72 56.24
+MoE-Adapter (Yu et al., 2024) 68.24 33.77 66.72 56.24
+RAIL (Xu et al., 2024) 68.24 33.77 66.72 56.24
+CoDyRA 68.95 36.52 68.52 58.00
+Table 3: Comparisons on 5-shot MTIL setting.
+Method
+Aircraft Caltech101 CIFAR100 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
 CLIP
-   Zero-shot [42] 24.3 88.4 68.2 44.6 54.9 71.0 88.5 59.4 89.0 64.7 65.2 65.3
+   Zero-shot (Radford et al., 2021) 24.3 88.4 68.2 44.6 54.9 71.0 88.5 59.4 89.0 64.7 65.2 65.3
 Transfer
-   Zero-shot [42] 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
-   LwF [29] 72.1 49.2 35.9 44.5 41.1 66.6 50.5 69.0 19.0 51.7 50.0
-   LwF-VR [13] 82.2 62.5 40.1 40.1 56.3 80.0 60.9 77.6 40.5 60.8 60.1
-   WiSE-FT [54] 77.6 60.0 41.3 39.4 53.0 76.6 58.1 75.5 37.3 58.2 57.7
-   ZSCL [66] 84.0 68.1 44.8 46.8 63.6 84.9 61.4 81.4 55.5 62.2 65.3
+   Zero-shot (Radford et al., 2021) 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
+   LwF (Li & Hoiem, 2017) 72.1 49.2 35.9 44.5 41.1 66.6 50.5 69.0 19.0 51.7 50.0
+   LwF-VR (Ding et al., 2022) 82.2 62.5 40.1 40.1 56.3 80.0 60.9 77.6 40.5 60.8 60.1
+   WiSE-FT (Wortsman et al., 2022) 77.6 60.0 41.3 39.4 53.0 76.6 58.1 75.5 37.3 58.2 57.7
+   ZSCL (Zheng et al., 2023) 84.0 68.1 44.8 46.8 63.6 84.9 61.4 81.4 55.5 62.2 65.3
 […]
@@
 Average
-   LwF [29] 23.5 77.4 43.5 41.7 43.5 52.2 54.6 63.4 68.0 21.3 52.6 49.2
-   LwF-VR [13] 24.9 89.1 64.2 53.4 54.3 70.8 79.2 66.5 79.2 44.1 61.6 62.5
-   WiSE-FT [54] 32.0 87.7 61.0 55.8 68.1 69.3 76.8 71.5 77.6 42.0 59.3 63.7
-   ZSCL [66] 28.2 88.6 66.5 53.5 56.3 73.4 83.1 56.4 82.4 57.5 62.9 64.4
+   LwF (Li & Hoiem, 2017) 23.5 77.4 43.5 41.7 43.5 52.2 54.6 63.4 68.0 21.3 52.6 49.2
+   LwF-VR (Ding et al., 2022) 24.9 89.1 64.2 53.4 54.3 70.8 79.2 66.5 79.2 44.1 61.6 62.5
+   WiSE-FT (Wortsman et al., 2022) 32.0 87.7 61.0 55.8 68.1 69.3 76.8 71.5 77.6 42.0 59.3 63.7
+   ZSCL (Zheng et al., 2023) 28.2 88.6 66.5 53.5 56.3 73.4 83.1 56.4 82.4 57.5 62.9 64.4
 […]
@@
 Last
-   LwF [29] 22.1 58.2 17.9 32.1 28.1 66.7 46.0 84.3 64.1 31.5 60.1 46.5
-   LwF-VR [13] 22.9 89.8 59.3 57.1 57.6 79.2 78.3 77.7 83.6 60.1 69.8 66.9
-   WiSE-FT [54] 30.8 88.9 59.6 60.3 80.9 81.7 77.1 94.9 83.2 62.8 70.0 71.9
-   ZSCL [66] 26.8 88.5 63.7 55.7 60.2 82.1 82.6 58.6 85.9 66.7 70.4 67.4
+   LwF (Li & Hoiem, 2017) 22.1 58.2 17.9 32.1 28.1 66.7 46.0 84.3 64.1 31.5 60.1 46.5
+   LwF-VR (Ding et al., 2022) 22.9 89.8 59.3 57.1 57.6 79.2 78.3 77.7 83.6 60.1 69.8 66.9
+   WiSE-FT (Wortsman et al., 2022) 30.8 88.9 59.6 60.3 80.9 81.7 77.1 94.9 83.2 62.8 70.0 71.9
+   ZSCL (Zheng et al., 2023) 26.8 88.5 63.7 55.7 60.2 82.1 82.6 58.6 85.9 66.7 70.4 67.4
 […]
@@
 […]
-4 Experiments
-4.1 Experimental Setting
-Datasets.
-We evaluate our method on two widely used continual learning settings for vision-language models: MTIL [66, 61] and X-TAIL [58].
-For the MTIL setting, a total of 11 datasets are used, where each dataset is treated as an individual task to be incrementally learned.
-For the X-TAIL setting, CIFAR100 is excluded to prevent domain overlap, following the protocol in [58].
-In line with [58], we use a 5-shot split for MTIL and a 16-shot split for X-TAIL.
-Evaluation Metrics.
-We adopt the same evaluation metrics as prior works, including “Transfer”, “Average” , and “Last”.
-“Transfer”evaluates the model’s zero-shot transfer capability on unseen data.
-“Last” assesses the model’s ability to retain knowledge from earlier tasks, and“Average” provides a composite score, representing the mean performance across both the “Transfer” and “Last”.
-Implementation Details.
-We follow the experimental setups in [66, 61, 58] and use the CLIP model with a ViT-B/16 backbone [42] for all experiments.
-By default, CoDyRA is applied to all pre-trained weight matrices in both the vision and text encoders, with an initial rank of 16.
-Each task in both experimental settings is trained over 500 iterations using the AdamW optimizer [33].
-The dense training iteration ratio is set to 0.5 for X-TAIL and 0.7 for MTIL, relative to the total training iterations.
-The maximum threshold for adaptive pruning $\kappa$ is fixed at 0.005 for both settings.
-4.2 Experimental Results
-Cross-domain task-agnostic incremental learning.
-We report the X-TAIL setting evaluation results in Table 1.
-In contrast to prior methods [61, 58] , which rely on zero-shot predictions to preserve pre-trained capabilities and mitigate forgetting, our method overcomes this limitation by continually refining the model.
-Notably, while these approaches are bounded by the upper limit of zero-shot performance, our approach surpasses this limit, achieving a higher average transfer accuracy.
-Although some predictions on unseen data might be slightly affected due to the continual parameter updates, our method ultimately improves overall transfer accuracy by adapting the pre-trained model to new tasks while enhancing its generalization capabilities.
-[…]
 Multi-domain task-incremental learning.
-We further evaluate our proposed method under the few-shot MTIL setting, as shown in Table 2, following [61, 58].
-Consistent with the results observed in the X-TAIL setting, our method demonstrates clear superiority in this scenario.
-Without incorporating any additional techniques, our approach outperforms previous methods across all evaluation metrics, achieving an average transfer accuracy that surpasses the upper bound set by zero-shot predictions in prior works.
-Furthermore, by integrating techniques from previous approaches [61, 58], our method can be further enhanced, leading to even greater performance gains.
+We further evaluate our method in the few-shot MTIL setting (Table 3), following (Yu et al., 2024; Xu et al., 2024).
+As with X-TAIL, our approach outperforms prior methods across all metrics without extra techniques and even surpasses their zero-shot upper bound in average transfer accuracy.
+Incorporating techniques from (Yu et al., 2024; Xu et al., 2024) further boosts performance.
 4.3 Discussions
 Pre-trained knowledge retention and potential pre-trained model improvement.
-To understand how our method improves transfer performance beyond the upper bound set by zero-shot predictions in prior works, we evaluated on hold-out datasets never encountered during continual learning, specifically CIFAR100 [27], Places365 [67], and ImageNet-1k [10], as shown in Table 3.
-Table 3: Average transfer accuracy on unseen datasets not encountered during CL.
-Method Average Transfer Accuracy on Unseen Data
-CIFAR100 Places365 ImageNet-1k Average
-Zero-Shot [42] 68.24 33.77 66.72 56.24
-MoE-Adapter [61] 68.24 33.77 66.72 56.24
-RAIL [58] 68.24 33.77 66.72 56.24
-CoDyRA 68.95 36.52 68.52 58.00
-Unlike prior approaches [61, 58], which rely exclusively on zero-shot predictions for unseen data—thereby constraining their performance to that of the original CLIP model—our method overcomes this limitation.
-By effectively integrating new knowledge into the pre-trained parameter space while preserving existing capabilities, our approach enhances overall model performance.
-As a result, it achieves superior zero-shot transfer accuracy, surpassing the zero-shot performance limits of prior works.
-This demonstrates that CoDyRA effectively incorporates general knowledge into its knowledge base, further improving its overall representation capability.
-Figure 5: Visualization and statistical analysis of rank activation on the Aircraft dataset.
-Figure 6: Visualization and statistical analysis of rank activation on the Oxford Pets dataset.
+To assess transfer beyond the zero-shot limits of prior works, we evaluated on unseen datasets—CIFAR100 (Krizhevsky et al., 2009), Places365 (Zhou et al., 2017), and ImageNet-1k (Deng et al., 2009) (Table 2).
+Unlike methods that rely solely on zero-shot predictions (Yu et al., 2024; Xu et al., 2024) and are constrained by CLIP’s baseline, our approach integrates new knowledge into the pre-trained space while preserving existing capabilities.
+This yields superior zero-shot transfer accuracy, demonstrating that CoDyRA effectively expands general knowledge and improves overall representations.
 Rank allocation across different datasets.
-We visualize and analyze the distribution of activated ranks merged into the pre-trained weights across different datasets in Fig.
+We visualize the ranks merged into the pre-trained weights for different datasets in Fig.
 6 and Fig.
-6.
-Overall, the Aircraft dataset generally requires more rank updates in the PTM than the Pets dataset.
-Table 4: Computation cost.
-Methods
-Training
-Params.
-Additional Infer.
-Params. / Mem.
-LWF [29] 129.6M None
-ZSCL [66] 129.6M None
-MoE-Adapters [61] 59.8M 13.35M
-RAIL [58] N/A 24.18M / 9.01M
-CoDyRA 4.4M None
+6 (more in Appendix Fig.
+11 to Fig.
+17).
+The results show that CoDyRA adapts the number of active ranks differently across locations and tasks, reflecting the varying demands of each dataset.
+These strike a balance between plasticity and stability, demonstrating the method’s ability to regulate forgetting while enabling effective task learning.
 For the Aircraft dataset, we observe that a greater number of ranks are allocated to the Output projection of the Attention module within the text encoder (Fig.
@@
 These observations highlight the adaptive nature of our approach, where rank allocation dynamically adjusts based on dataset characteristics, ensuring effective domain adaptation while preserving pre-trained knowledge.
-We provide more visualizations and analyses in Appendix.
+We provide more visualizations and analyses in Appendix B.4.
+Figure 5: Visualization and statistical analysis of rank activation on the Aircraft dataset.
+Figure 6: Visualization and statistical analysis of rank activation on the Oxford Pets dataset.
+Table 4: Computation cost.
+Methods
+Training
+Params.
+Additional Infer.
+Params. / Mem.
+LWF (Li & Hoiem, 2017) 129.6M None
+ZSCL (Zheng et al., 2023) 129.6M None
+MoE-Adapters (Yu et al., 2024) 59.8M 13.35M
+RAIL (Xu et al., 2024) N/A 24.18M / 9.01M
+CoDyRA 4.4M None
 Computation cost.
-We compare the computational costs in terms of trainable parameters and additional inference requirements, as summarized in Fig.
-4.
-For MoE-Adapter, the difference between training and testing parameters arises from its mixture-of-experts design: during training, all experts are active, whereas during inference, only the top-2 experts are selected, reducing the number of active parameters.
-RAIL-Primal incurs additional inference overhead due to its use of high-dimensional feature expansion through projection and regression layers.
-Meanwhile, RAIL-DUAL significantly increases memory usage by maintaining a memory bank of image features for all training samples across 1,100 classes.
-Notably, these methods experience linearly growing computational costs as the number of tasks increases.
-In contrast, our method minimizes training overhead by selectively retaining only ranks with high importance.
-Once training is complete, these updates are merged into the pre-trained weights, ensuring zero additional memory overhead during inference.
+We compare computational costs in terms of trainable parameters and inference overhead, as shown in Table 4.
+For MoE-Adapter, the gap between training and testing parameters stems from its mixture-of-experts design: all experts are active during training, while only the top-2 are used at inference.
+RAIL-Primal introduces extra overhead from high-dimensional feature expansion via projection and regression layers.
+RAIL-Dual further increases memory use by storing image features for all samples across 1,100 classes.
+These methods all incur linearly growing costs as tasks increase.
+In contrast, our method reduces training overhead by retaining only high-importance ranks; once merged into the pre-trained weights, the updates add zero memory cost at inference.
 Figure 7: Ablations on different insertion locations and the impact of varying the initial rank.
@@
 4.4 Ablation Studies
-Effects of insertion locations and number of initial ranks.
-We examine the impact of different insertion locations and initial ranks in our method, with results presented in Fig.
-7.
-These findings are consistent with our analyses shown in Fig.
-4.
-Our results indicate that when training is restricted to the vision encoder, transfer accuracy drops significantly, suggesting that the pre-trained capabilities are negatively impacted.
-Moreover, at specific locations, increasing the maximum rank improves adaptation ability while not always reducing transfer accuracy.
-Interestingly, restricting updates to only the Attention modules can surpass our reported state-of-the-art Transfer metric in Table 1 and Table 2.
-However, this comes at the expense of reduced adaptation capabilities.
-To strike a better balance between adaptation and retention, we enable updates across all modules by default, ensuring an optimal trade-off between learning new tasks and retaining the model’s pre-trained capabilities.
-Analyses of pruning threshold and dense-iteration ratio.
-To better understand the trade-off between downstream learning and forgetting, we control the number of adaptive ranks by varying the maximum threshold $\kappa_{\max}$ and the fraction of dense training iterations (Fig.
+Effects of insertion locations and initial ranks.
+We study the impact of insertion locations and initial ranks (Fig.
+7), consistent with our earlier analyses (Fig.
+4).
+Training restricted to the vision encoder significantly reduces transfer accuracy, indicating degraded pre-trained capabilities.
+At certain locations, increasing rank improves adaptation but does not always hinder transfer accuracy.
+Interestingly, restricting updates to only Attention modules can even surpass our reported state-of-the-art Transfer results (Table 1, Table 3), but at the cost of weaker adaptation.
+To balance adaptation and retention, we update across all modules by default, achieving an optimal trade-off between learning new tasks and preserving pre-trained knowledge.
+Analyses of pruning threshold and dense‐iteration ratio.
+To understand the trade‐off between learning and forgetting, we vary the maximum threshold $\kappa_{\max}$ and the dense training ratio (Fig.
 8).
-Gradually raising $\kappa_{\max}$ improves Last accuracy by pruning more useless or noisy ranks, whereas overly aggressive pruning limits task adaptation; a modest setting of 0.005 strikes a good balance.
-The dense-iteration ratio governs pruning onset: pruning from the start (ratio = 0) removes ranks before it can learn from the training data, while delaying pruning too long increases the risk of retaining useless or noisy ranks that do not get pruned eventually.
+Raising $\kappa_{\max}$ improves Last accuracy by pruning noisy ranks, but overly aggressive pruning limits adaptation.
+For simplicity, we do not tune the value of $\kappa_{\max}$ for each downstream task.
+The dense‐iteration ratio controls pruning onset: pruning from the start (ratio = 0) removes ranks before they learn useful knowledge, while delaying too long risks retaining noisy ranks that do not get pruned.
 5 Conclusion
-In this paper, we introduced CoDyRA, a novel approach to addressing continual learning challenges in vision-language models.
-By dynamically adjusting ranks based on their importance, our method optimizes knowledge retention while enabling efficient adaptation to new tasks without adding inference overhead.
-Extensive evaluations across various continual learning settings demonstrate that our method achieves state-of-the-art performance, effectively balancing downstream adaptation and pre-trained retention.
-Limitations and future works.
-Our approach merges updates into pre-trained weights without modeling inter-task correlations, mainly evaluated using pre-trained CLIP model.
-Future works could explore leveraging inter-task relationships and extending our method to other PTMs and settings.
-Acknowledgments and Disclosure of Funding
-This work was partially supported by the ARC DECRA Fellowship (DE230101591) awarded to D.
-Gong.
-H.
-Lu is affiliated with CSIRO Data61 through a PhD scholarship and acknowledges the support of the Google PhD Fellowship.
+We introduced CoDyRA, a continual learning framework for CLIP that dynamically selects and minimizes LoRA ranks based on their learned importance.
+Guided by sparsity-promoting regularization, CoDyRA balances plasticity with stability, without storing past data, task information, or relying on assumptions.
+By keeping updates close to the model’s previous state while preserving essential ranks, our method mitigates forgetting and enhances representation quality.
+Extensive experiments across diverse continual learning benchmarks demonstrate that CoDyRA achieves state-of-the-art performance, effectively balancing downstream adaptation and knowledge retention.
+Future works.
+CoDyRA updates the model continually by balancing learning and forgetting in the current task.
+While enabling a flexible and efficient CL process with strong results, future work will further refine the modeling of rank relationships, inter-task dependencies, and forgetting measures.
+Reproducibility Statement
+Detailed experiment settings and implementation details are provided in Sec. 4.1 and Appendix B.
 Appendix A Additional Details of the Proposed Method
@@
 where $\mathcal{L}_{\text{sup}}$ is the supervised training loss, $\Delta_{t}$ is the all trainable parameters $\{\mathbf{w}^{m}_{t},\mathbf{B}^{m}_{t},\mathbf{A}^{m}_{t}\}_{m=1}^{M}$ at $t$-th iteration, $M$ is the total number of pre-trained weight matrices with LoRA of rank $r$, $\mathbf{w}^{m}_{t}$ represents the importance weights of the $m$-th LoRA module, and $\lambda>0$ controls the strength of the $\ell_{1}$ regularization.
-We apply proximal gradient descent [4] to handle the non-differentiable $\ell_{1}$ regularization of the importance weights [12].
+We apply proximal gradient descent (Beck & Teboulle, 2009) to handle the non-differentiable $\ell_{1}$ regularization of the importance weights (Ding et al., 2023).
 At the $t$-th iteration, the update rule for the importance weights of each LoRA becomes:
@@
 such that the input to the operator $\mathcal{T}_{\kappa}$ is the elements in $\hat{\mathbf{w}}^{m}_{t}$, $\hat{\mathbf{w}}^{m}_{t}$ is the value of $\mathbf{w}^{m}_{t}$ after the gradient update using only $\mathcal{L}_{\text{sup}}(\Delta_{t})$, and $\kappa=\eta_{t}\cdot\lambda$ is the threshold.
-The indicator function $\mathbbm{1}(\cdot)$ returns 1 if the condition is met, and 0 otherwise, while $\mathrm{sign}(\cdot)$ denotes the sign (+/)(+/-)( + / - ) of the input.
+The indicator function $\mathbbm{1}(\cdot)$ returns 1 if the condition is met, and 0 otherwise, while $\mathrm{sign}(\cdot)$ denotes the sign (+/)(+/-) of the input.
 This concludes our derivation of the dynamic update for the importance weights as presented in the main paper Eq.
@@
 Appendix B More Experimental Results and Ablations
-Reproducibility Statement.
-The source code will be made publicly available upon acceptance of the paper.
-We report experiment results averaged over 3 independent runs for reproducibility.
-Experiments were conducted on a single NVIDIA GeForce RTX 3090 GPU.
-B.1 Effects of Additional Adaptive Prediction Head on Visual Representations
-The related continual learning methods focus on updating the representation with the backbone model (e.g., [53, 66, 61], and CoDyRA) or newly added representation alignment or prediction head [37, 58].
+Further details of the experiment settings.
+The MTIL setting consists of 1,201 classes drawn from 11 diverse datasets: Aircraft (Maji et al., 2013), Caltech101 (Fei-Fei et al., 2004), CIFAR100 (Krizhevsky et al., 2009), DTD (Cimpoi et al., 2014), EuroSAT (Helber et al., 2019), Flowers (Nilsback & Zisserman, 2008), Food (Bossard et al., 2014), MNIST (Deng, 2012), OxfordPet (Parkhi et al., 2012), Cars (Krause et al., 2013), and SUN397 (Xiao et al., 2010).
+In the X-TAIL setting, a total of 10 datasets are used, with CIFAR100 (Fei-Fei et al., 2004) excluded to prevent domain overlap, following the protocol in (Xu et al., 2024).
+In line with (Xu et al., 2024), we use a 5-shot split for MTIL and a 16-shot split for X-TAIL.
+For initialization, we follow the standard convention: matrix $\mathbf{A}$ is randomly initialized, matrix $\mathbf{B}$ is set to zero, and the importance weights $\mathbf{w}$ are randomly initialized.
+B.1 More Experiment Results
+Table 5: Extended results for Table 1, comparing continual fine-tuning with different LoRA settings and the LoRA variant of LAE (Gao et al., 2023).
+Method
+Aircraft Caltech101 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
+CLIP
+   Zero-shot 23.5 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 58.5
+Transfer
+   Zero-shot (Radford et al., 2021) 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
+   LoRA (rr=4) 74.9 39.4 41.8 67.7 83.3 44.1 87.1 64.2 62.7 62.8
+   LoRA (rr=8) 77 38.2 38.4 67.8 83.3 44.6 87.6 64 62.8 62.6
+   LoRA (rr=16) 76.4 39.1 37.5 65.9 82.7 41.9 86.9 63.4 62.7 61.8
+   LAE-LoRA (Gao et al., 2023) 76.2 39.9 43.8 66.5 83.3 42.7 87.5 63 62.9 62.9
+   CoDyRA 74.3 36.8 44.2 69.9 83.5 42.8 88.9 64.6 63.4 63.2
+[…]
+Average
+   LoRA (rr=4) 23.5 78.4 40.6 47.3 71.4 83.7 50.1 87.5 65.1 64 61.2
+   LoRA (rr=8) 23.7 78.1 39.7 46.6 71.3 83.7 50.7 87.8 65.4 64.1 61.1
+   LoRA (rr=16) 23.6 78.5 40.5 43.0 70.4 82.8 49.5 87.5 64.9 64.0 60.5
+   LAE-LoRA (Gao et al., 2023) 29.2 81.3 50.7 68.8 77.4 83.8 53.2 88.5 66.3 64.2 66.4
+   CoDyRA 41.4 81.0 58.7 77.8 83.4 84.6 64.5 90.4 67.2 64.4 71.3
+[…]
+Last
+   LoRA (rr=4) 9.75 79.5 40.7 37.5 65.1 82.6 47.1 84.1 57.5 76.3 58
+   LoRA (rr=8) 12.7 78.7 39.6 42.6 62.4 82.3 46.8 83.6 60.2 76.1 58.5
+   LoRA (rr=16) 13.7 78.6 40.3 39.4 66.1 82.7 50.9 85.7 60.3 76.1 59.4
+   LAE-LoRA (Gao et al., 2023) 25.6 81.5 50.2 73.2 79.5 85 70.4 91.1 72.9 75.7 70.5
+   CoDyRA 37.7 81.5 65.1 89.9 91.4 85.5 96.8 93.3 77.3 73.5 79.2
+[…]
+B.1.1 Comparison to fixed‐rank LoRA baselines.
+Table 5 reports results for fixed‐rank LoRA (r = 4, 8, 16) and the LoRA variant of LAE.
+Three key observations emerge: (1) As LoRA rank increases, Last accuracy improves while Average accuracy declines, aligning with Fig.
+4: a relatively higher-rank LoRA facilitates learning new tasks but tends to increase forgetting, while lower-rank LoRA mitigates forgetting but limits adaptation.
+(2) Fixed‐rank LoRA shows a substantial Last‐accuracy gap compared to our adaptive method.
+This suggests that without dynamic rank selection, low‐relevance ranks (potentially encoding noise) are retained in the model, impairing generalization to unseen domains and accumulating inter‐task interference over time, resulting in severe performance drop.
+(3) Leveraging fast online LoRA updates, slow EMA aggregation in offline LoRA, and energy-based logit ensembling, LAE substantially outperforms fixed-rank LoRA (r = 16).
+Nonetheless, unlike our method, fixed-rank LoRA lacks fine-grained control over the contribution of individual ranks within each module.
+Table 6: Comparisons on full-shot MTIL setting.
+Method
+Aircraft Caltech101 CIFAR100 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
+CLIP
+   Zero-shot (Radford et al., 2021) 24.3 88.4 68.2 44.6 54.9 71 88.5 59.4 89 64.7 65.2 65.2
+Transfer
+   Zero-shot (Radford et al., 2021) 74.5 56.9 39.1 51.1 52.6 72.8 60.6 75.1 30.3 55.9 56.9
+   LwF (Li & Hoiem, 2017) 56.6 44.6 32.7 39.3 46.6 68.0 46.0 77.4 31.9 60.5 50.4
+   iCaRL (Rebuffi et al., 2017a) 77.1 61.0 40.5 45.3 54.4 74.6 47.9 76.7 36.3 58.6 57.2
+   LwF-VR (Ding et al., 2022) 73.5 55.6 35.6 41.5 47.0 68.3 53.9 69.3 26.8 51.9 52.3
+   WiSE-FT (Wortsman et al., 2022) 86.0 67.4 45.4 50.4 69.1 87.6 61.8 86.8 60.1 66.8 68.1
+   ZSCL (Zheng et al., 2023) 55.7 50.9 30.4 41.4 49.3 71.8 36.3 77.5 55.3 53.4 53.2
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Average
+   LwF (Li & Hoiem, 2017) 36.3 86.9 72.0 59.0 73.7 60.0 73.6 74.8 80.0 37.3 58.1 64.7
+   iCaRL (Rebuffi et al., 2017a) 35.5 89.2 72.2 60.6 68.8 70.0 78.2 62.3 81.8 41.2 62.5 65.7
+   LwF-VR (Ding et al., 2022) 29.6 87.7 74.4 59.5 72.4 63.6 77.0 66.7 81.2 43.7 60.7 65.1
+   WiSE-FT (Wortsman et al., 2022) 26.7 86.5 64.4 57.1 65.7 58.7 71.1 70.5 75.8 36.9 54.6 60.7
+   ZSCL (Zheng et al., 2023) 45.1 92.0 80.1 64.3 79.5 81.6 89.6 75.2 88.9 64.7 68.0 75.4
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Last
+   LwF (Li & Hoiem, 2017) 26.3 87.5 71.9 66.6 79.9 66.9 83.8 99.6 92.1 66.1 80.4 74.6
+   iCaRL (Rebuffi et al., 2017a) 35.8 93.0 77.0 70.2 83.3 88.5 90.4 86.7 93.2 81.2 81.9 80.1
+   LwF-VR (Ding et al., 2022) 20.5 89.8 72.3 67.6 85.5 73.8 85.7 99.6 93.1 73.3 80.9 76.6
+   WiSE-FT (Wortsman et al., 2022) 27.2 90.8 68.0 68.9 86.9 74.0 87.6 99.6 92.6 77.8 81.3 77.7
+   ZSCL (Zheng et al., 2023) 40.6 92.2 81.3 70.5 94.8 90.5 91.9 98.7 93.9 85.3 80.2 83.6
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+B.1.2 Additional results on full-shot MTIL.
+To further extend our evaluation, we report full-shot results on the MTIL benchmark (Table 6).
+For fairness, we adopt the same domain-prediction design used in recent methods.
+Consistent with Tables 1 and 3, CoDyRA achieves state-of-the-art performance, surpassing all baselines on both Transfer (unseen-domain generalization) and Last (continual downstream learning).
+Table 6 also compares against prompt-based approaches (L2P (Wang et al., 2022e), DualPrompt (Wang et al., 2022d), S-Prompt (Wang et al., 2022c)).
+Although these methods mitigate task forgetting via key–value matching for prompt selection, they neglect the preservation of pre-trained capabilities—evidenced by substantial drops in transfer accuracy—and further underperform recent domain-prediction approaches on last accuracy.
+Table 7: Statistical significance of CoDyRA results corresponding to Table 1.
+    Aircraft Caltech101 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
+CoDyRA Transfer $74.3^{\pm 0.52}$ $36.8^{\pm 0.23}$ $44.2^{\pm 0.56}$ $69.9^{\pm 0.56}$ $83.5^{\pm 0.23}$ $42.8^{\pm 0.18}$ $88.9^{\pm 0.42}$ $64.6^{\pm 0.47}$ $63.4^{\pm 0.56}$ $63.2^{\pm 0.28}$
+CoDyRA Average $41.4^{\pm 0.28}$ $81^{\pm 0.38}$ $58.7^{\pm 0.26}$ $77.8^{\pm 0.47}$ $83.4^{\pm 0.39}$ $84.6^{\pm 0.28}$ $64.5^{\pm 0.14}$ $90.4^{\pm 0.4}$ $67.2^{\pm 0.23}$ $64.4^{\pm 0.47}$ $71.3^{\pm 0.18}$
+CoDyRA Last $37.7^{\pm 0.42}$ $81.5^{\pm 0.24}$ $65.1^{\pm 0.63}$ $89.9^{\pm 0.55}$ $91.4^{\pm 0.38}$ $85.5^{\pm 0.16}$ $96.8^{\pm 0.08}$ $93.3^{\pm 0.3}$ $77.3^{\pm 0.66}$ $73.5^{\pm 0.21}$ $79.2^{\pm 0.18}$
+B.1.3 Training stability.
+Table 7 reports the mean and standard deviation of CoDyRA’s results in Table 1, averaged over three independent runs.
+The consistently low variance across metrics demonstrates that our method is stable to random initialization, highlighting CoDyRA’s robust adaptive rank selection.
+B.2 Effects of Additional Adaptive Prediction Head on Visual Representations
+The related continual learning methods focus on updating the representation with the backbone model (e.g., (Wang et al., 2022e; Zheng et al., 2023; Yu et al., 2024), and CoDyRA) or newly added representation alignment or prediction head (McDonnell et al., 2024; Xu et al., 2024).
 These two types of methods can be seen as orthogonal.
-We study how the representation updated by the proposed can work with the additional adaptive prediction head, relying on the aggression-based adapters in [58].
+We study how the representation updated by the proposed can work with the additional adaptive prediction head, relying on the aggression-based adapters in (Xu et al., 2024).
 We provide a more detailed analysis of the effects of the additional prediction head on the visual representations in X-TAIL setting in Table 8.
-Note that the Dual version of RAIL [58] stores all visual representations of all samples for all seen classes and domains in memory, it is not included in Tables 1 and 2 for a fair comparison.
+Note that the Dual version of RAIL (Xu et al., 2024) stores all visual representations of all samples for all seen classes and domains in memory, it is not included in Tables 1 and 3 for a fair comparison.
 RAIL is a regression-based method that applies an additional prediction head on top of the visual representations extracted by the frozen pre-trained vision encoder of the CLIP model.
@@
 CoDyRA + RAIL-Dual 63.1 74.2 83.6
-B.2 Analysis of Changes/Interference w.r.t. Pre-Trained Weights
+B.3 Analysis of Changes/Interference w.r.t. Pre-Trained Weights
 In Sec.
@@
 6 visualize and analyze the dynamic ranks assigned to each module.
-Here, we extend the analysis by comparing parameter changes to pre-trained weights using the amplification factor as in [21].
+Here, we extend the analysis by comparing parameter changes to pre-trained weights using the amplification factor as in (Hu et al., 2021).
 The amplification factor is given by:
@@
 $\mathbf{U}$ and $\mathbf{V}$ represent the top $r$ singular vectors of $\text{SVD}(\Delta\mathbf{W})$, where $r$ is the number of interested ranks in analysis.
-A higher amplification factor indicates that low-rank parameter updates amplify directions less emphasized in the pre-trained weights [21].
+A higher amplification factor indicates that low-rank parameter updates amplify directions less emphasized in the pre-trained weights (Hu et al., 2021).
 We visualize the amplification factors for both our method and vanilla LoRA trained on the Aircraft, EuroSAT and Oxford Pets datasets in Fig.
@@
 It shows that the proposed method with adaptive rank selection can achieve more focused/concentrated updating and less interference.
-B.3 More Visualizations and Analyses on Rank allocation results of different datasets
+B.4 More Visualizations and Analyses on Rank allocation results of different datasets
 In Sec.
@@
 Figure 17: Visualization and statistical analysis of rank activation on the SUN397 dataset using our proposed method.
-B.4 Evaluation with Other VLMs
-To evaluate the broader applicability of our method, we conducted preliminary experiments with BLIP [28] on X-TAIL (Table 6).
+B.5 Evaluation with Other VLMs
+To evaluate the broader applicability of our method, we conducted preliminary experiments with BLIP (Li et al., 2022) on X-TAIL (Table 9).
 Our method consistently outperformed prior works and exceeded the pre-trained model’s upper-bound performance, demonstrating its general effectiveness across different VLMs.
@@
 CoDyRA  ($r_{init}=16$) 97.20 / 3.95M 81.61 / 3.75M 94.44 / 3.85M
-B.5 Evaluation on Standard Parameter-Efficient Fine-Tuning (PEFT) Task
+B.6 Evaluation on Standard Parameter-Efficient Fine-Tuning (PEFT) Task
 Leveraging our dynamic rank-adaptive parameter updates, our method naturally extends to Parameter-Efficient Fine-Tuning (PEFT) tasks.
@@
 Crucially, by amplifying important ranks through learned higher importance weights, our approach significantly enhances downstream adaptation.
-Broader Impacts
-CoDyRA’s sparse, adaptive, importance-aware low-rank subspace learning significantly reduces compute and memory demands for continual learning, enabling on-device personalization in domains such as mobile health and education and empowering resource-constrained users and organizations.
-As a fundamental research effort, CodyRA does not address potential biases or hazards in training data, and its incremental updates could introduce privacy risks or stealthy backdoors.
-We therefore recommend transparent update logging, regular audits, and differential-privacy safeguards.
-Although more efficient than full fine-tuning, large-scale deployment still incurs energy costs; developers should monitor resource use, leverage low-power hardware, and explore carbon offsets.

### v7

@@
-Adaptive Rank, Reduced Forgetting: Knowledge Retention in Continual Learning Vision-Language Models with Dynamic Rank-Selective LoRA
+Adaptive Rank, Reduced Forgetting: Continual Learning with Dynamic Rank-Selective LoRA
 Abstract
 Continual learning (CL) aims to accumulate knowledge from sequential tasks without catastrophic forgetting.
-Vision–language models like CLIP, with strong generalization, are widely used for CL.
-Existing methods often adapt isolated PTM components, adding inference complexity and limiting PTM improvement, or rely on replay, stored information, or assumptions, incurring high costs and limited applicability.
-To advance models as continual learners, we explore CL via natural, efficient PTM updates instead of complex task-specific additions.
-We thus study continual low-rank learning and systematically analyze how LoRA ranks and placements affect learning and forgetting.
-We find that a relatively higher-rank LoRA improves task learning (i.e., plasticity) but increases forgetting, while a relatively lower-rank LoRA reduces forgetting (i.e., stability) but limits adaptation.
-Crucially, we find a plasticity–stability balance tied to rank across parameters and tasks, with moderately small ranks maximizing CL benefits.
-Motivated by this, we propose Continual Dynamic Rank-Selective LoRA (CoDyRA), which continually updates PTMs with LoRA adapters for all parameters of adaptively optimized rank number.
-While the new-task objective drives learning, CoDyRA adaptively minimizes ranks with sparsity-promoting regularization to reduce interference and forgetting, achieving a plasticity–stability balance tailored to different parameters and tasks.
-Despite allowing updates to all parameters, the adaptively selected and minimized LoRA ranks help keep the updated model closer to its previous state, while still enabling comprehensive and effective learning of new tasks.
-CoDyRA enables efficient CL as a sequence of LoRA-based tasks without storing past data, task information, or relying on assumptions.
-It preserves the original model architecture and deployment pipeline, adding no inference overhead.
-Extensive experiments show CoDyRA improves new representations while retaining old knowledge, achieving state-of-the-art results.
+Many CL methods for pre-trained models (PTMs) achieve strong performance but rely on sophisticated architectures or additional assumptions.
+They adapt isolated modules—adding inference complexity and limiting PTM improvement—or depend on replay, stored data, or assumptions, incurring extra cost and reduced applicability.
+To advance usage, we explore natural, efficient PTM CL instead of complex task-specific additions.
+We study continual low-rank learning and systematically analyze how LoRA ranks and placements affect learning and forgetting.
+We find that a relatively higher-rank LoRA improves task learning (i.e., plasticity) but increases forgetting, while lower-rank LoRA reduces forgetting (i.e., stability) but limits adaptation.
+Crucially, we observe a plasticity–stability balance tied to ranks across parameters and tasks, with moderately small ranks maximizing CL benefits.
+Motivated by this, we propose Continual Dynamic Rank-Selective LoRA (CoDyRA), continually updating PTMs with LoRA of adaptively optimized ranks.
+While the new-task objective drives learning, CoDyRA adaptively minimizes ranks with sparsity-promoting regularization to reduce interference and forgetting, achieving plasticity–stability balance adaptively for different parameters and tasks.
+It preserves the original model architecture and serving pipeline, adding no inference overhead.
+Extensive experiments on both CLIP and LLMs show CoDyRA improves new-task learning with strong knowledge retention, delivering state-of-the-art results.
 Code is available.
 1 Introduction
-Continual learning (CL) (Hadsell et al., 2020; De Lange et al., 2021; Wang et al., 2024b) focuses on incrementally learning from new data streams without catastrophic forgetting (Nguyen et al., 2019; McCloskey & Cohen, 1989) or the need to retrain on all previously seen data.
-With the rise of large-scale pre-trained models (PTMs), recent studies increasingly perform CL in this context (Zhou et al., 2024; McDonnell et al., 2024; Wang et al., 2024a; Jha et al., 2024).
-Many PTM-based CL approaches primarily focus on continual adaptation for downstream tasks (e.g., class-incremental learning, CIL (Smith et al., 2023; Wortsman et al., 2022)), with an emphasis on downstream tasks.
-While task-specific and downstream-oriented designs (Zhou et al., 2024; McDonnell et al., 2024; Xu et al., 2024; Yu et al., 2024) show promising results on particular CL tasks, PTMs remain largely isolated from the CL process, mitigating forgetting but undermining interaction with new tasks and thus the potential for continual improvement.
+Continual learning (CL) [23, 11, 78] focuses on incrementally learning from new data streams without catastrophic forgetting [56, 51] or the need to retrain on all previously seen data.
+With the rise of pre-trained models (PTMs), recent studies increasingly perform CL in this context [101, 52, 75, 31].
+Following the context of class-incremental learning (CIL) tasks [68, 85], many PTM-based CL approaches primarily focus on or are evaluated through continual adaptation to downstream tasks.
+While task-specific and downstream-oriented designs [101, 52, 90, 93] show promising results on particular CL tasks, PTMs remain largely isolated from the CL process, mitigating forgetting but undermining interaction with new tasks and thus the potential for continual improvement.
 To advance toward models that function as continual learners, we consider CL with PTMs that enable flexible and efficient updates, preserving or even enhancing generalizable representations while acquiring new knowledge.
-Leveraging the strong generalization and flexibility of jointly learned vision–text embeddings, CLIP as a vision–language model has been widely adopted in CL (Jha et al., 2024; Zheng et al., 2023; Wang et al., 2023; Zhou et al., 2023), including continual updating (Garg et al., 2023) and multi-domain CL (Zheng et al., 2023).
+Leveraging the strong generalization and flexibility of jointly learned vision–text embeddings, CLIP as a vision–language model has been widely adopted in CL [31, 99, 79, 103], including continual updating [21] and multi-domain CL [99].
 Beyond downstream performance, the retention of knowledge in PTMs is also evaluated.
-Some approaches, such as ZSCL (Zheng et al., 2023), aim to continually update PTM, i.e., CLIP, through full fine-tuning while mitigating forgetting via memory replay (Li & Hoiem, 2017; Wortsman et al., 2022; Rebuffi et al., 2017b) using additional reference data.
+Some approaches, such as ZSCL [99], aim to continually update the pretrained CLIP through full fine-tuning while mitigating forgetting via memory replay [38, 85, 64] using additional reference data.
 However, this leads to high training costs and still struggles with catastrophic forgetting (Fig.
 1(a)).
-To directly control the forgetting, some methods (Yu et al., 2024; Xu et al., 2024) introduce task-specific components isolated from the frozen PTMs, requiring task/domain prediction or gating at inference.
+To directly control the forgetting, some methods [93, 90] introduce task-specific components isolated from the frozen PTMs, requiring task/domain prediction or gating for routing the specific components at inference.
 These methods can deliver strong results with task-specific gating and branching; however, by isolating tasks and relying heavily on per-task modules and accurate task prediction, they become restricted to certain domains and limited in both general CL applicability and inference efficiency.
-Moreover, with restricted interaction between PTMs and CL process (e.g., isolation) limits PTMs’ generalization improvement potential (Fig.
+Moreover, the restricted interaction between PTMs and CL process (e.g., isolation) limits PTMs’ generalization improvement potential (Fig.
 1(b)).
-Figure 1: (a) Continual full fine-tuning leverages reference data to reduce forgetting of pre-trained knowledge (e.g., zero-shot performance (Zheng et al., 2023)) but remains costly and prone to forgetting.
-(b) Task-specific modular methods (e.g., (Yu et al., 2024; Xu et al., 2024)) add isolated components requiring domain prediction or gating, increasing complexity and limiting adaptability and generalization.
+Figure 1: (a) Continual full fine-tuning leverages reference data to reduce forgetting of pre-trained knowledge (e.g., zero-shot performance [99]) but remains costly and prone to forgetting.
+(b) Task-specific modular methods (e.g., [93, 90]) add isolated components requiring domain prediction or gating, increasing complexity and limiting adaptability and generalization.
 (c) Our approach enables efficient, universal continual PTM updating without reference data or task-specific designs, retaining PTM knowledge and improving performance on unseen data.
-Straightforward adaptation of PTMs can be achieved through direct model updating through full fine-tuning (FT) or parameter-efficient fine-tuning (PEFT) methods (Zhou et al., 2022c; b), such as Low-Rank Adaptation (LoRA) (Hu et al., 2021).
+General-purpose PTM updating can be achieved through direct model updating via straightforward full fine-tuning (FT) or parameter-efficient fine-tuning (PEFT) methods [105, 104], such as Low-Rank Adaptation (LoRA) [27].
 This raises a fundamental question: instead of relying on complex additions or task-specific components, can CL be performed by leveraging these natural PTM update mechanisms?
-Our work investigates how to mitigate catastrophic forgetting within these standard frameworks without imposing extensive assumptions.
+Our work investigates how to mitigate catastrophic forgetting within these standard frameworks without imposing additional assumptions.
 We propose an efficient CL approach for CLIP that enables the model to be updated naturally across a stream of tasks, while retaining foundational pre-trained knowledge.
 We build upon LoRA for efficiency.
-While there have been some LoRA-based CL methods (Liang & Li, 2024; Wu et al., 2025; Zhu et al., 2025), they rely on specific regularization (e.g., orthogonal property) for specific settings.
+There have been some LoRA-based CL methods [40, 87, 106], which rely on specific regularization (e.g., orthogonal property) for specific settings.
 We first systematically analyze and ablate how LoRA adapters at different locations in the network with different ranks influence learning and forgetting (Sec.
@@
 CoDyRA performs new task learning while jointly minimizing the rank of each LoRA adaptively.
-Rank selection is guided by sparsity-promoting regularization and learnable importance weights, which dynamically adjust the contribution of each rank during training.
-By updating with LoRA adapters of minimized ranks, CoDyRA inherently biases models toward remaining closer to their previous state, thereby alleviating forgetting in principle without task-specific assumptions.
+Rank selection is guided by sparsity-promoting regularization and learnable importance weights, which dynamically adjust the contribution of each rank and decide a balanced capacity in training.
+By updating with LoRA adapters of minimized ranks, CoDyRA inherently regularizes models toward remaining closer to their previous state, thereby alleviating forgetting in principle without task-specific assumptions.
 Our contributions are fourfold: (1) We study PTM-based CL with natural and efficient continual update of the model, not relying on complex task-specific additions, without storing past data, task information, or relying on assumptions.
@@
 It handles different parameters and tasks, yielding regularized updates that reduce interference and keep the model closer to its previous state with less forgetting.
-(4) Extensive experiments across multiple CL benchmarks validate our method’s superior performance in both downstream learning and mitigating forgetting, achieving a balanced trade-off between plasticity and stability.
+(4) Extensive experiments across multiple CL benchmarks demonstrate superior downstream performance and reduced forgetting, achieving a strong plasticity–stability balance.
+We further extend CoDyRA to LLMs beyond CLIP, validating its effectiveness on broader applications and pre-trained models.
 2 Related Work
@@
 Continual learning enables models to sequentially learn new tasks while retaining previously acquired knowledge.
-Experience replay (ER) methods (Luo et al., 2023; Aljundi et al., 2019b; Chaudhry et al., 2018a; Liu et al., 2020; Chaudhry et al., 2018b; Yan et al., 2022; 2021) store subsets of past data to refresh the model on prior tasks during new training.
-Parameter regularization approaches (Kirkpatrick et al., 2017; Aljundi et al., 2018; Zenke et al., 2017; Aljundi et al., 2019a; Jha et al., 2023) constrain updates to important weights to preserve past knowledge.
-Meanwhile, dynamic networks (Wang et al., 2022a; b; Zhou et al., 2022a; Wang et al., 2024a; McDonnell et al., 2024; Liang & Li, 2024; Wang et al., 2022e; d; Smith et al., 2023; Wang et al., 2022c) adjust their architecture on the fly, balancing the acquisition of new information with the retention of old knowledge.
+Experience replay (ER) methods [48, 3, 8, 42, 9, 91, 92] store subsets of past data to refresh the model on prior tasks during new training.
+Parameter regularization approaches [33, 1, 94, 2, 32, 39] constrain updates to important weights to preserve past knowledge.
+Dynamic networks [73, 74, 102, 75, 52, 40, 84, 83, 68, 82, 65] adjust their architecture on the fly, balancing the acquisition of new information with the retention of old knowledge.
 Continual Learning of CLIP.
-Continual learning of CLIP (Jha et al., 2024; Zhang et al., 2024b) aims to enable PTMs to sequentially learn across diverse domains while retaining their pre-trained generalization capabilities for previously seen tasks.
-(Garg et al., 2023) continually adapt CLIP to temporal distribution shifts via timestamped data updates.
-The approach introduced in ZSCL (Zheng et al., 2023) addresses the challenge of preserving zero-shot capabilities while adapting to new tasks, mitigating the risk of catastrophic forgetting.
-Subsequent works (Yu et al., 2024; Tang et al., 2025; Xu et al., 2024; Huang et al., 2025) have focused primarily on continual adaptation to downstream tasks while leveraging pre-trained predictions for unseen data.
-Recently, (Xu et al., 2024) proposed a more challenging setting, where the labels from different domains are mixed during testing.
+CL of CLIP [31, 97] aims to enable PTMs to sequentially learn across diverse domains while retaining their pre-trained generalization capabilities for previously seen tasks.
+[21] continually adapt CLIP to temporal distribution shifts via timestamped data updates.
+ZSCL [99] was proposed to address catastrophic forgetting while preserving zero-shot capabilities during task adaptation.
+Subsequent works [93, 69, 90, 29, 41] have focused primarily on continual adaptation to downstream tasks while leveraging pre-trained predictions for unseen data.
+Recently, a more challenging setting was proposed, where the labels from different domains are mixed during testing [90].
 Low-Rank Adaptation.
-Low-Rank Adaptation (LoRA) (Hu et al., 2021) has gained popularity as a parameter-efficient method for fine-tuning large pre-trained models.
-Similar to LoRA, (Wu et al., 2024b) learns a small set of task‐specific representation vectors at each layer.
-Building on LoRA’s framework, recent works enhance LoRA by reformulating parameter updates, focusing on initialization based on the pre-trained weights (Zhang et al., 2024a; Meng et al., 2024), treating each rank independently (Ding et al., 2023; Zhang et al., 2023; Liu et al., 2024), or employing a mixture of subspaces (Wu et al., 2024a).
-These advancements highlight LoRA’s effectiveness in efficiently fine-tuning large models while preserving generalization.
+Low-Rank Adaptation (LoRA) [27] has gained popularity as a parameter-efficient method for fine-tuning large pre-trained models.
+Similar to LoRA, [88] learns a small set of task‐specific representation vectors at each layer.
+Building on LoRA’s framework, recent works enhance LoRA by reformulating parameter updates, focusing on initialization based on the pre-trained weights [95, 53], treating each rank independently [14, 96, 43], or employing a mixture of subspaces [86].
+LoRA has been shown to be effective even with small ranks [5, 66], while our work investigates how to perform low-rank learning in CL and how to dynamically and adaptively decide capacity.
+Owing to their effectiveness and efficiency, LoRA-based methods have been extended to CL of LLMs, often relying on specialized regularization or architectural designs [80, 61, 60].
+For instance, O-LoRA [80] mitigates interference via orthogonal regularization; TreeLoRA [60] routes inputs to task-specific adapters using a gradient-similarity tree with bandit search.
 3 Methodology
@@
 The matrices $\mathbf{W}^{\text{Q}}_{l},\mathbf{W}^{\text{K}}_{l},\mathbf{W}^{\text{V}}_{l},\mathbf{W}^{\text{O}}_{l}\in\mathbb{R}^{d\times d}$ correspond to the Query, Key, Value, and Output projection matrices, respectively.
-The MLP in each transformer block consists of two linear layers separated by an activation function $\text{Act}(\cdot)$, such as ReLU or GELU (Hendrycks & Gimpel, 2016).
+The MLP in each transformer block consists of two linear layers separated by an activation function $\text{Act}(\cdot)$, such as ReLU or GELU [26].
 The MLP module at the $l$-th transformer layer is defined as: $\text{MLP}(x)=\text{Act}(x\mathbf{W}^{\text{FC}}_{l}+b^{\text{FC}}_{l})\mathbf{W}^{\text{Proj}}_{l}+b^{\text{Proj}}_{l},$ where $\mathbf{W}^{\text{FC}}_{l}\in\mathbb{R}^{d\times d_{m}}$ and $\mathbf{W}^{\text{Proj}}_{l}\in\mathbb{R}^{d_{m}\times d}$ are the weight matrices, with $d_{m}$ being the hidden dimension.
-Figure 2: Overview of CoDyRA: we propose dynamic rank-selection LoRA, enabling each pre-trained weight matrix to adaptively add necessary ranks for downstream adaptation while retaining pre-trained capabilities.
-After each task, dynamic rank updates are merged into the pre-trained weights with no inference overhead.
 CLIP model.
-In CLIP (Radford et al., 2021; Ilharco et al., 2021; Fang et al., 2024; Li et al., 2022), both the image encoder $f_{\theta}$ and text encoder $g_{\psi}$ are transformers that learn a joint vision–language embedding.
+In CLIP [62, 30, 17], both the image encoder $f_{\theta}$ and text encoder $g_{\psi}$ are transformers that learn a joint vision–language embedding.
 Given an image ${\mathbf{x}}\in\mathbb{R}^{H\times W\times C}$, the vision encoder produces features ${\mathbf{z}}_{\text{V}}=f_{\theta}({\mathbf{x}})\in\mathbb{R}^{D}$.
@@
 At inference, the probability of classifying $\mathbf{x}$} into class $y_{i}\in\{1,\dots,C\}$ is: $p(y_{i}|{\mathbf{x}})=\frac{\exp(\mathtt{sim}({\mathbf{z}}_{\text{V}},{\mathbf{z}}_{\text{T}}^{y_{i}})/\tau)}{\sum_{c=1}^{C}\exp(\mathtt{sim}({\mathbf{z}}_{\text{V}},{\mathbf{z}}_{\text{T}}^{y_{c}})/\tau)},$ where $\mathtt{sim}(\cdot)$ is cosine similarity and $\tau$ is a temperature parameter.
+CL with CLIP.
+We study continual learning of CLIP with multi-domain data, e.g., multi-domain task-incremental learning (MTIL) [99, 93] and cross-domain task-agnostic incremental learning (X-TAIL) [90].
+In MTIL, the model sequentially learns from $T$ tasks, where each task $t$ has dataset $\mathcal{D}^{t}=\{({\mathbf{x}}^{t}_{i},y^{t}_{i})\}_{i=1}^{N^{t}}$, with input image ${\mathbf{x}}^{t}_{i}\in\mathbb{R}^{H\times W\times C}$, label $y^{t}_{i}\in\mathcal{C}^{t}$, and $M^{t}$ classes in $\mathcal{C}^{t}=\{y^{t}_{j}\}_{j=1}^{M^{t}}$.
+In X-TAIL, the test-time category set includes both seen and unseen domains, $\mathcal{C}=\mathcal{C}_{\text{seen}}\cup\mathcal{C}_{\text{unseen}}$, where $\mathcal{C}_{\text{seen}}=\bigcup_{i=1}^{n}\mathcal{C}_{i}$ contains all classes from prior tasks and $\mathcal{C}_{\text{unseen}}$ denotes novel classes never encountered.
 Figure 2: Task adaptation and zero-shot capability retention after training CLIP with different LoRA insertion points and ranks.
@@
 “T” and “V” indicate text-encoder and vision-encoder modules, respectively.
-CL with CLIP.
-We study continual learning of CLIP with multi-domain data, e.g., multi-domain task-incremental learning (MTIL) (Zheng et al., 2023; Yu et al., 2024) and cross-domain task-agnostic incremental learning (X-TAIL) (Xu et al., 2024).
-In MTIL, the model sequentially learns from $T$ tasks, where each task $t$ has dataset $\mathcal{D}^{t}=\{({\mathbf{x}}^{t}_{i},y^{t}_{i})\}_{i=1}^{N^{t}}$, with input image ${\mathbf{x}}^{t}_{i}\in\mathbb{R}^{H\times W\times C}$, label $y^{t}_{i}\in\mathcal{C}^{t}$, and $M^{t}$ classes in $\mathcal{C}^{t}=\{y^{t}_{j}\}_{j=1}^{M^{t}}$.
-In X-TAIL, the test-time category set includes both seen and unseen domains, $\mathcal{C}=\mathcal{C}_{\text{seen}}\cup\mathcal{C}_{\text{unseen}}$, where $\mathcal{C}_{\text{seen}}=\bigcup_{i=1}^{n}\mathcal{C}_{i}$ contains all classes from prior tasks and $\mathcal{C}_{\text{unseen}}$ denotes novel classes never encountered.
 3.2 Analyses on LoRA Location and Rank
 To enable efficient and natural continual updating of PTMs, we employ LoRA and analyze its learning and forgetting capability for optimizing its usage.
-We examine how LoRA’s rank and placement affect downstream learning and forgetting through two types of analyses (Pearl, 2022; Meng et al., 2022), providing a comprehensive view of their overall impact.
+We examine how LoRA’s rank and placement affect downstream learning and forgetting through two types of analyses [59, 54], providing a comprehensive view of their overall impact.
 (1) We first apply LoRA components with varying ranks at different locations in the pre-trained CLIP, for measurement of importance or contribution (Fig.
-4); (2) We apply LoRA across all PTM weights and systematically prune them for evaluation, for testing the standalone capability (Fig.
-4).
+3); (2) We apply LoRA across all PTM weights and systematically prune them for evaluation, for testing the standalone capability (Fig.
+3).
 We train the model on downstream datasets and assess its performance on both (1) the new task, i.e., learning new knowledge, and (2) an additional reference dataset to evaluate the retention and forgetting of existing capabilities.
-We achieve three main results to guide methodology design.
 Without loss of generality, the learning–forgetting analyses can be extended to each learning step in straightforward, assumption-free CL.
 Applying LoRA on all weights.
-Unlike prior works (Hu et al., 2021; Zhang et al., 2024a; Meng et al., 2024) that only apply LoRA to limited components by default, we first examine the impact of different LoRA placements.
+Unlike prior works [27, 95, 53] that only apply LoRA to limited components by default, we first examine the impact of different LoRA placements.
 As shown in Fig.
 3 and Fig.
-4, applying LoRA to only specific modules alone may limit learning capacity in different ways: (1) LoRA placement strongly impacts downstream performance and retention of pre-trained capabilities; (2) Updating only the vision encoder causes more forgetting; (3) Updating only attention modules mitigates forgetting but limits learning on new tasks.
+3, applying LoRA to only specific modules alone may limit learning capacity in different ways: (1) LoRA placement strongly impacts downstream performance and retention of pre-trained capabilities; (2) Updating only the vision encoder causes more forgetting; (3) Updating only attention modules mitigates forgetting but limits learning on new tasks.
 These observations reflect the distinct roles of PTM components.
-Takeaway 1: Instead of manual or fixed LoRA placement at specific parameters of PTM like previous methods (Hu et al., 2021; Liang & Li, 2024; Yu et al., 2024; Zhang et al., 2024a; Meng et al., 2024), we apply LoRA to all weights and optimize for an adaptive configuration.
+Takeaway 1: Instead of manual or fixed LoRA placement [27, 95, 53], we apply LoRA to all weights and optimize for an adaptive configuration.
 The effects of LoRA ranks on balancing learning and forgetting.
-We identified that the rank number of LoRA plays a critical role in balancing learning and forgetting.
+We identified the rank number of LoRA plays a critical role for balancing learning and forgetting.
 As shown in Fig.
-4, we can observe that a relatively higher-rank LoRA facilitates learning new tasks but tends to increase forgetting, while a lower-rank LoRA mitigates forgetting but limits adaptation.
+3, we can observe that a relatively higher-rank LoRA facilitates learning new tasks but tends to increase forgetting, while a lower-rank LoRA mitigates forgetting but limits adaptation.
 This suggests that lower-rank LoRA interferes less with existing knowledge (i.e., reduces forgetting) while learning less new knowledge.
 Fortunately, we identify a balance (top-right in Fig.
-4), where a moderately small rank reduces forgetting while enabling enough adaptability.
-During training, low-rank parameter updates tend to enable the updated model to be closer to the original model, with less interference and forgetting.
+3), where a moderately small rank reduces forgetting while enabling enough adaptability.
+During training, low-rank updates tend to enable the updated model to be closer to the original model, with less interference and forgetting.
 In the extreme case, a rank-zero LoRA learns nothing and introduces no forgetting.
@@
 On the other hand, the task-learning objective emphasizes new knowledge acquisition, enabling a balance.
-Takeaway 2: A plasticity–stability balance exists and associates with LoRA rank (high vs.
+Takeaway 2: A plasticitystability balance exists and associates with LoRA rank (high vs.
 low), which can be adaptively achieved by jointly optimizing the task objective and minimizing LoRA ranks.
@@
 3 shows that the optimal rank for balance depends on both parameter placement and the learning task, consistent with the pruning-based sufficiency analysis in Fig.
-4.
+3.
 For example, removing updates to the Value projection in the Attention module of the vision encoder had little impact on adaptation but substantially restored pre-trained performance.
 Interestingly, pruning the MLP layer slightly reduced task performance while enhancing pre-trained capabilities.
-Takeaway 3: The learning–forgetting balance point tied to rank varies across modules and tasks, necessitating adaptive optimization.
+Takeaway 3: The learningforgetting balance point tied to rank varies across modules and tasks, necessitating adaptive optimization.
+Figure 4: Overview of CoDyRA: we propose dynamic rank-selection LoRA, enabling each pre-trained weight matrix to adaptively add necessary ranks for downstream adaptation while retaining pre-trained capabilities.
+After each task, dynamic rank updates are merged into the pre-trained weights with no inference overhead.
 3.3 Proposed CoDyRA with Sparsity-Induced Rank Selection and Minimization
@@
 We introduce adaptive rank-selective updates for all pre-trained weight matrices in the Attention and MLP modules of both vision and text encoders (Fig.
-2).
+4).
 The sparsity-promoting regularization enables adaptive rank selection and minimization for each LoRA adapter on each task.
-For each task in CL, beyond adaptive rank minimization, CoDyRA requires no additional priors or assumptions and does not store or maintain information (e.g., data, parameters, or statistics) from previous steps, unlike prior methods (Zheng et al., 2023; Yu et al., 2024; Xu et al., 2024).
+For each task in CL, beyond adaptive rank minimization, CoDyRA requires no additional priors or assumptions and does not store or maintain information (e.g., data, parameters, or statistics) from previous steps, unlike prior methods [99, 93, 90].
 During training on task $t$, we denote the pre-trained weights, or those updated from previous tasks, as $\{\mathbf{W}^{t,m}_{0}\}_{m=1}^{M}$, where $M$ is the total number of weight matrices updated by LoRA.
@@
 Regularized low-rank updates with rank importance.
-To update weight matrix $\mathbf{W}_{0}^{t,m}\in\mathbb{R}^{d\times k}$, LoRA (Hu et al., 2021) introduces two low-rank matrices, $\mathbf{B}^{t,m}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}^{t,m}\in\mathbb{R}^{r\times k}$, such that $\Delta\mathbf{W}^{t,m}=\mathbf{B}^{t,m}\mathbf{A}^{t,m}$, where $r\ll\min(d,k)$.
-The updated weight matrix is then defined as:
+To update weight matrix $\mathbf{W}_{0}^{t,m}\in\mathbb{R}^{d\times k}$, LoRA [27] introduces two low-rank matrices, $\mathbf{B}^{t,m}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}^{t,m}\in\mathbb{R}^{r\times k}$, such that $\Delta\mathbf{W}^{t,m}=\mathbf{B}^{t,m}\mathbf{A}^{t,m}$, where $r\ll\min(d,k)$.
+The updated weight matrix is defined as:
 $\mathbf{W}^{t,m}=\mathbf{W}_{0}^{t,m}+\Delta\mathbf{W}^{t,m}=\mathbf{W}_{0}^{t,m}+\mathbf{B}^{t,m}\mathbf{A}^{t,m}.$ (1)
@@
 (2) enables the model to dynamically optimize the importance of each rank and minimize active ranks through gradient descent.
-[^1] To align the training process with the pre-trained model, we use CLIP’s contrastive loss as the optimization objective during pre-training.
-To only retain the essential ranks for learning new tasks (thus reducing forgetting), we can achieve optimizing $R(\Delta\mathbf{W}^{t,m})$ through an $\ell_{1}$ norm-based sparsity-promoting regularization on $\mathbf{w}^{t,m}$, i.e., $\|\mathbf{w}^{t,m}\|_{1}$.
+To only retain the essential ranks for learning new tasks (thus reducing forgetting), we can achieve optimizing $R(\Delta\mathbf{W}^{t,m})$ through an $\ell_{1}$ norm-based sparsity-promoting regularization [71, 14] on $\mathbf{w}^{t,m}$, i.e., $\|\mathbf{w}^{t,m}\|_{1}$.
 The optimization objective for parameters $\{\mathbf{w}^{t,m},\mathbf{B}^{t,m},\mathbf{A}^{t,m}\}_{m=1}^{M}$ learned on task $t$ is:
-$\mathcal{L}_{\text{train}}^{t}:={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{L}_{\text{sup}}^{t}}}+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\lambda\sum\nolimits_{m=1}^{M}\|\mathbf{w}^{t,m}\|_{1}}},$ (3)
+$\mathcal{L}_{\text{train}}^{t}:={{\color[rgb]{0.55,0.14,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.55,0.14,1}\pgfsys@color@cmyk@stroke{0.45}{0.86}{0}{0}\pgfsys@color@cmyk@fill{0.45}{0.86}{0}{0}\mathcal{L}_{\text{sup}}^{t}}}+{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\lambda\sum\nolimits_{m=1}^{M}\|\mathbf{w}^{t,m}\|_{1}}},$ (3)
 where $\mathcal{L}_{\text{sup}}^{t}$ is the supervised training loss for current task, $M$ is the number of weight matrices inserted $r$-rank update, and $\mathbf{w}^{t,m}$ represents the importance weights of each rank added to the weight matrix $\mathbf{W}^{t,m}_{0}$.
 The $\ell_{1}$ regularization strength is controlled by $\lambda$.
-To handle the non-differentiable $\ell_{1}$ regularization applied to the importance weights, we adopt the proximal gradient method (Beck & Teboulle, 2009).
+Minimizing $\|\mathbf{w}^{t,m}\|_{1}$ encourages minimizing of $\|\Delta\mathbf{W}^{t,m}\|_{F}$ if given $\mathbf{A}^{t,m}$ and $\mathbf{B}^{t,m}$ (according to Eq.
+(2)).
+It softly reduces capacity in updates conflicting with critical information encoded in the learned parameters, without introducing additional assumptions.
+More discussions are in Appendix A.3.
+To handle the non-differentiable $\ell_{1}$ regularization applied to the importance weights, we adopt the proximal gradient method [4].
 The $i$-th element of $\hat{\mathbf{w}}^{t,m}$ for rank $i$ is updated via soft-thresholding:
-$\mathbf{w}_{i}^{t,m}:=\mathbbm{1}(\lvert\hat{\mathbf{w}}^{t,m}_{i}\rvert>\kappa)\cdot(\hat{\mathbf{w}}^{t,m}_{i}+\mathrm{sign}(\hat{\mathbf{w}}^{t,m}_{i})\cdot\kappa),$ (4)
+$\mathbf{w}_{i}^{t,m}:=\mathbbm{1}(\lvert\hat{\mathbf{w}}^{t,m}_{i}\rvert>\kappa)\cdot(\hat{\mathbf{w}}^{t,m}_{i}{-}\mathrm{sign}(\hat{\mathbf{w}}^{t,m}_{i})\cdot\kappa),$ (4)
 where $\hat{\mathbf{w}}^{t,m}_{i}$ denotes the value of $\mathbf{w}^{t,m}_{i}$} after applying the gradient update from the supervised loss $\mathcal{L}_{\text{sup}}^{t}$.
@@
 More details are left in Appendix A.
-This approach preserves only the ranks with significant importance for learning the current task and prunes those with low relevance, adapting across weights and tasks.
+This approach preserves only the ranks with significant importance for learning current task and prunes those with low relevance, adapting across weights and tasks.
 The importance weights on preserved ranks softly highlight their relative significance.
@@
 (4), allowing all ranks to capture task-relevant information before sparsity is enforced.
-Continually learn on new tasks.
-CoDyRA views CL as a sequence of tasks with LoRA, without storing past data Zheng et al.
-(2023), task information Yu et al.
-(2024); Xu et al.
-(2024), or relying on assumptions (Liang & Li, 2024).
-For each new task, new CoDyRA modules are initialized and optimized; low-importance ranks are pruned, and non-zero ranks merged into the weights.
-Sparsity-promoting regularization controls task capacity, keeping the model close to its previous state and reducing forgetting.
-This enables efficient continual updating with only sparsity-based regularization.
-Table 1: Comparisons on X-TAIL for each domain.
+Table 1: Comparisons on 5-shot MTIL setting.
 The best and the second best results are highlighted in red and blue, respectively.
-Methods marked with $\dagger$ indicate the use of domain prediction or distribution detection (Yu et al., 2024; Xu et al., 2024).
-Method
-Aircraft Caltech101 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
+Methods marked with $\dagger$ indicate the use of domain prediction or distribution detection [93, 90].
+    Method Aircraft Caltech101 CIFAR100 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
 CLIP
-   Zero-shot 23.5 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 58.5
+   Zero-shot [62] 24.3 88.4 68.2 44.6 54.9 71.0 88.5 59.4 89.0 64.7 65.2 65.3
 Transfer
-   Zero-shot (Radford et al., 2021) 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
-   LwF (Li & Hoiem, 2017) 66.6 26.9 19.5 51.0 78.4 26.6 68.9 35.5 56.1 47.7
-   WiSE-FT (Wortsman et al., 2022) 70.1 31.9 25.3 56.3 79.8 29.9 74.9 45.6 56.8 52.3
-   iCaRL (Rebuffi et al., 2017b) 71.7 35.0 43.0 63.4 86.9 43.9 87.8 63.7 60.0 61.7
-   ZSCL (Zheng et al., 2023) 73.3 32.6 36.8 62.1 83.8 42.1 83.6 56.5 60.2 59.0
-[…]
-[…]
-   CoDyRA 74.3 36.8 44.2 69.9 83.5 42.8 88.9 64.6 63.4 63.2
-[…]
-Average
-   LwF (Li & Hoiem, 2017) 24.7 79.7 38.3 36.9 63.9 81.0 36.5 71.9 42.7 56.7 53.2
-   WiSE-FT (Wortsman et al., 2022) 27.1 76.5 40.9 31.3 68.7 81.6 31.4 74.7 51.7 58.4 54.2
-   iCaRL (Rebuffi et al., 2017b) 25.4 72.1 37.5 51.6 65.1 87.1 59.1 88.0 63.7 60.1 61.0
-   ZSCL (Zheng et al., 2023) 36.0 75.0 40.7 40.5 71.0 85.3 46.3 83.3 60.7 61.5 60.0
-[…]
-[…]
-   CoDyRA 41.4 81.0 58.7 77.8 83.4 84.6 64.5 90.4 67.2 64.4 71.3
-[…]
-Last
-   LwF (Li & Hoiem, 2017) 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
-   WiSE-FT (Wortsman et al., 2022) 21.8 76.8 42.9 20.8 77.5 84.9 30.7 76.6 75.8 72.5 58.0
-   iCaRL (Rebuffi et al., 2017b) 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
-   ZSCL (Zheng et al., 2023) 33.1 75.3 43.5 35.2 74.6 87.4 50.4 84.2 77.3 73.4 63.4
-[…]
-[…]
-   CoDyRA 37.7 81.5 65.1 89.9 91.4 85.5 96.8 93.3 77.3 73.5 79.2
-[…]
-4 Experiments
-4.1 Experimental Setting
-Datasets.
-We evaluate our method on two widely used continual learning settings for vision-language models: MTIL (Zheng et al., 2023; Yu et al., 2024) and X-TAIL (Xu et al., 2024).
-A wide range of fine-grained datasets is used, where each dataset is treated as an individual task to be incrementally learned.
-We leave detailed experimental settings in Appendix B.
-Evaluation Metrics.
-We use the same metrics as prior work (Zheng et al., 2023; Yu et al., 2024; Xu et al., 2024): “Transfer” for zero-shot performance on unseen data, “Last” for retention of earlier tasks, and “Average” for mean accuracy across all datasets of all learning tasks.
-Implementation Details.
-We follow the setups in (Zheng et al., 2023; Yu et al., 2024; Xu et al., 2024) using CLIP with a ViT-B/16 backbone (Radford et al., 2021).
-By default, CoDyRA is applied to all pre-trained weight matrices in the vision and text encoders with an initial rank of 16.
-Each task is trained for 500 iterations with AdamW (Loshchilov & Hutter, 2017), using dense training ratios of 0.5 for X-TAIL and 0.7 for MTIL.
-$\kappa_{\text{max}}$ is set as 0.005.
-4.2 Experimental Results
-Cross-domain task-agnostic incremental learning.
-We report X-TAIL results in Table 1.
-Unlike prior methods (Yu et al., 2024; Xu et al., 2024) that rely on zero-shot predictions to preserve pre-trained capabilities, our approach continually updates the model with minimal forgetting.
-As those methods are bounded by zero-shot performance, ours surpasses this limit, achieving higher transfer accuracy.
-While continual updates may slightly affect some unseen predictions, the overall effect is positive: the model adapts to new tasks while modestly improving generalization.
-[…]
-Table 2: Average transfer accuracy on unseen datasets not encountered during CL.
-Method Average Transfer Accuracy on Unseen Data
-CIFAR100 Places365 ImageNet-1k Average
-Zero-Shot (Radford et al., 2021) 68.24 33.77 66.72 56.24
-MoE-Adapter (Yu et al., 2024) 68.24 33.77 66.72 56.24
-RAIL (Xu et al., 2024) 68.24 33.77 66.72 56.24
-CoDyRA 68.95 36.52 68.52 58.00
-Table 3: Comparisons on 5-shot MTIL setting.
-Method
-Aircraft Caltech101 CIFAR100 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
-CLIP
-   Zero-shot (Radford et al., 2021) 24.3 88.4 68.2 44.6 54.9 71.0 88.5 59.4 89.0 64.7 65.2 65.3
-Transfer
-   Zero-shot (Radford et al., 2021) 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
-   LwF (Li & Hoiem, 2017) 72.1 49.2 35.9 44.5 41.1 66.6 50.5 69.0 19.0 51.7 50.0
-   LwF-VR (Ding et al., 2022) 82.2 62.5 40.1 40.1 56.3 80.0 60.9 77.6 40.5 60.8 60.1
-   WiSE-FT (Wortsman et al., 2022) 77.6 60.0 41.3 39.4 53.0 76.6 58.1 75.5 37.3 58.2 57.7
-   ZSCL (Zheng et al., 2023) 84.0 68.1 44.8 46.8 63.6 84.9 61.4 81.4 55.5 62.2 65.3
+   Zero-shot [62] 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
+   LwF [38] 72.1 49.2 35.9 44.5 41.1 66.6 50.5 69.0 19.0 51.7 50.0
+   LwF-VR [15] 82.2 62.5 40.1 40.1 56.3 80.0 60.9 77.6 40.5 60.8 60.1
+   WiSE-FT [85] 77.6 60.0 41.3 39.4 53.0 76.6 58.1 75.5 37.3 58.2 57.7
+   ZSCL [99] 84.0 68.1 44.8 46.8 63.6 84.9 61.4 81.4 55.5 62.2 65.3
 […]
@@
 Average
-   LwF (Li & Hoiem, 2017) 23.5 77.4 43.5 41.7 43.5 52.2 54.6 63.4 68.0 21.3 52.6 49.2
-   LwF-VR (Ding et al., 2022) 24.9 89.1 64.2 53.4 54.3 70.8 79.2 66.5 79.2 44.1 61.6 62.5
-   WiSE-FT (Wortsman et al., 2022) 32.0 87.7 61.0 55.8 68.1 69.3 76.8 71.5 77.6 42.0 59.3 63.7
-   ZSCL (Zheng et al., 2023) 28.2 88.6 66.5 53.5 56.3 73.4 83.1 56.4 82.4 57.5 62.9 64.4
+   LwF [38] 23.5 77.4 43.5 41.7 43.5 52.2 54.6 63.4 68.0 21.3 52.6 49.2
+   LwF-VR [15] 24.9 89.1 64.2 53.4 54.3 70.8 79.2 66.5 79.2 44.1 61.6 62.5
+   WiSE-FT [85] 32.0 87.7 61.0 55.8 68.1 69.3 76.8 71.5 77.6 42.0 59.3 63.7
+   ZSCL [99] 28.2 88.6 66.5 53.5 56.3 73.4 83.1 56.4 82.4 57.5 62.9 64.4
 […]
@@
 Last
-   LwF (Li & Hoiem, 2017) 22.1 58.2 17.9 32.1 28.1 66.7 46.0 84.3 64.1 31.5 60.1 46.5
-   LwF-VR (Ding et al., 2022) 22.9 89.8 59.3 57.1 57.6 79.2 78.3 77.7 83.6 60.1 69.8 66.9
-   WiSE-FT (Wortsman et al., 2022) 30.8 88.9 59.6 60.3 80.9 81.7 77.1 94.9 83.2 62.8 70.0 71.9
-   ZSCL (Zheng et al., 2023) 26.8 88.5 63.7 55.7 60.2 82.1 82.6 58.6 85.9 66.7 70.4 67.4
+   LwF [38] 22.1 58.2 17.9 32.1 28.1 66.7 46.0 84.3 64.1 31.5 60.1 46.5
+   LwF-VR [15] 22.9 89.8 59.3 57.1 57.6 79.2 78.3 77.7 83.6 60.1 69.8 66.9
+   WiSE-FT [85] 30.8 88.9 59.6 60.3 80.9 81.7 77.1 94.9 83.2 62.8 70.0 71.9
+   ZSCL [99] 26.8 88.5 63.7 55.7 60.2 82.1 82.6 58.6 85.9 66.7 70.4 67.4
 […]
@@
 […]
-Multi-domain task-incremental learning.
-We further evaluate our method in the few-shot MTIL setting (Table 3), following (Yu et al., 2024; Xu et al., 2024).
-As with X-TAIL, our approach outperforms prior methods across all metrics without extra techniques and even surpasses their zero-shot upper bound in average transfer accuracy.
-Incorporating techniques from (Yu et al., 2024; Xu et al., 2024) further boosts performance.
+Continual learning framework.
+For each task, we initialize and optimize CoDyRA modules, low-importance ranks are pruned via sparsity-promoting regularization, and remaining active ranks are merged into the backbone weights after training.
+This regularization naturally constrains update capacity, preserving prior knowledge while enabling efficient adaptation.
+Following standard conventions, we initialize LoRA matrix $\mathbf{A}^{t,m}$ randomly and $\mathbf{B}^{t,m}$ to zero, and our proposed importance $\mathbf{w}^{t,m}$ is randomly initialized.
+4 Experiments
+We evaluate our approach on continual learning tasks for both vision-language models and LLMs.
+For vision-language CLIP [62, 30] model, we utilize the MTIL [99] and X-TAIL [90] benchmarks.
+For LLMs, we assess performance on the TRACE benchmark [81] using LLaMA [72, 16] and Gemma [70].
+Detailed experiment settings and additional results are provided in Appendix B.
+4.1 Experimental Results
+[…]
+Experiments on language model continual learning.
+Table 2 evaluates CoDyRA on the TRACE benchmark [81] using LLaMA [16, 72] and Gemma [70].
+Traditional regularization (EWC [33], OGD [18]), rehearsal (GEM [44]), and prompt-based methods (L2P [84], DualPrompt [83]) often underperform simple In-Context Learning (ICL [7]).
+Among parameter-efficient approaches, naive SeqLoRA suffers from significant forgetting.
+While O-LoRA [80], HiDeLoRA [76, 77], and TreeLoRA [60] improve stability, they impose rigid orthogonality, complex decomposition, or structural expansion.
+In contrast, CoDyRA achieves state-of-the-art performance without such constraints.
+By leveraging adaptive rank selection via sparsity-promoting regularization, it dynamically allocates capacity (minimizing ranks for stability while expanding for plasticity) to optimize the trade-off between learning and forgetting.
+Table 2: Comparison with a broad range of CL methods on the TRACE benchmark.
+We report Overall Performance (OP (%) $\uparrow$) and Backward Transfer (BWT (%) $\downarrow$).
+Results are averaged over three runs with standard deviations.
+The best results are highlighted in bold.
+FIX(ICL) SeqLoRA OGD GEM EWC L2P DualPrompt HiDeLoRA O-LoRA TreeLoRA CoDyRA
+meta-llama / LLaMA-2-7B-Chat
+OP 38.94 $\pm$ 0.3 34.3 $\pm$ 1.2 42.09 $\pm$ 1.6 40.08 $\pm$ 1.6 42.36 $\pm$ 1.2 36.23 $\pm$ 0.8 37.69 $\pm$ 1.2 41.60 $\pm$ 0.8 42.78 $\pm$ 0.8 43.52 $\pm$ 1.0 43.82 $\pm$ 0.9
+BWT 18.5 $\pm$ 0.8 8.06 $\pm$ 1.2 6.77 $\pm$ 1.2 5.97 $\pm$ 0.8 8.25 $\pm$ 0.8 8.03 $\pm$ 0.8 7.12 $\pm$ 0.4 7.16 $\pm$ 0.4 3.46 $\pm$ 0.4 3.25 $\pm$ 0.5
+google / Gemma-2B-it
+OP 32.3 $\pm$ 0.2 31.89 $\pm$ 0.8 32.85 $\pm$ 1.4 26.48 $\pm$ 1.5 28.35 $\pm$ 1.6 31.14 $\pm$ 1.2 32.42 $\pm$ 1.0 33.25 $\pm$ 0.9 33.73 $\pm$ 0.8 33.41 $\pm$ 0.9 33.96 $\pm$ 1.0
+BWT 15.28 $\pm$ 0.4 12.27 $\pm$ 0.9 18.25 $\pm$ 0.9 16.96 $\pm$ 1.2 15.77 $\pm$ 0.7 14.25 $\pm$ 0.5 13.66 $\pm$ 0.5 12.36 $\pm$ 0.4 8.50 $\pm$ 0.5 7.94 $\pm$ 0.8
+meta-llama / LLaMA-3-1B-Instruct
+OP 31.16 $\pm$ 0.4 29.73 $\pm$ 1.6 30.12 $\pm$ 2.0 32.19 $\pm$ 2.0 31.96 $\pm$ 1.6 29.38 $\pm$ 1.2 30.76 $\pm$ 1.2 33.73 $\pm$ 1.2 32.94 $\pm$ 0.8 36.14 $\pm$ 0.7 37.46 $\pm$ 0.8
+BWT 17.03 $\pm$ 1.2 15.2 $\pm$ 1.6 10.74 $\pm$ 1.6 11.62 $\pm$ 1.2 13.57 $\pm$ 0.8 11.34 $\pm$ 0.8 12.36 $\pm$ 0.8 12.89 $\pm$ 1.2 7.36 $\pm$ 0.8 5.11 $\pm$ 0.8
+Table 3: Pre-trained retention on unseen datasets.
+Method Zero-shot Retention on Unseen Data
+CIFAR100 Places365 ImageNet-1k Average
+Zero-Shot [62] 68.24 33.77 66.72 56.24
+MoE-Adapter [93] 68.24 33.77 66.72 56.24
+RAIL [90] 68.24 33.77 66.72 56.24
+CoDyRA 68.95 36.52 68.52 58.00
 4.2 Discussions
-Pre-trained knowledge retention and potential pre-trained model improvement.
-To assess transfer beyond the zero-shot limits of prior works, we evaluated on unseen datasets—CIFAR100 (Krizhevsky et al., 2009), Places365 (Zhou et al., 2017), and ImageNet-1k (Deng et al., 2009) (Table 2).
-Unlike methods that rely solely on zero-shot predictions (Yu et al., 2024; Xu et al., 2024) and are constrained by CLIP’s baseline, our approach integrates new knowledge into the pre-trained space while preserving existing capabilities.
-This yields superior zero-shot transfer accuracy, demonstrating that CoDyRA effectively expands general knowledge and improves overall representations.
-Rank allocation across different datasets.
-We visualize the ranks merged into the pre-trained weights for different datasets in Fig.
+Pre-trained knowledge retention and potential improvement.
+We assess transfer capability on unseen datasets: CIFAR100 [35], Places365 [100], and ImageNet-1k [12] (Table 3).
+Unlike prior methods [93, 90] that freeze the backbone and remain bounded by CLIP’s initial zero-shot performance, CoDyRA continually integrates new knowledge into the parameter space.
+This allows our model to surpass the original zero-shot baseline, demonstrating that CoDyRA effectively expands general knowledge and potentially improves overall representations.
+Rank allocation analysis.
+We visualize rank allocation across datasets in Fig.
 6 and Fig.
-6 (more in Appendix Fig.
-11 to Fig.
-17).
+6 (extended in Appendix Fig.
+1218).
 The results show that CoDyRA adapts the number of active ranks differently across locations and tasks, reflecting the varying demands of each dataset.
 These strike a balance between plasticity and stability, demonstrating the method’s ability to regulate forgetting while enabling effective task learning.
-For the Aircraft dataset, we observe that a greater number of ranks are allocated to the Output projection of the Attention module within the text encoder (Fig.
-6(b)).
-Additionally, a higher concentration of ranks is assigned to deeper layers, particularly layers 6 to 10 (Fig.
+For Aircraft, ranks concentrate in the text encoder’s Attention Output projection (Fig.
+6(b)) and deeper layers 6–10 (Fig.
 6(c)).
-In contrast, for the Pets dataset, more ranks are allocated to the Attention module’s Value and Output projections and the MLP’s Projection layer within the vision encoder (Fig.
-6(b)).
-Additionally, rank allocation is primarily concentrated in the first and last layers (Fig.
-6(c)), indicating a distinct adaptation pattern that balances knowledge retention and adaptation across datasets.
-These observations highlight the adaptive nature of our approach, where rank allocation dynamically adjusts based on dataset characteristics, ensuring effective domain adaptation while preserving pre-trained knowledge.
-We provide more visualizations and analyses in Appendix B.4.
+In contrast, for Pets, allocation shifts to the vision encoder (Attn Value/Output and MLP Projection; Fig.
+6(b)), primarily targeting the first and last layers (Fig.
+6(c)).
+These distinct patterns highlight CoDyRA’s ability to dynamically adjust capacity for effective domain adaptation while preserving existing knowledge. Further analysis is provided in Appendix B.8.
 Figure 5: Visualization and statistical analysis of rank activation on the Aircraft dataset.
 Figure 6: Visualization and statistical analysis of rank activation on the Oxford Pets dataset.
-Table 4: Computation cost.
-Methods
-Training
-Params.
-Additional Infer.
-Params. / Mem.
-LWF (Li & Hoiem, 2017) 129.6M None
-ZSCL (Zheng et al., 2023) 129.6M None
-MoE-Adapters (Yu et al., 2024) 59.8M 13.35M
-RAIL (Xu et al., 2024) N/A 24.18M / 9.01M
-CoDyRA 4.4M None
+Figure 7: Ablations on different insertion locations and the impact of varying the initial rank.
+Figure 8: Analyses (a) maximum pruning threshold and (b) ratios of dense training iterations.
 Computation cost.
@@
 In contrast, our method reduces training overhead by retaining only high-importance ranks; once merged into the pre-trained weights, the updates add zero memory cost at inference.
-Figure 7: Ablations on different insertion locations and the impact of varying the initial rank.
-Figure 8: Analyses (a) maximum pruning threshold and (b) ratios of dense training iterations.
+Table 4: Computation cost.
+Methods
+Training
+Params.
+Additional Infer.
+Params. / Mem.
+LWF [38] 129.6M None
+ZSCL [99] 129.6M None
+MoE-Adapters [93] 59.8M 13.35M
+RAIL [90] N/A 24.18M / 9.01M
+CoDyRA 4.4M None
 4.3 Ablation Studies
@@
 7), consistent with our earlier analyses (Fig.
-4).
+3).
 Training restricted to the vision encoder significantly reduces transfer accuracy, indicating degraded pre-trained capabilities.
@@
 5 Conclusion
-We introduced CoDyRA, a continual learning framework for CLIP that dynamically selects and minimizes LoRA ranks based on their learned importance.
-Guided by sparsity-promoting regularization, CoDyRA balances plasticity with stability, without storing past data, task information, or relying on assumptions.
-By keeping updates close to the model’s previous state while preserving essential ranks, our method mitigates forgetting and enhances representation quality.
-Extensive experiments across diverse continual learning benchmarks demonstrate that CoDyRA achieves state-of-the-art performance, effectively balancing downstream adaptation and knowledge retention.
+We introduced CoDyRA, a dynamic rank-selective framework for continual learning in large pre-trained models, including both CLIP and LLMs.
+Guided by sparsity-promoting regularization, CoDyRA automatically identifies the ranks needed, keeping updates close to the model’s previous state while preserving essential ranks.
+This approach eliminates the need for data replay, task identifiers, or complex architectural expansion.
+Extensive evaluation across diverse benchmarks confirms that CoDyRA achieves superior performance, effectively balancing learning and knowledge retention across modalities and scales.
 Future works.
 CoDyRA updates the model continually by balancing learning and forgetting in the current task.
-While enabling a flexible and efficient CL process with strong results, future work will further refine the modeling of rank relationships, inter-task dependencies, and forgetting measures.
-Reproducibility Statement
-Detailed experiment settings and implementation details are provided in Sec. 4.1 and Appendix B.
+While enabling a flexible and efficient CL process with strong results, future work will further refine modeling of rank relationships, inter-task dependencies, and forgetting measures.
+Impact Statement
+CoDyRA’s sparse, adaptive, importance‐aware low‐rank subspace learning significantly reduces compute and memory demands for continual learning, enabling on‐device personalization in domains such as mobile health and education and empowering resource‐constrained users and organizations.
+As a fundamental research effort, CodyRA does not address potential biases or hazards in training data, and its incremental updates could introduce privacy risks or stealthy backdoors.
+We therefore recommend transparent update logging, regular audits, and differential‐privacy safeguards.
+Although more efficient than full fine‐tuning, large-scale deployment still incurs energy costs; developers should monitor resource use, leverage low-power hardware, and explore carbon offsets.
+Acknowledgments and Disclosure of Funding
+This work was partially supported by the ARC DECRA Fellowship (DE230101591) awarded to D.
+Gong.
+H.
+Lu is affiliated with CSIRO Data61 through a PhD scholarship and acknowledges the support of the Google PhD Fellowship.
 Appendix A Additional Details of the Proposed Method
@@
 where $\mathcal{L}_{\text{sup}}$ is the supervised training loss, $\Delta_{t}$ is the all trainable parameters $\{\mathbf{w}^{m}_{t},\mathbf{B}^{m}_{t},\mathbf{A}^{m}_{t}\}_{m=1}^{M}$ at $t$-th iteration, $M$ is the total number of pre-trained weight matrices with LoRA of rank $r$, $\mathbf{w}^{m}_{t}$ represents the importance weights of the $m$-th LoRA module, and $\lambda>0$ controls the strength of the $\ell_{1}$ regularization.
-We apply proximal gradient descent (Beck & Teboulle, 2009) to handle the non-differentiable $\ell_{1}$ regularization of the importance weights (Ding et al., 2023).
+We apply proximal gradient descent [4] to handle the non-differentiable $\ell_{1}$ regularization of the importance weights [14].
 At the $t$-th iteration, the update rule for the importance weights of each LoRA becomes:
@@
 where the soft-thresholding operator $\mathcal{T}$ is defined as:
-$\mathcal{T}_{\kappa}(x)=\mathbbm{1}(\lvert x\rvert>\kappa)\cdot(x+\mathrm{sign}(x)\cdot\kappa),$ (7)
+$\mathcal{T}_{\kappa}(x)=\mathbbm{1}(\lvert x\rvert>\kappa)\cdot(x{-}\mathrm{sign}(x)\cdot\kappa),$ (7)
 such that the input to the operator $\mathcal{T}_{\kappa}$ is the elements in $\hat{\mathbf{w}}^{m}_{t}$, $\hat{\mathbf{w}}^{m}_{t}$ is the value of $\mathbf{w}^{m}_{t}$ after the gradient update using only $\mathcal{L}_{\text{sup}}(\Delta_{t})$, and $\kappa=\eta_{t}\cdot\lambda$ is the threshold.
@@
 (4).
-Appendix B More Experimental Results and Ablations
-Further details of the experiment settings.
-The MTIL setting consists of 1,201 classes drawn from 11 diverse datasets: Aircraft (Maji et al., 2013), Caltech101 (Fei-Fei et al., 2004), CIFAR100 (Krizhevsky et al., 2009), DTD (Cimpoi et al., 2014), EuroSAT (Helber et al., 2019), Flowers (Nilsback & Zisserman, 2008), Food (Bossard et al., 2014), MNIST (Deng, 2012), OxfordPet (Parkhi et al., 2012), Cars (Krause et al., 2013), and SUN397 (Xiao et al., 2010).
-In the X-TAIL setting, a total of 10 datasets are used, with CIFAR100 (Fei-Fei et al., 2004) excluded to prevent domain overlap, following the protocol in (Xu et al., 2024).
-In line with (Xu et al., 2024), we use a 5-shot split for MTIL and a 16-shot split for X-TAIL.
-For initialization, we follow the standard convention: matrix $\mathbf{A}$ is randomly initialized, matrix $\mathbf{B}$ is set to zero, and the importance weights $\mathbf{w}$ are randomly initialized.
-B.1 More Experiment Results
-Table 5: Extended results for Table 1, comparing continual fine-tuning with different LoRA settings and the LoRA variant of LAE (Gao et al., 2023).
-Method
-Aircraft Caltech101 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
+A.2 More Discussions on Analytical Results of Sec. 3.2
+A.2.1 Takeaway 1 - On Applying LoRA to All Modules
+The original LoRA paper applies LoRA only to selected modules (mainly the attention components).
+Many subsequent LoRA and LoRA-based continual learning works (e.g., InfoLoRA, MoE-Adapters, O-LoRA, etc.) also apply LoRA to only part of the Transformer modules, though the specific choices vary.
+To clarify this design space, we conducted an analysis to determine which module configuration is most effective.
+Our results show that applying LoRA to all layers yields clear benefits.
+While the theoretical analysis of standard Transformers and full LoRA training dynamics is complex, the empirical Neural Tangent Kernel (eNTK) [50] offers a practical framework for theoretical investigation.
+The necessity of applying LoRA to all layers, rather than just attention modules, can be explained through the lens of the eNTK .
+The eNTK of LoRA approximates that of Full Fine-Tuning (FullFT) only when LoRA is applied to the layers that contain the majority of the parameters.
+Since the kernel is based on the dot products of gradients ($K(i,j)=g_{i}\cdot g_{j}$), layers with more parameters (e.g., MLPs) typically exert the most influence on the kernel.
+Restricting updates to attention layers limits the model’s ability to approximate the learning dynamics of FullFT, resulting in "slower learning" or reduced plasticity.
+A.2.2 Takeaway 2 - On the Rank-Dependent Plasticity-Stability Balance
+We formalize the causality of this balance by defining the trade-off as an informational dilemma rooted in the subspace dimensionality:
+High Rank $\rightarrow$ High Plasticity / Low Stability.
+The rank ($r$) dictates the intrinsic dimensionality (degrees of freedom) of the update subspace $\Delta W$.
+A high rank maximizes the informational capacity of the parameter shift.
+In the NTK context, maximizing rank pushes the update closer to the FullFT regime.
+While this allows for rapid acquisition of new task knowledge (High Plasticity), the larger informational perturbation maximizes disruption to the pre-trained knowledge manifold, leading to catastrophic forgetting (Low Stability).
+Low Rank $\rightarrow$ High Stability / Low Plasticity.
+Conversely, restricting rank constrains the degrees of freedom.
+This forces the model to encode only the minimal, essential information for the new task.
+This minimization of total informational perturbation ensures the model remains informationally proximal to its previous state—a necessary condition for stability.
+A lower rank $r$ of the LoRA implies a smaller dimension for the update $\Delta W$.
+A low rank acts as a structural regularizer.
+It forces the update to be the "simplest" change necessary to solve the current task.
+This minimizes the norm $\Delta W||_{F}$ and reduces the probability that the update vector on the critical subspace of previous tasks.
+A.2.3 Takeaway 3 - On the Necessity of Adaptive Optimization
+The need for adaptive rank selection stems from the functional variance of model components and the different information requirements of tasks.
+Following the theory on intrinsic dimension of objective landscapes [36], different tasks require different degrees of freedom to be solved.
+And, not all layers contribute equally to the objectives.
+Adaptivity allows the model to allocate higher ranks to layers with high gradient influence (e.g., specific MLPs) while enforcing sparsity on less critical layers.
+A.3 Connections Between Rank Minimizing and Forgetting
+As shown in Eq.
+(2) and (3), we minimize the rank number to regularize the updating strength $\|\Delta\mathbf{W}^{t,m}\|$, $R(\Delta\mathbf{W}^{t,m})$.
+Minimizing $\|\mathbf{w}^{t,m}\|_{1}$ with a small rank number encourages minimizing of $\|\Delta\mathbf{W}^{t,m}\|_{F}$ if given $\mathbf{A}^{t,m}$ and $\mathbf{B}^{t,m}$ (according to Eq.
+(2)).
+It limits the update $\Delta\mathbf{W}^{t,m}$ to the minimal dimension necessary.
+It softly reduces capacity in updates conflicting with critical information encoded in the learned parameters, without introducing additional assumptions.
+Another line of work regularizes updates (e.g., LoRA weights or gradients) to lie in subspaces orthogonal to existing parameters [80, 40].
+Orthogonal-regularization methods and CoDyRA both aim to mitigate interference, but adopt different strategies with distinct trade-offs, which can be summarized along three aspects.
+(1) Orthogonal regularization models and aims to constrain updates to lie in subspaces orthogonal to previous parameters, which effectively reduces interference but also limits the ability to acquire new knowledge.
+This rigid geometric constraint may suppresses learning new knowledge, and thus is typically softened or relaxed in practice.
+In contrast, CoDyRA regularizes update simplicity (via rank sparsity) rather than geometry, allowing alignment with prior knowledge when beneficial while keeping updates compact to avoid unintended overwriting.
+Both of them aims to alleviate interference (for stability) but have trade-off for learning new tasks (for plasticity).
+(2) Explicit orthogonal-regularization methods typically require storing the information of already the learned subspace (e.g., projection matrices or singular vectors) to compute orthogonality losses, introducing additional memory overhead and dependence on past task statistics.
+In contrast, CoDyRA mitigates interference implicitly through sparsity over rank importance.
+This strictly local objective requires no access to previous tasks, aligning with our goal of natural, memory-free PTM updates.
+(3) Standard orthogonal methods typically assume a fixed rank and focus on rotating it.
+When this rank exceeds a task’s intrinsic dimensionality, the extra dimensions behave as noise, causing unnecessary weight drift.
+CoDyRA actively prunes such dimensions: a rank pruned to zero induces, by definition, “zero interference” with previous tasks.
+This yields “zero overlap” for irrelevant dimensions while preserving flexible updates for task-relevant ones.
+By mitigating forgetting through different mechanisms and trade-offs, CoDyRA and orthogonal-based methods are complementary.
+For example, the task-relevant ranks selected by CoDyRA (with nonzero ${\textnormal{w}}^{t,m}$) can be further regularized using orthogonal constraints.
+Appendix B More Experiments
+B.1 Further Details of Experiments.
+Details of CLIP CL experiment settings.
+The MTIL setting consists of 1,201 classes drawn from 11 diverse datasets: Aircraft [49], Caltech101 [19], CIFAR100 [35], DTD [10], EuroSAT [25], Flowers [57], Food [6], MNIST [13], OxfordPet [58], Cars [34], and SUN397 [89].
+In the X-TAIL setting, a total of 10 datasets are used, with CIFAR100 [19] excluded to prevent domain overlap, following the protocol in [90].
+In line with [90], we evaluate MTIL under both 5-shot and full-shot settings, while X-TAIL is evaluated using a 16-shot setting.
+Details of LLM CL experiment settings.
+We adopt the TRACE benchmark [81], designed for evaluating continual learning in LLMs such as LLaMA [72, 16] and Gemma [70].
+TRACE comprises eight diverse datasets standardized into a unified format, spanning domain-specific understanding, multilingual processing, code generation, and mathematical reasoning.
+Specifically, the tasks are: C-STANCE [98], FOMC [67], MeetingBank [28], Py150 [47], ScienceQA [46], NumGLUE-cm [55], NumGLUE-ds [55], and 20Minuten [22].
+In our setup, each task contributes 5,000 training instances and 2,000 testing instances, totaling 40,000 training and 16,000 testing samples.
+Given this diversity, we employ task-specific metrics to capture performance nuances: Accuracy is used for classification and reasoning tasks (C-STANCE, FOMC, ScienceQA, NumGLUE); ROUGE-L for summarization (MeetingBank); SARI for multilingual text simplification (20Minuten); and a similarity score for code generation (Py150).
+B.2 Evaluation Metrics
+To strictly evaluate the plasticity and stability of our method, we employ the following metrics.
+Let $N$ denote the total number of learned tasks, and $A_{i,j}$ denote the performance (e.g., accuracy, ROUGE, or similarity score) on task $j$ after training on task $i$.
+CLIP Continual Learning Metrics.
+For the CLIP experiments (MTIL and X-TAIL), we report the following metrics as in [99, 93, 90]:
+  • Transfer Accuracy (Transfer): indicates the model’s zero-shot performance on future domains before they are learned, measures the extent to which the pre-trained zero-shot ability is preserved (or improved) throughout incremental learning.
+  • Average Accuracy (Average): indicates the average accuracy of all learning steps across all domains.
+It captures the comprehensive performance stability throughout the entire incremental training process.
+  • Last Accuracy (Last): represent the model’s performance on all seen domains after the training is fully completed.
+LLM Continual Learning Metrics.
+For the TRACE benchmark, we adopt the standard metrics as in [81]:
+  • Overall Performance (OP): The average performance across all learned tasks at the end of training.
+    $\text{OP}=\frac{1}{N}\sum_{j=1}^{N}A_{N,j},$ (8)
+    where $A_{i,j}$ represents the performance on task $j$ after learning task $i$.
+  • Backward Transfer (BWT): Measures the average performance degradation (forgetting) of previous tasks $j<N$ after learning new tasks. A lower BWT indicates less forgetting (better stability).
+    $\text{BWT}=\frac{1}{N-1}\sum_{j=1}^{N-1}(A_{j,j}-A_{N,j}).$ (9)
+B.3 Implementation Details.
+CLIP experiments.
+Following established protocols [99, 93, 90], we employ the CLIP ViT-B/16 backbone [62].
+CoDyRA is applied to all pre-trained weight matrices in both vision and text encoders, initialized with a rank of 16.
+CoDyRA is optimized for 500 iterations per task using AdamW [45], with dense training ratios of 0.5 for X-TAIL and 0.7 for MTIL, setting $\kappa_{\text{max}}=0.005$.
+LLM experiments.
+We adhere to the variable epoch schedule of prior work [60] ({2, 1, 3, 2, 1, 2, 2, 3} epochs).
+We use an initial rank of 16, a dense training ratio of 0.5, and set $\kappa_{\text{max}}=0.005$.
+Optimization employs a learning rate of $1\times 10^{-4}$ and a batch size of 4.
+Training utilizes DeepSpeed ZeRO-2 with BF16 mixed-precision and a maximum context length of 1024 tokens.
+All experiments were conducted on a cluster of four Nvidia H100 GPUs.
+B.4 More Experiment Results
+B.4.1 Cross-domain task-agnostic incremental learning.
+Table 5 presents results on the X-TAIL benchmark [90], a more challenging setting where inference is performed over a mixed label space across all domains.
+CoDyRA outperforms prior methods across all metrics without requiring auxiliary components, surpassing the zero-shot upper bound in transfer accuracy.
+Table 5: Comparisons on X-TAIL for each domain.
+The best and the second best results are highlighted in red and blue, respectively.
+Methods marked with $\dagger$ indicate the use of domain prediction or distribution detection [93, 90].
+    Method Aircraft Caltech101 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
 CLIP
@@
 Transfer
-   Zero-shot (Radford et al., 2021) 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
+   Zero-shot [62] 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
+   LwF [38] 66.6 26.9 19.5 51.0 78.4 26.6 68.9 35.5 56.1 47.7
+   WiSE-FT [85] 70.1 31.9 25.3 56.3 79.8 29.9 74.9 45.6 56.8 52.3
+   iCaRL [64] 71.7 35.0 43.0 63.4 86.9 43.9 87.8 63.7 60.0 61.7
+   ZSCL [99] 73.3 32.6 36.8 62.1 83.8 42.1 83.6 56.5 60.2 59.0
+[…]
+[…]
+   CoDyRA 74.3 36.8 44.2 69.9 83.5 42.8 88.9 64.6 63.4 63.2
+[…]
+Average
+   LwF [38] 24.7 79.7 38.3 36.9 63.9 81.0 36.5 71.9 42.7 56.7 53.2
+   WiSE-FT [85] 27.1 76.5 40.9 31.3 68.7 81.6 31.4 74.7 51.7 58.4 54.2
+   iCaRL [64] 25.4 72.1 37.5 51.6 65.1 87.1 59.1 88.0 63.7 60.1 61.0
+   ZSCL [99] 36.0 75.0 40.7 40.5 71.0 85.3 46.3 83.3 60.7 61.5 60.0
+[…]
+[…]
+   CoDyRA 41.4 81.0 58.7 77.8 83.4 84.6 64.5 90.4 67.2 64.4 71.3
+[…]
+Last
+   LwF [38] 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
+   WiSE-FT [85] 21.8 76.8 42.9 20.8 77.5 84.9 30.7 76.6 75.8 72.5 58.0
+   iCaRL [64] 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
+   ZSCL [99] 33.1 75.3 43.5 35.2 74.6 87.4 50.4 84.2 77.3 73.4 63.4
+[…]
+[…]
+   CoDyRA 37.7 81.5 65.1 89.9 91.4 85.5 96.8 93.3 77.3 73.5 79.2
+[…]
+Table 6: Extended results for Table 5 on X-TAIL for each domain, comparing continual fine-tuning with different LoRA settings and other LoRA variants.
+    Method Aircraft Caltech101 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
+CLIP
+   Zero-shot 23.5 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 58.5
+Transfer
+   Zero-shot [62] 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
    LoRA (rr=4) 74.9 39.4 41.8 67.7 83.3 44.1 87.1 64.2 62.7 62.8
@@
    LoRA (rr=16) 76.4 39.1 37.5 65.9 82.7 41.9 86.9 63.4 62.7 61.8
-   LAE-LoRA (Gao et al., 2023) 76.2 39.9 43.8 66.5 83.3 42.7 87.5 63 62.9 62.9
+   LAE-LoRA [20] 76.2 39.9 43.8 66.5 83.3 42.7 87.5 63 62.9 62.9
+   TreeLoRA [60] - 76.0 36.3 34.0 58.5 77.2 43.3 82.2 49.8 55.8 57.0
+   InfLoRA [40] - 75.8 34.5 29.2 58.1 73.4 38.6 79.5 47.7 50.3 54.1
+   CL-LoRA [24] - 73.3 33.7 29.5 58.5 80.3 43.1 85.5 61.5 59.2 58.3
    CoDyRA 74.3 36.8 44.2 69.9 83.5 42.8 88.9 64.6 63.4 63.2
@@
    LoRA (rr=16) 23.6 78.5 40.5 43.0 70.4 82.8 49.5 87.5 64.9 64.0 60.5
-   LAE-LoRA (Gao et al., 2023) 29.2 81.3 50.7 68.8 77.4 83.8 53.2 88.5 66.3 64.2 66.4
+   LAE-LoRA [20] 29.2 81.3 50.7 68.8 77.4 83.8 53.2 88.5 66.3 64.2 66.4
+   TreeLoRA [60] 19.3 81.4 55.3 63.9 77.6 80.4 64.3 85.0 54.8 57.6 64.0
+   InfLoRA [40] 38.3 84.2 60.1 45.8 79.0 77.5 61.8 82.8 52.2 52.2 63.4
+   CL-LoRA [24] 39.7 79.2 56.7 58.5 79.1 81.1 64.0 78.0 63.6 60.6 66.0
    CoDyRA 41.4 81.0 58.7 77.8 83.4 84.6 64.5 90.4 67.2 64.4 71.3
@@
    LoRA (rr=16) 13.7 78.6 40.3 39.4 66.1 82.7 50.9 85.7 60.3 76.1 59.4
-   LAE-LoRA (Gao et al., 2023) 25.6 81.5 50.2 73.2 79.5 85 70.4 91.1 72.9 75.7 70.5
+   LAE-LoRA [20] 25.6 81.5 50.2 73.2 79.5 85 70.4 91.1 72.9 75.7 70.5
+   TreeLoRA [60] 16.0 79.1 59.2 62.3 83.0 83.6 95.0 89.9 71.9 74.2 71.4
+   InfLoRA [40] 38.3 85.2 66.6 52.9 92.9 81.6 96.6 90.4 70.1 69.3 74.4
+   CL-LoRA [24] 39.7 79.8 62.5 70.9 92.9 81.9 95.2 90.5 72.4 73.1 75.9
    CoDyRA 37.7 81.5 65.1 89.9 91.4 85.5 96.8 93.3 77.3 73.5 79.2
 […]
-B.1.1 Comparison to fixed‐rank LoRA baselines.
-Table 5 reports results for fixed‐rank LoRA (r = 4, 8, 16) and the LoRA variant of LAE.
-Three key observations emerge: (1) As LoRA rank increases, Last accuracy improves while Average accuracy declines, aligning with Fig.
-4: a relatively higher-rank LoRA facilitates learning new tasks but tends to increase forgetting, while lower-rank LoRA mitigates forgetting but limits adaptation.
+B.4.2 Comparison to fixed‐rank LoRA baselines.
+Table 6 reports results for fixed‐rank LoRA (r = 4, 8, 16) and the LoRA variants, including LAE [20], TreeLoRA [60], InfLoRA [40] and CL-LoRA [24].
+Several key observations emerge: (1) As LoRA rank increases, Last accuracy improves while Average accuracy declines, aligning with Fig.
+3: a relatively higher-rank LoRA facilitates learning new tasks but tends to increase forgetting, while lower-rank LoRA mitigates forgetting but limits adaptation.
 (2) Fixed‐rank LoRA shows a substantial Last‐accuracy gap compared to our adaptive method.
@@
 Table 7: Comparisons on full-shot MTIL setting.
-Method
-Aircraft Caltech101 CIFAR100 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
+    Method Aircraft Caltech101 CIFAR100 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
 CLIP
-   Zero-shot (Radford et al., 2021) 24.3 88.4 68.2 44.6 54.9 71 88.5 59.4 89 64.7 65.2 65.2
+   Zero-shot [62] 24.3 88.4 68.2 44.6 54.9 71 88.5 59.4 89 64.7 65.2 65.2
 Transfer
-   Zero-shot (Radford et al., 2021) 74.5 56.9 39.1 51.1 52.6 72.8 60.6 75.1 30.3 55.9 56.9
-   LwF (Li & Hoiem, 2017) 56.6 44.6 32.7 39.3 46.6 68.0 46.0 77.4 31.9 60.5 50.4
-   iCaRL (Rebuffi et al., 2017a) 77.1 61.0 40.5 45.3 54.4 74.6 47.9 76.7 36.3 58.6 57.2
-   LwF-VR (Ding et al., 2022) 73.5 55.6 35.6 41.5 47.0 68.3 53.9 69.3 26.8 51.9 52.3
-   WiSE-FT (Wortsman et al., 2022) 86.0 67.4 45.4 50.4 69.1 87.6 61.8 86.8 60.1 66.8 68.1
-   ZSCL (Zheng et al., 2023) 55.7 50.9 30.4 41.4 49.3 71.8 36.3 77.5 55.3 53.4 53.2
+   Zero-shot [62] 74.5 56.9 39.1 51.1 52.6 72.8 60.6 75.1 30.3 55.9 56.9
+   LwF [38] 56.6 44.6 32.7 39.3 46.6 68.0 46.0 77.4 31.9 60.5 50.4
+   iCaRL [63] 77.1 61.0 40.5 45.3 54.4 74.6 47.9 76.7 36.3 58.6 57.2
+   LwF-VR [15] 73.5 55.6 35.6 41.5 47.0 68.3 53.9 69.3 26.8 51.9 52.3
+   WiSE-FT [85] 86.0 67.4 45.4 50.4 69.1 87.6 61.8 86.8 60.1 66.8 68.1
+   ZSCL [99] 55.7 50.9 30.4 41.4 49.3 71.8 36.3 77.5 55.3 53.4 53.2
 […]
@@
 Average
-   LwF (Li & Hoiem, 2017) 36.3 86.9 72.0 59.0 73.7 60.0 73.6 74.8 80.0 37.3 58.1 64.7
-   iCaRL (Rebuffi et al., 2017a) 35.5 89.2 72.2 60.6 68.8 70.0 78.2 62.3 81.8 41.2 62.5 65.7
-   LwF-VR (Ding et al., 2022) 29.6 87.7 74.4 59.5 72.4 63.6 77.0 66.7 81.2 43.7 60.7 65.1
-   WiSE-FT (Wortsman et al., 2022) 26.7 86.5 64.4 57.1 65.7 58.7 71.1 70.5 75.8 36.9 54.6 60.7
-   ZSCL (Zheng et al., 2023) 45.1 92.0 80.1 64.3 79.5 81.6 89.6 75.2 88.9 64.7 68.0 75.4
+   LwF [38] 36.3 86.9 72.0 59.0 73.7 60.0 73.6 74.8 80.0 37.3 58.1 64.7
+   iCaRL [63] 35.5 89.2 72.2 60.6 68.8 70.0 78.2 62.3 81.8 41.2 62.5 65.7
+   LwF-VR [15] 29.6 87.7 74.4 59.5 72.4 63.6 77.0 66.7 81.2 43.7 60.7 65.1
+   WiSE-FT [85] 26.7 86.5 64.4 57.1 65.7 58.7 71.1 70.5 75.8 36.9 54.6 60.7
+   ZSCL [99] 45.1 92.0 80.1 64.3 79.5 81.6 89.6 75.2 88.9 64.7 68.0 75.4
 […]
@@
 Last
-   LwF (Li & Hoiem, 2017) 26.3 87.5 71.9 66.6 79.9 66.9 83.8 99.6 92.1 66.1 80.4 74.6
-   iCaRL (Rebuffi et al., 2017a) 35.8 93.0 77.0 70.2 83.3 88.5 90.4 86.7 93.2 81.2 81.9 80.1
-   LwF-VR (Ding et al., 2022) 20.5 89.8 72.3 67.6 85.5 73.8 85.7 99.6 93.1 73.3 80.9 76.6
-   WiSE-FT (Wortsman et al., 2022) 27.2 90.8 68.0 68.9 86.9 74.0 87.6 99.6 92.6 77.8 81.3 77.7
-   ZSCL (Zheng et al., 2023) 40.6 92.2 81.3 70.5 94.8 90.5 91.9 98.7 93.9 85.3 80.2 83.6
+   LwF [38] 26.3 87.5 71.9 66.6 79.9 66.9 83.8 99.6 92.1 66.1 80.4 74.6
+   iCaRL [63] 35.8 93.0 77.0 70.2 83.3 88.5 90.4 86.7 93.2 81.2 81.9 80.1
+   LwF-VR [15] 20.5 89.8 72.3 67.6 85.5 73.8 85.7 99.6 93.1 73.3 80.9 76.6
+   WiSE-FT [85] 27.2 90.8 68.0 68.9 86.9 74.0 87.6 99.6 92.6 77.8 81.3 77.7
+   ZSCL [99] 40.6 92.2 81.3 70.5 94.8 90.5 91.9 98.7 93.9 85.3 80.2 83.6
 […]
@@
 […]
-B.1.2 Additional results on full-shot MTIL.
+B.4.3 Additional results on full-shot MTIL.
 To further extend our evaluation, we report full-shot results on the MTIL benchmark (Table 7).
 For fairness, we adopt the same domain-prediction design used in recent methods.
-Consistent with Tables 1 and 3, CoDyRA achieves state-of-the-art performance, surpassing all baselines on both Transfer (unseen-domain generalization) and Last (continual downstream learning).
-Table 6 also compares against prompt-based approaches (L2P (Wang et al., 2022e), DualPrompt (Wang et al., 2022d), S-Prompt (Wang et al., 2022c)).
+Consistent with Tables 5 and 1, CoDyRA achieves state-of-the-art performance, surpassing all baselines on both Transfer (unseen-domain generalization) and Last (continual downstream learning).
+Table 7 also compares against prompt-based approaches (L2P [84], DualPrompt [83], S-Prompt [82]).
 Although these methods mitigate task forgetting via key–value matching for prompt selection, they neglect the preservation of pre-trained capabilities—evidenced by substantial drops in transfer accuracy—and further underperform recent domain-prediction approaches on last accuracy.
@@
 CoDyRA Last $37.7^{\pm 0.42}$ $81.5^{\pm 0.24}$ $65.1^{\pm 0.63}$ $89.9^{\pm 0.55}$ $91.4^{\pm 0.38}$ $85.5^{\pm 0.16}$ $96.8^{\pm 0.08}$ $93.3^{\pm 0.3}$ $77.3^{\pm 0.66}$ $73.5^{\pm 0.21}$ $79.2^{\pm 0.18}$
-B.1.3 Training stability.
+B.4.4 Training stability.
 Table 8 reports the mean and standard deviation of CoDyRA’s results in Table 5, averaged over three independent runs.
 The consistently low variance across metrics demonstrates that our method is stable to random initialization, highlighting CoDyRA’s robust adaptive rank selection.
-B.2 Effects of Additional Adaptive Prediction Head on Visual Representations
-The related continual learning methods focus on updating the representation with the backbone model (e.g., (Wang et al., 2022e; Zheng et al., 2023; Yu et al., 2024), and CoDyRA) or newly added representation alignment or prediction head (McDonnell et al., 2024; Xu et al., 2024).
+B.4.5 Analysis in forgetting
+We provide a dedicated forgetting analysis in Table 9, quantifying forgetting via Backward Transfer (BWT).
+It reports BWT of CoDyRA and some relevant CL methods with representation updating.
+Fixed-rank LoRA serves as a baseline and exhibits significant forgetting due to unconstrained updates without regularization.
+ZSCL relies on the replay of reference data to help mitigate forgetting.
+MoE-Adapters isolates task-wise knowledge within specific adapters and adopts domain prediction to further reduce interference.
+InfLoRA constrains parameter updates to a direction orthogonal to prior tasks to reduce interference.
+CoDyRA effectively mitigates forgetting by minimizing the rank of updates via sparsity regularization.
+This ensures the model learns effectively while remaining close to its original state, demonstrating strong retention capabilities with a BWT of only $1.87\%$.
+Moreover, in Table 2, we validate our method in the language domain. The low BWT scores observed in this setting further support our claims of reduced forgetting across architectures.
+Table 9: Forgetting analysis. We report Backward Transfer (BWT) averaged across all X-TAIL tasks, where lower values indicate better retention.
+Methods ZSCL MoE-Adapter InfLoRA LoRA (rr=4) LoRA (rr=8) LoRA (rr=16) CoDyRA
+BWT (%) $\downarrow$ 8.78 5.43 3.12 13.94 14.24 14.86 1.87
+B.5 Effects of Additional Adaptive Prediction Head on Visual Representations
+The related continual learning methods focus on updating the representation with the backbone model (e.g., [84, 99, 93], and CoDyRA) or newly added representation alignment or prediction head [52, 90].
 These two types of methods can be seen as orthogonal.
-We study how the representation updated by the proposed can work with the additional adaptive prediction head, relying on the aggression-based adapters in (Xu et al., 2024).
+We study how the representation updated by the proposed can work with the additional adaptive prediction head, relying on the aggression-based adapters in [90].
 We provide a more detailed analysis of the effects of the additional prediction head on the visual representations in X-TAIL setting in Table 10.
-Note that the Dual version of RAIL (Xu et al., 2024) stores all visual representations of all samples for all seen classes and domains in memory, it is not included in Tables 1 and 3 for a fair comparison.
+Note that the Dual version of RAIL [90] stores all visual representations of all samples for all seen classes and domains in memory, it is not included in Tables 5 and 1 for a fair comparison.
 RAIL is a regression-based method that applies an additional prediction head on top of the visual representations extracted by the frozen pre-trained vision encoder of the CLIP model.
@@
 CoDyRA + RAIL-Dual 63.1 74.2 83.6
-B.3 Analysis of Changes/Interference w.r.t. Pre-Trained Weights
+B.6 Analysis of Changes/Interference w.r.t. Pre-Trained Weights
 In Sec.
@@
 6 visualize and analyze the dynamic ranks assigned to each module.
-Here, we extend the analysis by comparing parameter changes to pre-trained weights using the amplification factor as in (Hu et al., 2021).
+Here, we extend the analysis by comparing parameter changes to pre-trained weights using the amplification factor as in [27].
 The amplification factor is given by:
@@
 $\mathbf{U}$ and $\mathbf{V}$ represent the top $r$ singular vectors of $\text{SVD}(\Delta\mathbf{W})$, where $r$ is the number of interested ranks in analysis.
-A higher amplification factor indicates that low-rank parameter updates amplify directions less emphasized in the pre-trained weights (Hu et al., 2021).
+A higher amplification factor indicates that low-rank parameter updates amplify directions less emphasized in the pre-trained weights [27].
 We visualize the amplification factors for both our method and vanilla LoRA trained on the Aircraft, EuroSAT and Oxford Pets datasets in Fig.
@@
 6.
-The number of activated ranks reflects the quantity of significant ranks contributing to parameter updates for leaning each task.
+The number of activated ranks reflects the quantity of significant ranks contributing to parameter updates for learning each task.
 In this analysis, the amplification factor directly measures the parameter change to each pre-trained weight matrix.
@@
 It shows that the proposed method with adaptive rank selection can achieve more focused/concentrated updating and less interference.
-B.4 More Visualizations and Analyses on Rank allocation results of different datasets
+B.7 Parameter-space Directional Analysis
+To further investigate how CoDyRA structures its parameter updates across sequential tasks, we conduct a parameter-space directional analysis following prior work on directional interference in LoRA-based continual learning.
+Specifically, we compute the cosine similarity between the LoRA $\mathbf{A}$-matrix updates obtained after each task, comparing every task’s update direction against those from all other tasks.
+This metric captures the degree to which updates for different tasks align or interfere: higher similarity indicates that tasks induce overlapping update directions, while lower similarity reflects more orthogonal and thus less interfering modifications.
+As shown in Figure 10, CoDyRA consistently produces lower cross-task similarity than a baseline that trains a new LoRA adapter per task and merges them without additional constraints.
+Despite imposing no explicit orthogonality regularization, CoDyRA naturally encourages more decorrelated update directions, revealing an inherent structural property that helps mitigate destructive interference between tasks.
+B.8 More Visualizations and Analyses on Rank allocation results of different datasets
 In Sec.
-4.3, Fig.
+4.2, Fig.
 6 and Fig.
@@
 12 to Fig.
-17.
+18.
 The visualizations, along with statistical analyses of transformer modules and layers, reveal distinct rank allocation patterns across datasets.
 These findings suggest that the rank of parameter updates needed for achieving a good downstream improvements and knowledge retention is distinct across each kind of data.
+Figure 10: Parameter-space directional analysis of $\mathbf{A}$-matrix updates.
+For each task, the column displays the average cosine similarity between the model’s LoRA updates after a specific task and the LoRA updates from models trained on all other tasks.
+This analysis quantifies directional interference among sequential updates.
+Lower cosine similarity reflects more orthogonal (less interfering) updates.
+CoDyRA produces lower similarity than the simple LoRA–merging baseline, demonstrating stronger implicit orthogonality of parameter updates even without explicit orthogonal regularization.
 Figure 11: Visualization and statistical analysis of rank activation on the Caltech101 dataset using our proposed method.
@@
 Figure 18: Visualization and statistical analysis of rank activation on the SUN397 dataset using our proposed method.
-B.5 Evaluation with Other VLMs
-To evaluate the broader applicability of our method, we conducted preliminary experiments with BLIP (Li et al., 2022) on X-TAIL (Table 9).
-Our method consistently outperformed prior works and exceeded the pre-trained model’s upper-bound performance, demonstrating its general effectiveness across different VLMs.
 Table 11: Averaged performance using BLIP.
@@
 CoDyRA  ($r_{init}=16$) 97.20 / 3.95M 81.61 / 3.75M 94.44 / 3.85M
-B.6 Evaluation on Standard Parameter-Efficient Fine-Tuning (PEFT) Task
+B.9 Evaluation with Other VLMs
+To evaluate the broader applicability of our method, we conducted preliminary experiments with BLIP [37] on X-TAIL (Table 11).
+Our method consistently outperformed prior works and exceeded the pre-trained model’s upper-bound performance, demonstrating its general effectiveness across different VLMs.
+B.10 Evaluation on Standard Parameter-Efficient Fine-Tuning (PEFT) Task
 Leveraging our dynamic rank-adaptive parameter updates, our method naturally extends to Parameter-Efficient Fine-Tuning (PEFT) tasks.